### v1

# MnLargeSymbols’164 MnLargeSymbols’171

Abstract.
Secure collaborative analytics (SCA) enable the processing of analytical SQL queries across multiple owners’ data, even when direct data sharing is not feasible. Although essential for strong privacy, the large overhead from data-oblivious primitives in traditional SCA has hindered its practical adoption. Recent SCA variants that permit controlled leakages under differential privacy (DP) show a better balance between privacy and efficiency. However, they still face significant challenges, such as potentially unbounded privacy loss, suboptimal query planning, and lossy processing.
To address these challenges, we introduce SPECIAL, the first SCA system that simultaneously ensures bounded privacy loss, advanced query planning, and lossless processing. SPECIAL employs a novel synopsis-assisted secure processing model, where a one-time privacy cost is spent to acquire private synopses (table statistics) from owner data. These synopses then allow SPECIAL to estimate (compaction) sizes for secure operations (e.g., filter, join) and index encrypted data without extra privacy loss. Crucially, these estimates and indexes can be prepared before runtime, thereby facilitating efficient query planning and accurate cost estimations. Moreover, by using one-sided noise mechanisms and private upper bound techniques, SPECIAL ensures strict lossless processing for complex queries (e.g., multi-join). Through a comprehensive benchmark, we show that SPECIAL significantly outperforms cutting-edge SCAs, with up to 80×80$\times$80 × faster query times and over 900×900$\times$900 × smaller memory for complex queries. Moreover, it also achieves up to an 89×89$\times$89 × reduction in privacy loss under continual processing.
1. Introduction
Organizations, such as hospitals, frequently hold sensitive data in separate silos to comply with privacy laws, despite the valuable insights that could be gained from sharing this information. Recent advancement of Secure Collaborative Analytics (SCA) (Eskandarian and Zaharia, 2017; Qin et al., 2022; Roy Chowdhury et al., 2020; Wang et al., 2021, 2022; Bater et al., 2017, 2018; Liagouris et al., 2023; Poddar et al., 2021; Bater et al., 2020; He et al., 2015) provides an exciting solution to tackle this dilemma. These systems leverage advanced multi-party secure computation (MPC) (Yao, 1986) primitives to empower multiple data owners, who previously could not directly share data, to collaboratively process analytical queries over their combined data while ensuring the privacy of each individual’s data.
While MPC can effectively conceal data values (Yao, 1986), its security guarantees do not immediately extend to the protection of execution transcripts. Consequently, data-dependent processing patterns such as memory traces and read/write volumes can still reveal critical information, risking privacy breaches (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016) even when the core data remains encrypted. To ensure strong privacy, modern SCA systems also utilize data-oblivious primitives that exhaustively pad query processing complexities to a worst-case, data-independent upper bound (Bater et al., 2017; Liagouris et al., 2023; Poddar et al., 2021). However, such stringent protections can largely reduce system efficiency and hinder the generalization of conventional optimization techniques to SCA, which are typically data-dependent (Liagouris et al., 2023). To address this, recent efforts (Qin et al., 2022; Wang et al., 2021, 2022; Bater et al., 2018) have introduced differentially private SCA (DPSCA). This approach allows controlled information leakage under differential privacy (Dwork et al., 2014) to mitigate constant worst-case overhead. For instance, systems under this model can dynamically compact intermediate query sizes to a noisy estimate close to actual sizes, avoiding exhaustive padding. As such, queries under DPSCA experience largely boosted efficiencies (e.g., up to $10^{5}\times$ faster (Wang et al., 2022)) compared to their “no leakage” counterparts. Despite these substantial performance gains, existing DPSCAs still face critical limitations that impede their practical uses, as elaborated below:
$\bullet$ L-1. Unbounded privacy loss. Most DPSCA systems utilize a per-operator privacy expenditure model (Bater et al., 2018; Wang et al., 2021; Qin et al., 2022; Chu et al., 2021; Chan et al., 2022; Wang et al., 2022), meaning each query operator (e.g., join, filter) independently consumes a portion of the privacy budget. This approach can lead to either unbounded privacy loss or the forced cessation of query responses upon budget exhaustion. To mitigate this, some studies (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., [n.d.]) propose private, locality-sensitive grouping, incurring a one-time privacy cost to pre-group data based on specific attributes. Subsequent queries on those attributes can be directly applied to a smaller subset and need no additional privacy budget. However, this method only supports simple queries (e.g., point and range); complex queries like joins still suffer from unbounded privacy loss.
$\bullet$ L-2. Suboptimal execution plan. Conventional query planners can pre-estimate sizes for equivalent plans of a given query and select the most efficient plan with minimized intermediate sizes before execution (Silberschatz et al., 2011; Blasgen et al., 1981). In contrast, SCA systems lack this capability, and even DPSCA designs (Qin et al., 2022; Wang et al., 2022; Zhang et al., 2023; Wagh et al., 2021; Bater et al., 2018, 2020) can only reactively determine plan sizes during runtime. This inherent limitation often forces existing systems to settle for less efficient query plans, such as suboptimal join orders, which lead to significantly inflated intermediate sizes (§ 7.2) and substantially hinder performance.
$\bullet$ L-3. Lossy processing. Noise from randomized mechanisms in DPSCA also introduces a unique accuracy issue (e.g., losing tuples after compaction), and unfortunately, no existing design can provide deterministic accuracy guarantees for complex queries (Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Zhang et al., 2023; Groce et al., 2019). Furthermore, stronger privacy settings can further increase noise variance, which amplifies errors, significantly impacting the utility of SCA systems.
1.1. Overview of SPECIAL.
In this work, we introduce SPECIAL, an innovative SCA system designed to address all these challenges simultaneously through a new paradigm called synopsis-assisted secure processing. At its core, SPECIAL incurs a one-time privacy cost to gather DP synopses (statistics of base tables) from owners’ data. These synopses are then used to compact query intermediate results, index encrypted data, and enhance SCA query planning. Notably, SPECIAL is the first system to provide all of the following benefits: (1) Bounded privacy—it manages complex queries, such as multi joins, within strict privacy limits; (2) Advanced query planning—it builds an advanced SCA planner that can exploit plan sizes before runtime; and (3) Lossless processing—it ensures exact results with no data omissions. An overview of SPECIAL’s architecture is shown in Figure 1.
Figure 1. Overview of SPECIAL workflow.
SPECIAL operates under a standard server-aided MPC model (Kamara et al., 2011) with three key participants: data owners, at least two SPECIAL servers, and a vetted analyst. The process starts with data owners securely outsourcing their data (e.g. using secret sharing (Kamara et al., 2011; Wang et al., 2021; Zhang et al., 2023; Wang et al., 2022)), and privately release a set of DP synopses (table statistics) to the servers. Once the data and synopses are ready, analysts can then submit queries on the outsourced data. To process queries, SPECIAL uses synopses to privately estimate tight result compaction bounds or build private indexes for faster access, thus avoiding exhaustive padding. Crucially, this information is exploit before runtime, so that SPECIAL can pre-identify an optimal secure execution plan that minimizes intermediate sizes and overall costs. Finally, the optimized plan is executed via MPC across the secret-shared data, and the results are securely returned to the analyst.
In the SPECIAL prototype, we adopt the same privacy model used in existing DPSCA designs, assuming a computationally bounded, “honest-but-curious” adversary (Goldreich, 2004) who can compromise up to $n-1$ out of $n$ owners and $m-1$ out of $m$ servers. SPECIAL ensures that such an adversary can only learn a limited leakage profile $\mathsf{Lkg}$ about honest parties’ data, which strictly adheres to DP constraints. Note that, SPECIAL’s core components, including synopses management and query planning do not leverage raw data at all. Thus, any of the servers, alone or jointly, can manage these tasks. Also, once analysts are authorized by data owners, they are considered trusted, and additional randomization of query results is not required. It is worth mentioning that SPECIAL’s corruption model (e.g. all except one) is entirely determined by the underlying MPC, and SPECIAL’s design does not modify them at all. Hence, one may also opt to a weaker corruption model, such as a supermajority of owners and servers must remain uncorrupted (Liagouris et al., 2023; Tan et al., 2021), to enhance efficiency.
1.2. Unique challenges and key contributions
Leveraging DP synopses in SCA holds significant promise for achieving our desired objectives. However, this also introduces unique challenges. Below, we highlight the key challenges and summarize our non-trivial contributions to address them:
$\bullet$ C-1. How to select proper synopses? Even a single relation can have numerous options for generating synopses, such as using different attributes or their combinations. However, improper selection can lead to large errors (e.g. using too many synopses or leverage high-dimensional attributes (Zhang et al., 2021)), or reduced functionalities (e.g., using only simple attributes (Zhang et al., 2023)). Hence, a key challenge is selecting a limited set of DP synopses to optimize the privacy budget for complex query processing. Our approach is informed by two observations: (i) secure joins are resource-intensive and need prioritized acceleration, and (ii) synopses for common filtering predicates are vital as they allow pre-built indexes on base relations for fast access. Consequently, we propose a focused strategy (§ 4) that targets low-dimensional (1D and 2D) attributes frequently involved in joins and filters within a representative workload.
$\bullet$ C-2. How to enforce lossless processing? Private synopses do not immediately implies lossless guarantees. Thus, a second challenge is designing practical approaches to achieve lossless results without violating privacy goals. To address this, we employ one-sided DP noise (either strictly positive or negative, § 4) for generating synopses and design novel primitives (§ 5) based on them to pessimistically estimate filter cardinalities and intervals of index structures. This enable us to efficiently discard unnecessary records during processing without breaking privacy and lossless guarantees. To ensure lossless processing of complex joins, we extend upon cutting-edge join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join compaction sizes using available synopses. To our knowledge, this is the first study to support private join upper bound estimation.
$\bullet$ C-3. How DP synopses can empower efficient query processing? The use of DP synopses in SCA is largely underexplored, leaving a knowledge gap regarding their potential to enhance query efficiency. To navigate this potential, we explore various use of synopses in accelerating secure processing including private indexes SPEidx (§ 5.2), and compacted oblivious operations SPEop (§ 5.3). We then develop a Selinger-style query planner (§ 6) that supports both standard secure operators and SPEop. Employing a customized cost model and tailored heuristics, this planner efficiently generates optimal execution plans for queries using a mix of standard and SPEop operators. This in essence, also address the limitations of sub-optimal execution plans inherent in standard DPSCA.
$\bullet$ C-4. How to systematically evaluate SPECIAL? A major challenge in SCA design evaluation is the absence of open-source benchmarks. We address this by initiating an open-source evaluation platform accessible to the public, utilizing public Financial data (Rel, [n.d.]) (1.1M rows and 59 columns), and designing eight test queries ranging from simple linear to complex 5-way joins. We evaluated our prototype SPECIAL alongside the state-of-the-art DPSCA system, Shrinkwrap (Bater et al., 2018), and conventional SCA, SMCQL (Bater et al., 2017). Results indicate that SPECIAL outperforms Shrinkwrap, reducing query latency by up to 80.3×80.3$\times$80.3 ×, and SMCQL, with at least a 114×114$\times$114 × reduction in query latency. Additionally, SPECIAL improves memory efficiency in complex join processing by more than 900×900$\times$900 × compared to both systems. Moreover, scaling experiments show that SPECIAL can effectively scale linear and binary joins up to 8×8$\times$8 × (8.8M rows) and 5-way joins up to 4×4$\times$4 × (4.4M rows) workloads. All benchmarks, including our prototype implementation, are open-sourced and available at (Lovingmage, 2024).
The remainder of the paper is structured as follows: We begin with essential background in § 2 and a formal definition of our privacy model in § 3 before delving into design details. § 4 discusses the SPECIAL catalog design, including synopses selection and generation. § 5 details novel synopses-assisted secure database operators, and § 6 describes our planner design. Benchmark evaluations are presented in § 7, followed related works and conclusions.
2. Background
General notations. We consider the logical database $\mathcal{D}$ to contain multiple private relations $\{D_{1},D_{2},...\}$, where each relation $D_{i}$ is belongs owned by a specific party $P_{i}$, where $D\in\mathcal{D}$ is owned by a specific owner and has a set of attributes as $attr(D)$. The domain for attribute $A\in attr(D)$ is given by $dom(A)$, and the combined domain for a collection of attributes $\mathbf{A}=\{A_{1},A_{2},...\}\subseteq attr(D)$ is denoted as $dom(\mathbf{A})=\prod_{A\in\mathbf{A}}dom(A)$. For a tuple $t\in D$, and $\mathbf{A}\subseteq attr(D)$, we denote $t.\mathbf{A}$ as the attribute value of $\mathbf{A}$ in $t$
Frequency (count). Given $D$, and $\mathbf{A}\subseteq attr(D)$, and $\mathbf{v}\in dom(\mathbf{A})$, the frequency (count) of $\mathbf{v}$ in $D$ is the total number of tuples $t\in D$ with $t.\mathbf{A}=\mathbf{v}$. In addition, the max frequency moments (MF) of $\mathbf{A}$ is defined as ${\mathtt{mf}}(\mathbf{A},D)=\max_{\mathbf{v}\in dom(\mathbf{A})}|\{t\in D\mid t.\mathbf{A}=\mathbf{v}\}|$.
Histograms. Given $D$, and $\mathbf{A}\subseteq attr(D)$, the (equal-width) histogram $\mathbf{h}(\mathbf{A},D)=(c_{1},c_{2},...,c_{m})$ is a list of counts for the attribute values in $\mathbf{A}$. Specifically, $\mathbf{h}$ partitions [^1] We use a condensed notation that omits arguments when $\mathbf{A}$ and $D$ are obvious. $dom(\mathbf{A})$ into $m$ “equal-sized” domain intervals $(B_{1},...,B_{m})$, and each count $c_{i}\in\mathbf{h}$ corresponds to the total number of $t\in D$ such that $t.\mathbf{A}$ fall within $B_{i}$.
Multi-party secure computation (MPC). MPC (Yao, 1986; Goldreich, 2009; Micali et al., 1987; Ben-Or et al., 2019) is a cryptographic techniques to allow a group of participants $P_{1},P_{2},...$ to jointly compute a function $f(x_{1},x_{2},...)$ over their respective private input $x_{i}$. MPC ensures that no unauthorized information is revealed to any involved party, excepted the desired output of function $f$, emulating a computation as if performed by a trusted third party. Earlier MPC typically require all parties to participate in intensive computations. Recent schemes (Kamara et al., 2011; Scholl et al., 2017; Mohassel and Zhang, 2017) allow these computations to be offloaded to multiple powerful servers without sacrificing original security guarantees, known as the server-aided MPC. In this approach, each party, say $P_{i}$, secretly shares their inputs, $x_{i}$ to servers. These servers then jointly evaluate an MPC that functionally replicates the logic of first reconstructing the secrets $x_{1},x_{2},\ldots$ and then evaluating $f(x_{1},x_{2},\ldots)$. Depending on designs, the result may be directly returned to the parties or re-shared among the servers for later retrieval. By default, SPECIAL considers server-aided MPC with two non-colluding servers.
Secret sharing and secure array. SPECIAL uses the $2$-out-of-$2$ boolean secret share (Araki et al., 2016) over ring $\mathbb{Z}_{2^{32}}$ for securely outsourcing owners’ data and storing query execution results. Specifically, each data, $x$, is divided into two shares: $x1,x2$ that are uniformly distributed over the ring $\mathbb{Z}_{2^{32}}$ such that $x=x1\oplus x2$. Each server $S_{i}$ receives one secret shares, $s_{i}$, where $i\in\{0,1\}$. By retrieving shares from any two servers, an authorized party can successfully reconstruct the value of $x$. However, a single server alone learns nothing about $x$. For clarity and to abstract out the lower-level details, we leverage a logically unified data structure, namely the secure array (Wang et al., 2022; Bater et al., 2018), denoted as $\langle\mathbf{x}\rangle=\left(\langle x_{1}\rangle,\langle x_{2}\rangle,...\right)$, which is a collection of secret-shared relational tuples.
Oblivious (relational) operators. Oblivious operators are specific MPC protocols, applied over secure arrays, that implement database operator functionalities such as filter, join, and aggregation. Specifically, an oblivious operator $\langle\mathbf{o}\rangle\leftarrow\mathsf{\textbf{op}}(\langle\mathbf{a}\rangle,\langle\mathbf{b}\rangle,...)$ accepts one or more secure arrays as inputs, performs database operations over the inputs, and then writes the result to an output secure array. Moreover, the oblivious property stipulates that absolutely no memory access pattern information is disclosed (Chang et al., 2022; Ramachandran and Shi, 2021; Liagouris et al., 2023) (or in other words, the memory access is fully independent from input data) during the protocol’s execution. Since revealing output size can also enable numerous reconstruction attacks, recent designs of oblivious operators also consider hiding output volumes, i.e., by padding the outputs to their maximum sizes. Thus, unless otherwise stated, we assume that all the subsequent oblivious operators mentioned have considered volume-hiding.
Differential privacy (Dwork et al., 2014). DP ensures that modifying a single input data tuple to a mechanism produces only a negligible change in its output. To elaborate, consider $D$ and $D^{\prime}$ as two databases differing by just one tuple—termed as neighboring databases. Then:
Definition 2.1 ($(\epsilon,\delta)$-differential privacy).
Given $\epsilon>0$, and $\delta\in(0,1)$. A randomized mechanism $\mathcal{M}$ is said to be $(\epsilon,\delta)$-DP if for all pairs of neighboring databases $D,D^{\prime}$, and any possible output $o\subset Range(\mathcal{M})$, the following holds:
$\textup{Pr}\left[\mathcal{M}(D)\in o\right]\leq e^{\epsilon}\textup{Pr}\left[\mathcal{M}(D^{\prime})\in o\right]+\delta$
3. Privacy Model
In this section, we formally define the privacy definitions of SPECIAL and explain the corresponding semantics. Similar to previous work on DPSCA (Wang et al., 2022; Zhang et al., 2023), we follow computational DP (SIM-CDP) (Mironov et al., 2009) framework for privacy model formulation.
Definition 3.1 (Secure protocol with DP leakage).
Given a party (owner) $P$ with private data $D$ and a randomized mechanism $\mathsf{Lkg}(D)$, referred to as the leakage profile of $P$’s data. Consider $\Pi$ a secure query processing protocol applied to $D$. The protocol $\Pi$ is said to be secure with DP leakage if, for any honest party $P$ with data $D$ and any probabilistic polynomial time (p.p.t.) adversary $\mathcal{A}$:
  • $\mathsf{Lkg}(D)$ satisfies $(\epsilon,\delta)$-DP (definition 2.1).
  • There exists a p.p.t. simulator $\mathcal{S}$ with only access to public parameters $\mathsf{pp}$ and the output of $\mathsf{Lkg}(D)$ that satisfies:
    (1) $\begin{split}&\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\Pi}(\mathcal{D},\mathsf{pp})=1\right)\right]\\ &\leq\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\mathcal{S}}(F(\mathcal{D}),\mathsf{pp})\right)=1\right]+\mathsf{negl}(\kappa)\end{split}$
where $\mathsf{VIEW}^{\Pi}$, and $\mathsf{VIEW}^{\mathcal{S}}$ denotes the adversary’s view against the protocol execution and the simulator’s outputs, respectively. And $\mathsf{negl}(\kappa)$ is a negligible function related to a security parameter $\kappa$.
Simply put, Definition 3.1 specifies that as long as there is at least one honest owner, the knowledge any p.p.t. adversary can obtain about this owner’s data from observing $\Pi$ execution transcripts is bounded to what can be inferred from the outputs of an $(\epsilon,\delta)$-DP mechanism $\mathsf{Lkg}$. By default, we define $\mathsf{Lkg}$ within $(\epsilon,\delta)$-tuple level DP (Dwork et al., 2010), that’s said primary privacy objective is to protect individual tuples. However, due to group-privacy properties (Kifer and Machanavajjhala, 2011; Xiao and Xiong, 2015) of DP, privacy can extend across multiple tuples and to user-level, as long as a user owns a limited number of tuples. Throughout this paper, our focus remains on developing algorithms that guarantee tuple-level privacy, while upholding all stated privacy guarantees, albeit with potentially different privacy parameters.
Theorem 3.2.
The privacy of SPECIAL adheres to Definition 3.1
Proof.
Due to space concerns, we defer the complete privacy proof of SPECIAL to Appendix A. ∎
4. SPECIAL Synopses
SPECIAL maintains independent DP synopses for all outsourced data. Each synopsis holds noisy statistics that reflect the corresponding data’s distributional traits. As per our core design idea, these statistics will be used to dynamically optimize query processing—for example, by estimating tight compaction bounds or enabling fast access to eliminate unnecessary tuples, which significantly compacts intermediate query sizes, all without any extra privacy cost.
Challenges. A major design challenge is selecting proper synopses to achieve our design goals. This involves decisions such as attribute selection, combinations thereof, and the types of statistics used (e.g., histograms, min/max/avg, frequency counts, etc.). Using too many attributes or leveraging complex (high-dimensional) combinations can introduce significant noise (Zhang et al., 2021) and increase storage costs. Another challenge involves privately releasing synopses to aid in lossless processing, such as result compaction and private indexing. Ensuring lossless compaction requires consistently overestimating the output sizes of subplans. Supporting lossless indexes is even more complex: for a given indexed range $D[\mathsf{lo},\mathsf{hi}]$, we need to overestimate the upper index $\mathsf{hi}$ and underestimate the lower index $\mathsf{lo}$. Unfortunately, traditional DP methods, which typically use symmetric additive noises, struggle to ensure these conditions.
Key ideas. We observe that joins are more resource-intensive than any other operations. For instance a $k$-way oblivious join can incur a $O(n^{k})$ complexity if without optimisations (Bater et al., 2017, 2018; Wang et al., 2022; Zheng et al., 2017; Eskandarian and Zaharia, 2017). Hence, joins require prioritized acceleration. Moreover, synopses available for frequently queried filter attributes on base relations are vital (Wang et al., 2022) as well. These statistics not only help to compact input data at an early stage before heavy joins, but also enable fast indexing (§ 5.2) for quick data access. Hence, to address the first challenge, our key idea is to focus on commonly queried join and base relation filter attributes, as well as their low-dimensional combinations.
To tackle the second challenge, our key idea is twofold. First, to ensure lossless result compaction of filters and indexing, we employ one-sided DP noise to generate private histograms that either overestimate or underestimate attribute distributions consistently. Second, to achieve lossless compaction of join outputs, we integrate noisy max frequency moments (MF) into the synopses. This integration enables us to extend upon advanced join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join compaction sizes without losing data (§ 5.2).
4.1. Synopses generation
Steered by our key ideas, we now elaborate on our synopsis generation details. The general process is outlined as follows: Initially, the SPECIAL servers selects appropriate attributes to generate private synopses, which are then distributed to owners. The owners create the synopses using a DP mechanism and upload them to SPECIAL servers. Finally, the server applies post-processing to these synopses to aid in query processing.
Attributes selection (servers). The first step is to identify a smaller set of attributes for deriving synopses. In general, we consider the existence of a representative workload, $Q_{\mathsf{R}}$ (Kotsogiannis et al., 2019), which can be sourced from a warm-up run of SPECIAL or annotated by the administrator. Note that this process do not incur any private data thus is leakage-free. The servers first identify representative attribute pairs, $\mathsf{pair}=\{\mathsf{pair}_{k}\}_{k\geq 1}$, for each private relation $D\in\mathcal{D}$ via $Q_{\mathsf{R}}$. The designated pairs include: (i) 2-way attribute pairs, which correspond to frequently queried filter-join key combinations; (ii) frequently queried individual attributes not covered by these pairs. By default, each $\mathsf{pair}_{k}=(A_{\mathsf{ft}},A_{\mathsf{j}})$ contains two valid attributes (case i), but either $A_{\mathsf{ft}}$ or $A_{\mathsf{j}}$ may be empty (case ii).
Synopses release (owners). Next, servers pushes the identified pairs to each owner, and subsequently, the owner independently dispatches a private synopsis and returns it to servers. In what follows, we focus on the DP synopses generation mechanism evaluated by each owner, Algorithm 1 illustrates the overall flow.
Algorithm 1 DP synopsis generation $\mathcal{M}_{\mathtt{synop}}$ (in the view of $P$)
1:Input: $\mathsf{pair}=\{\mathsf{pair}_{k}\}_{k\geq 1}$ from servers; private data $D$.
2:$P$ self-determines privacy parameters $\epsilon,\delta$, and init $\mathtt{synop}\leftarrow\emptyset$
3:for each $\mathsf{pair}_{k}$ do
4:     $\mathbf{h}(\mathsf{pair}_{k},D)\leftarrow\mathsf{HistGen}(\mathsf{pair}_{k},D)$
5:DP histograms:
6:     $\mathbf{h}^{+}(\mathsf{pair}_{k},D)\leftarrow\mathbf{h}+\mathsf{Lap}^{+}(\epsilon,\delta,\mathbf{h}.\mathsf{shape})$
7:     $\mathbf{h}^{-}(\mathsf{pair}_{k},D)\leftarrow\mathbf{h}+\mathsf{Lap}^{-}(\epsilon,\delta,\mathbf{h}.\mathsf{shape})$
8$\triangleright$ adding independently sampled noise to every bin of $\mathbf{h}^{+}$, $\mathbf{h}^{-}$
9:DP max frequencies:
10:     if $A_{\mathsf{j}}\in\mathsf{pair}_{k}=\emptyset$ or $A_{\mathsf{j}}$ is unique valued then   $\mathtt{MF}_{k}=\emptyset$
11:     else if $A_{\mathsf{ft}}\in\mathsf{pair}_{k}\neq\emptyset$ then
12$\triangleright$ assuming $\mathbf{h}$ partitions $dom(A_{\mathsf{ft}})$ into $\{B_{1},...,B_{m}\}$
13:         $D^{\ell}\leftarrow\sigma_{A_{\mathsf{ft}}\in B_{\ell}}(D)$ for $\ell=1,2,...,m$
14:         compute noisy MF table, $\mathtt{MF}_{k}=\{\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\}_{1\leq\ell\leq m}$
15:     else $\mathtt{MF}_{k}=\widehat{\mathtt{mf}}(A_{\mathsf{j}},D)$
16:     $\mathtt{synop}\leftarrow\mathtt{synop}\cup(\mathsf{pair}_{k},\{{\bf h}^{+},{\bf h}^{-}\},\mathtt{MF}_{k})$
17:release $\mathtt{synop}$, $\epsilon$, $\delta$ to servers
Generally, we expect owners to set a desired privacy budget for their data (using parameters $\epsilon$ and $\delta$) before initiating the synopsis generation. For each $\mathsf{pair}_{k}$, the owner first creates a histogram $\mathbf{h}(\mathsf{pair}_{k},D)$. By default, we assume there exist global parameters, i.e. bin sizes of each attribute, which guide all owners to partition attributes consistently. Next, the owner derives two noisy histograms, $\mathbf{h}^{+}(\mathsf{pair}_{k},D)$, and $\mathbf{h}^{-}(\mathsf{pair}_{k},D)$ by adding independently sampled one-sided Laplace noises to every bin of $\mathbf{h}(\mathsf{pair}_{k},D)$.
Definition 4.1 (One-sided Laplace variable).
$\mathsf{Lap}^{+}(\epsilon,\delta)=\max(0,z)$ (resp. $\mathsf{Lap}^{-}(\epsilon,\delta)=\min(0,z)$) is a one-sided Laplace random variable in the range of $[0,\infty)$ (resp. $(-\infty,0]$) if $z$ is drawn from a distribution with the following density function
(2) $\begin{split}\textup{Pr}\left[z=x\right]=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon|x-\mu|}\end{split}$
where $\mu=1-\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))$ (resp. $\mu=\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))-1$).
The one-sided Laplace noise can be strictly positive ($\mathsf{Lap}^{+}$) or negative ($\mathsf{Lap}^{-}$). This ensures that $\mathbf{h}^{+}$ always overestimates the actual histogram, while $\mathbf{h}^{-}$ consistently underestimates it. For the ease of description, we refer $\mathbf{H}(\mathsf{pair}_{k},D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$ as the bounding histogram of attribute $\mathsf{pair}_{k}$, with $\mathbf{h}^{+}$, and $\mathbf{h}^{-}$ to be the upper and lower histograms, respectively.
We continue with the details of the (noisy) MFs generation. Assume both $A_{\mathsf{ft}},A_{\mathsf{j}}\neq\emptyset$, and the $\mathbf{h}(\mathsf{pair}_{k},D)$ partitions $dom(A_{\mathsf{ft}})$ into $\{B_{1},...,B_{m}\}$. The mechanism $\mathcal{M}_{\mathtt{synop}}$ generates a table of noisy MFs, $\{\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\}_{\ell\geq 1}$, where each entry $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})$ is an independent MF statistic for $A_{\mathsf{j}}$ over specific filtered data on $A_{\mathsf{ft}}$, such that
(3) $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\leftarrow\widehat{\max}_{\epsilon}\left(\mathcal{G}_{\mathbf{count}(A_{\mathsf{j}})}\left(\sigma_{A_{\mathsf{ft}}\in B_{\ell}}(D)\right)\right)$
here $\mathcal{G}_{\mathbf{count}(A_{\mathsf{j}})}$ is a group-by-count operation over $A_{\mathsf{j}}$, and $\widehat{\max}_{\epsilon}$ is a report noisy max mechanism (Dwork et al., 2014). It first adds i.i.d. noise from the exponential distribution $\mathsf{Exp}(\frac{2}{\epsilon})$ to each grouped count, then outputs the largest noisy count. We stress that $\mathcal{M}_{\mathtt{synop}}$ will not generate noisy MFs for non-join key attributes, and when $A_{\mathsf{ft}}$ is empty, a global MF will be generated instead of MF tables (Alg 1:10). Moreover, since SPECIAL enables owners to label attributes as unique-valued, if $A_{\mathsf{j}}$ is known to be unique-valued, then $\smash{\widehat{\mathtt{mf}}(A_{\mathsf{j}},\cdot)}$ is always 1. Nevertheless, as exponential noises are non-negative, thus $\widehat{\mathtt{mf}}\geq\mathtt{mf}$ holds for all cases. Given the discussion so far, we summarize SPECIAL synopses as:
Definition 4.2 (SPECIAL synopses).
Given representative workload $Q_{\mathsf{R}}$, we consider for each relation $D$, its corresponding synopsis $\mathtt{synop}$ to be the collection of $\{(\mathsf{pair}_{k},\mathbf{H}(\mathsf{pair}_{k},D),\mathtt{MF}_{k})\}_{k\geq 1}$, such that
  • $\mathsf{pair}_{k}$ is a frequently queried attribute (pair) over $D$ found in $Q_{\mathsf{R}}$.
  • $\mathbf{H}$ is the private bounding histogram for $\mathsf{pair}_{k}$ over $D$.
  • $\mathtt{MF}_{k}$ represents a collection of privately overestimated join key MFs that corresponds to a group of tuples in $\mathbf{h}$, categorized by $\mathsf{pair}_{k}.A_{\mathsf{ft}}$.
Theorem 4.3.
Given $|\mathsf{pair}|=c$, $\epsilon,\delta>0$, the synopsis generation (Algorithm 1) is $(\hat{\epsilon},\hat{\delta})$-DP where $\hat{\epsilon}\leq 6\epsilon\sqrt{c\ln(1/\delta)}$, and $\hat{\delta}=(c+1)\delta$
For space concern, we move the complete proof to Appendix A. In a sketch, adding $\mathsf{Lap}^{+}$ (or $\mathsf{Lap}^{-}$) to a single bin is $(\epsilon,\delta)$-DP. By parallel and sequential composition, generating $\mathbf{{H}}$ is $(2\epsilon,2\delta)$-DP. Moreover, each report noisy max is $(\epsilon,0)$-DP, and by parallel composition, the generation of the entire MF table is also $(\epsilon,0)$-DP. In this way, we know that the generation of each $(\mathsf{pair}_{k},\mathbf{H},\mathtt{MF}_{k})$ is at most $(3\epsilon,\delta)$-DP. Given there are in total $c$ such pairs, and thus the total privacy loss is subject to $c$-fold advanced composition (Dwork et al., 2014).
Synopsis transformations. We say that one can perform transformations on released synopses without incurring extra privacy loss, per the post-processing theorem of DP (Dwork et al., 2014). Now, we briefly outline key synopsis transformations relevant to the SPECIAL design. First, given any (2d) bounding histogram $\mathbf{H}(\mathsf{pair},D)$ with both $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}$ are non-empty, one can derive the (1d) bounding histograms, i.e. $\mathbf{H}(A_{\mathsf{j}},D)$ and $\mathbf{H}(A_{\mathsf{ft}},D)$, for any single attribute $A_{\mathsf{ft}}$ or $A_{\mathsf{j}}$ by marginal sums $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over $A_{\mathsf{j}}$ or $A_{\mathsf{ft}}$, respectively. This enables the creation of statistics on individual attributes, even when $A_{\mathsf{ft}}$ and $A_{\mathsf{j}}$ are not included as a standalone synopsis attribute. Moreover, it’s possible to derive relevant join key statistics even following a selection on the base relation. For example, given $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}\neq\emptyset$, and let $D^{\prime}\leftarrow\sigma_{A_{\mathsf{ft}}\in vals}(D)$, one can obtain the (1d) bounding histogram $\mathbf{H}(A_{\mathsf{j}},D^{\prime})$ by conducting a selective marginal sum of $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over bins of $A_{\mathsf{ft}}$ that intersect with $vals$. Beyond bounding histograms, join key MFs over pre-filtered data can also be computed by
(4) ${}\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\prime})=\min\left(\textstyle\sum_{B_{\ell}\cap vals\neq\emptyset}\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell}),\widehat{\mathtt{mf}}(A_{\mathsf{j}},D)\right)$
Note that $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D)$ exists if $A_{\mathsf{j}}$ is also included as a standalone synopsis attribute; otherwise, Eq 4 yields only the first term.
5. SPECIAL Primitives
We now outline key secure database operations offered by SPECIAL. A major challenge in designing synopsis-assisted secure processing is the evident knowledge gap on how private synopses could potentially accelerate oblivious query processing. To bridge this gap, we explore various uses of synopses, such as creating private indexes (SPEidx 5.2) and designing compacted oblivious operations (SPEop 5.3). Moreover, as discussed earlier, joins are the most intensive operations. Hence, we extensively optimize join algorithms, synergizing both private indexing and compaction techniques to develop a novel, parallel-friendly oblivious join (§ 5.3). Another challenge in primitive design arises from ensuring lossless guarantees. To address this, we integrate mechanisms into our designs that pessimistically estimate selection cardinalities, indexing ranges, and complex join sizes, using a combination of available synopses and advanced upper bound techniques (Hertzschuch et al., 2021). For clarity, we’ll assume, WLOG., that all input relations below are of size $n$ and all 1d histograms contain $m$ bins.
5.1. Basic Operations
SPECIAL supports conventional fully-oblivious operators (Bater et al., 2017). In general, these operators function logically the same as non-private ones, but their execution transcripts are completely data independent, with results consistently padded to the worst-case maximum.
Default data access SeqACC. By default, all query executions proceed with a standard data access, loading all data into a secure array using a sequential scan [^2] In the context of a secret-shared array $\langle D\rangle$, this step can be each server sequentially load the respective secret shares of $D$ into memory, readying them for MPC.. Additionally, each loaded tuple gets a secret bit (initially ‘0’), $\mathsf{ret}$, used to mark tuple validity.
SELECT. is the secure version of filter $\sigma_{p}(R)$. It conducts a linear scan over secure array $\langle R\rangle$ for relation $R$ and updates the $\mathsf{ret}$ bit: ‘1’ for tuples satisfying predicate $p$ and ‘0’ for others. Every tuple’s $\mathsf{ret}$ bit is updated, regardless of whether it matches the predicate.
PROJECT. discards irrelevant attributes for relation stored in $\langle R\rangle$. Note that the secret bit $\mathsf{ret}$ cannot be discarded by a PROJECT.
JOIN. implements the secure version of $\theta$-join, $R_{0}\Join_{\theta}R_{1}$. It processes two secure arrays $\langle R_{0}\rangle$ and $\langle R_{1}\rangle$, computing their cartesian product $\langle R_{0}\times R_{1}\rangle$ and then using SELECT to mark the joined tuples. It’s important to note that JOIN expands output, always padding it to the worst-case max (product of input sizes).
COUNT, SUM, MIN/MAX are the aggregation primitives. The operators linearly scan a secure array and continuously update a secret-shared aggregation value after each tuple access.
ORDER-BY, DISTINCT, GROUP-BY(AGG). are built on the oblivious sort primitive (Batcher, 1968). Specifically, ORDER-BY obliviously sorts a secure array according to a given attribute. DISTINCT sorts a secure array first, then performs a linear scan to identify unique tuples. For instance, among consecutively sorted identical tuples, it sets the $\mathsf{ret}$ bit of the last tuple to ‘1’ and the rest to ‘0’. GROUP-BY(AGG) starts by using DISTINCT to identify unique tuples. For each distinct tuple ($\mathsf{ret}$ set to ‘1’), it appends an aggregation value derived from the tuple and a dummy attribute (e.g., ‘-1’) for non-distinct tuples.
5.2. SPECIAL Index SPEidx
In conventional databases, indexes are powerful data structures that map attribute values to positions within a sorted array, enabling fast data access. However, generalizing such feature to SCA remains largely underexplored with existing solutions facing significant limitations such as high data loss (Roy Chowdhury et al., 2020), dependence on complex data structures filled with dummy tuples (Zhang et al., 2023; Bogatov et al., 2021), and limited query capabilities (i.e., linear queries across fixed attributes). Moreover, these methods only index base relations. SPEidx innovatively overcomes these issues by enabling the creation of lossless indexes directly on outsourced data or query intermediate results, without the need for additional structures or dummy injections.
In general, SPEidx builds upon the typical indexing model that utilizes cumulative frequencies (CF) (Kraska et al., 2018). Specifically, given $D$ sorted by $A\in attr(D)$, all records $t\in D$ where $t.A=x$ can be indexed by the interval $[g(x-1),g(x)]$, where $g(x)=|\{t\mid t.A\leq x\}|$ is the CF function. For better illustration, we show an example index lookup in Figure 2: to get all records with an attribute value of 24, one may compute $[g(23),g(24)]=[217,248]$ and access the relevant data from the subset $D[217:248]$. To make this indexing method private and lossless, the key idea of SPEidx is to derive two noisy CF curves from synopses (e.g. bounding histograms). One curve, $g^{+}(x)$, consistently overestimates $g(x)$, while the other, $g^{-}(x)$, consistently underestimates it. Then for any attribute value $x$, we can now derive a private interval $[g^{-}(x-1),g^{+}(x)]$ that losslessly indexes all desired records. For instance, as illustrated in Figure 2, the SPECIAL index might estimate the index range for attribute value 24 as $[g^{-}(23),g^{+}(24)]=[198,267]$.
Figure 2. True index (left) vs. SPECIAL index (right) for $x=24$
In what follows, we provide the formal explanations on how SPEidx derives indexes from DP synopses. Specifically, SPEidx first determines the bounding histograms $\mathbf{H}(A,D)$, which may be either transformed from an available 2D histogram $\mathbf{H}(\mathsf{pair},D)$ with $A\in\mathsf{pair}$, or sourced directly if $\mathbf{H}(A,D)$ is already included in the synopses. It then constructs the noisy mapping as follows:
Definition 5.1 (SPECIAL index).
Given $D$ sorted by $A$, the bounding histogram $\mathbf{H}(A,D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$, and assume $\mathbf{h}^{+}=(c^{+}_{1},...,c^{+}_{m})$, $\mathbf{h}^{-}=(c^{-}_{1},...,c^{-}_{m})$ partitions $dom(A)$ into $\{B_{1},...,B_{m}\}$. We say $\textsc{SPEidx}(A,D)=\{\mathsf{idx}_{i}=[\mathsf{lo}_{i},\mathsf{hi}_{i}]\}_{1\leq i\leq m}$ is the SPECIAL index of $D$ over $A$ with:
  • $\forall~{}i\geq 1$, $\mathsf{hi}_{i}=\min(|D|,\sum^{i}_{k=1}c^{+}_{k})$.
  • $\mathsf{lo}_{1}=0$, and $\forall~{}i\geq 2$, $\mathsf{lo}_{i}=\sum^{i-1}_{k=1}\max(0,c^{-}_{k})$.
By this construction, all tuples $t\in D$ such that $t.A\in B_{i}$ will be organized into the subset $D[\mathsf{idx}_{i}]\subseteq D$. This subset can be quickly accessed if $D$ is already sorted, without the need for special data structures or inclusion of dummy tuples. Depending on how bounding histograms are constructed, $\textsc{SPEidx}(A,D)$ can support indexing lookups with varying granularity. This can range from indexing individual attribute values (where each $B_{i}$ corresponds to a single domain value) to indexing a range of of values. The bounding histogram’s pessimistic estimation ensures that all tuples where $t.A\in B_{i}$ are accurately contained within $D[\mathsf{idx}_{i}]$, thereby achieving lossless indexing. In contrast to existing methods that are limited to indexing base relations (Roy Chowdhury et al., 2020; Bogatov et al., 2021; Zhang et al., 2023), SPEidx extends its capabilities to create private indexes on query intermediate results. For instance, consider $D^{\prime}\leftarrow\sigma_{A^{*}\in vals}(D)$ where the attribute pair $(A^{*},A)$ is included in $\mathtt{synop}$. Here, SPEidx can derive $\mathbf{H}(A,D^{\prime})$ from $\mathbf{H}(A^{*},D)$ and subsequently build indexes on $D^{\prime}$. Importantly, since index creation is a post-processing activity using available DP synopses, it incurs no additional privacy loss.
Indexed store and fast data access IdxAcc. SPEidx enables the exploration of a new storage layout for organizing outsourced data, namely indexed datastore, that facilities private data access. Specifically, by analyzing a representative workload $Q_{\mathsf{R}}$, one may identify the “hottest” attribute per base relation, sort them according to the “hottest” attribute, and then build indexes over the sorted data. This storage layout enables fast indexed access (IdxAcc) to retrieve a compact subset of data from the outsourced relations, thereby eliminating the need for a full table sequential scan (SeqAcc) potentially bypassing selection operations (§ 5.3). We emphasize that the SPECIAL design does not require replicating the outsourced datastore to accommodate multiple query types (Bogatov et al., 2021; Zhang et al., 2023). However, creating compact replicas (e.g., column replicas (Huang et al., 2020) over frequently queried attributes) can be optionally employed to enhance query processing speed. Moreover, the generation of all aforementioned objects (indexed store and column replicas) requires only three primitives: projection, oblivious sorting, and SPEidx, all achieved without extra privacy loss. In other words, this implies that one can selectively adjust these objects to align with dynamic query workloads, without incurring any privacy costs.
5.3. SPECIAL Operators SPEop
We introduce SPEop, a set of novel synopsis-assisted operators that maintain full obliviousness (e.g., operators’ execution cause no privacy loss) while enabling lossless compaction. To our knowledge, SPEop is the first primitive of its kind in any SCA system.
Oblivious compaction: OPAC. We first introduce the fundamental operation of oblivious compaction, a key component for many SPEop primitives. Given input $\langle R\rangle$, OPAC sorts it based on the secret bit $\mathsf{ret}$, moving tuples with $\mathsf{ret}=$ ‘1’ to the front. Then, OPAC retains only the first $k$ tuples from the sorted array. The compaction is lossless if $k$ is greater than or equal to the number of tuples with $\mathsf{ret}=$ ‘1’; otherwise, it is lossy.
SPECIAL selections: (OP)SELECT, (SP)SELECT, (DC)SELECT. Let $R$ to be a relation and $A\in attr(R)$, we now introduce three advanced selections that implements $\sigma_{A\in vals}(R)$.
Algorithm 2 $\mathtt{CardEst}(\sigma_{A\in vals}(R),\mathtt{synop})$
1:$\mathbf{h}=\emptyset$, $c=0$
2:if $\mathbf{H}(A,R)\in\mathtt{synop}$ then $\mathbf{h}\leftarrow\mathbf{h}^{+}\in\mathbf{H}(A,R)$
3:else if $\exists~{}\mathsf{pair}\in\mathtt{synop}$, s.t. $A\in\mathsf{pair}$ then
4:     $\mathbf{h}\leftarrow$ marginal sum $\mathbf{h}^{+}\in\mathbf{H}(A,R)$ over $\left(\mathsf{pair}\setminus A\right)$.
5:else return $cs=|R|$
6:return $cs=\min(|R|,\sum_{i=1}^{m}c_{i}\in\mathbf{h}:\left(B_{m}\cap vals\neq\emptyset\right))$
(OP)SELECT. is mainly implemented based on the oblivious compaction (OPAC) operation. Specifically, the operation first conducts a standard SELECT on the input secure array $\langle R\rangle$ to label selected tuples, followed by an OPAC to to eliminate a large portion of non-matching tuples. To determine the compaction size $cs$, (OP)SELECT examines the synopsis of $R$ and pessimistically estimates the cardinality of $\sigma_{A\in vals}(R)$ as shown in Algorithm 2. Since $cs$ never underestimates the actual cardinality, and thus, the compaction is lossless with no missing tuples. Moreover, as OPAC is fully oblivious and $cs$ is determined completely from post-processing over DP synopsis, thus, (OP)SELECT causes no privacy loss.
(SP)SELECT. The running complexity of (OP)SELECT depends on OPAC, which is typically linearithmic (see § 6.1 or (Sasy et al., 2022)). However, when $\mathtt{CardEst}(\sigma_{A\in vals}(R),\mathtt{synop})$ is relatively small, oblivious selection can be achieved without necessarily incurring linearithmic cost. Specifically, we consider (SP)SELECT, which first creates an empty output array $\langle R_{o}\rangle$ with size equals to $cs$ before any computations. Next, it evaluates two linear scans over $\langle R\rangle$, where the first scan obliviously marks all selected tuples, and in the second scan, it privately writes all marked tuples into $\langle R_{o}\rangle$. Specifically, in the second scan, (SP)SELECT internally maintains the last actual write position $\mathsf{idx}$ in $\langle R_{o}\rangle$. Then for every newly accessed tuple $\langle t\rangle$ in $\langle D\rangle$, a write action occurs on all tuples in $\langle R_{o}\rangle$. If $\langle t\rangle$ is selected, then an actual write is made that writes $\langle t\rangle$ to $\langle R_{o}[\mathsf{idx}+1]\rangle$ and a dummy write [^3] In the context of the secret-shared secure array $\langle\mathbf{a}\rangle$, a dummy write to $\langle\mathbf{a}[i]\rangle$ is simply a re-sharing of $\mathbf{a}[i]$ through secure protocols without changing its value. is made to elsewhere. If not, dummy writes are made throughout $\langle R_{o}\rangle$. Similarly, (SP)SELECT does not yield additional privacy loss, and the complexity is linear in its input.
(DC)SELECT. Finally, if the underlying data is already indexable on $A$, a direct selection over the data can be applied that bypasses secure computations. The operator simply looks up $\textsc{SPEidx}(A,R)$, and accesses $R[a,b]$, where $a=\min_{i}(\mathsf{idx}_{i}.\mathsf{lo}),b=\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})$, where $\mathsf{idx}_{i}$ dentoes the index in $\textsc{SPEidx}(A,R)$ where $B_{i}\cap vals\neq\emptyset$.
SPECIAL join: (MX)JOIN. We now introduce a novel MF-Index based oblivious join operation. The advancements of (MX)JOIN stand out in two key aspects compared to existing oblivious joins. First, compared to the standard JOIN, (MX)JOIN stands out for its ability to significantly compact the output size, coupled with a highly parallelizable fast processing mode. Second, existing oblivious joins with DP leakages typically require additional privacy budget to learn join sensitivity (Dong et al., 2022) or necessitate truncations on joined tuples (Wang et al., 2022; Bater et al., 2018). Moreover, (MX)JOIN is unique as the first oblivious join that enables lossless output compaction without extra privacy loss. We illustrate the construction details in Algorithm 3.
Algorithm 3 (MX)JOIN (base and pre-filtered relations)
1:Input: relations $R_{0}$, $R_{1}$; join attribute $A_{\mathsf{j}}$.
2:if $\mathsf{MXReady}(R_{0},R_{1})==\mathsf{True}$  then  $\mathsf{BucketJoin}(R_{0},R_{1},A_{\mathsf{j}})$
3:else if $R_{0}$, $R_{1}$ are either base or pre-filtered relation then
4:     for $b\in\{0,1\}$ do
5:         derive $\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{b})$ from $\mathtt{synop}_{b}$ (§ 4)
6:         build index $\textsc{SPEidx}(A_{\mathsf{j}},R_{b})=\{\mathsf{idx}_{i}\}_{i=1,..,m}$ (§ 5.2)
7:     if $\forall b~{}$, $\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{b})$, and $\textsc{SPEidx}(A_{\mathsf{j}},R_{b})\neq\mathsf{null}$  then
8:         oblivious sort $R_{0}$, $R_{1}$ on $A_{\mathsf{j}}$, $\mathsf{BucketJoin}(R_{0},R_{1},A_{\mathsf{j}})$
9:     else assert “not applicable for (MX)JOIN”
10:$\mathsf{BucketJoin}(R_{0},R_{1},A_{\mathsf{j}})$:
11:for $i=1,2,...,m$ do
12:     compute $O_{i}\leftarrow(R_{0}[\mathsf{idx}_{i}]\Join_{A_{\mathsf{j}}}R_{1}[\mathsf{idx}_{i}])$ via basic JOIN
13:     $cs_{i}\leftarrow\min\left(\frac{|R_{0}[\mathsf{idx}_{i}]|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})},\frac{|R_{1}[\mathsf{idx}_{i}]|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}\right)\times{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})}\cdot{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}$
14:     $R_{\mathsf{out}}\leftarrow R_{\mathsf{out}}\cup{\texttt{OPAC}}(O_{i},cs_{i})$
15:return $R_{\mathsf{out}}$
In general, (MX)JOIN can be applied to two types of data: the base and pre-filtered relations where the join key attribute is contained by the synopses. Specifically, (MX)JOIN starts with computing the join key MFs (Alg 3:4) and constructing private indexes (Alg 3:5) for both inputs. All these operations are conducted through “privacy cost-free” transformations using available DP synopses. Once these objects are obtained, the algorithm employs oblivious sort to rearrange both inputs (Alg 3:6,7), rendering them indexable with tuples logically distributed into independent buckets by join key values. Next, (MX)JOIN simply adopts standard JOIN to join tuples exclusively within the same buckets (Alg 3:10). Finally, (MX)JOIN performs per-bucket output compaction, where it first determines the MF join bound (Hertzschuch et al., 2021) for each bucket join and invokes OPAC to compact the output according to the learned size (Alg 3:11,12). As bucket-wise operations are independent, the aforementioned steps lend themselves well to parallelized processing. For better illustration, we visualize (MX)JOIN’s processing flow in Figure 3.
Figure 3. Example processing flow of (MX)JOIN.
Since (MX)JOIN derives join compaction sizes completely from post-processing of DP synopses, it thus incurs no extra privacy loss. Moreover this design choice also eliminates the need for computing join sensitivities — a major hurdle in existing — as accurate join sensitivity estimation is hard and typically requires a significant privacy budget (Dong et al., 2022). Additionally, the noisy MF bounds guarantee that compaction sizes are consistently overestimated, ensuring lossless compaction of join results.
6. SPECIAL Planner
A major limitation of existing DPSCA designs is sub-optimal execution planning. Without the ability to pre-estimate intermediate result sizes, traditional methods cannot identify the most efficient execution strategies. To address this, SPECIAL employs an innovative query planner that leverages available synopses for plan size estimation, enabling optimal execution planning.
At a high level, our planner is modeled on the Selinger-style (cost-based) optimizer (Blasgen et al., 1981). It uses a bottom-up, dynamic programming approach to enumerate all equivalent secure execution plans for a given query and estimate their costs to select the optimal one. However, challenges in the design space still remain. First, the introduction of SPECIAL primitives can significantly impact the cost modeling for oblivious operations, rendering existing models (Bater et al., 2018; Liagouris et al., 2023) ineffective. Additionally, the equivalent plan search space is extensive, requiring further strategies to simplify the search space. To address these challenges, we first conduct a comprehensive analysis of the complexities of SPECIAL primitives and develop a new cost model (§ 6.1) for SPECIAL planner. Next, we design protocol-specific heuristics (§ 6.2) tailored for the SPECIAL planner to simplify the plan search space. Note that existing features of the Selinger-style optimizer, such as plan enumerations, are not within our contributions. For detailed understanding of these features, we encourage readers to explore the original IBM System R paper (Blasgen et al., 1981).
6.1. Cost Model
Similar to the cost model proposed by Shrinkwrap (Bater et al., 2018), we view the cost of a secure execution plan cost as a combination of each operator’s I/O and secure evaluation costs. In addition, we formulate these costs as functions over their processed data sizes. Specifically, given a plan with $\ell$ operators, $\mathsf{op}_{1},...,\mathsf{op}_{\ell}$, and let $\mathbf{I}=\{{I}_{1},...,{I}_{\ell}\}$, $\mathbf{O}=\{{O}_{1},...,{O}_{\ell}\}$, to be the input and output sizes of each operators. The plan cost is formulated as:
(5) $\mathsf{Cost}=\textstyle\sum_{i=1}^{\ell}C^{\mathsf{op}_{i}}_{\mathsf{in}}(I_{i})+C_{\mathsf{eval}}^{\mathsf{op}_{i}}(I_{i})+C^{\mathsf{op}_{i}}_{\mathsf{out}}(O_{i})$
Here, $C_{\mathsf{in}}$ represents the data access cost (input I/O), primarily capturing the expenses when moving data from persistent storage to an in-memory secure array. If the accessed data has already loaded into memory, $C_{\mathsf{in}}$ becomes negligible. $C_{\mathsf{out}}$ denotes the output I/O cost, modeling the expenses when writing operator results into output arrays. $C_{\mathsf{eval}}$ accounts for the secure computing expenses for evaluating each operators, typically constituting the dominant cost. Note that, in practice, the exact formulas for $C_{\mathsf{in}},C_{\mathsf{out}}$, and $C_{\mathsf{eval}}$ can vary depending on the specific secure protocol employed (garbled circuits (Yao, 1986), secret sharing (Micali et al., 1987), etc.) as well as the particular hardware configurations in use. Nonetheless, the understanding of the asymptotic costs is adequate for comprehending the principles of our query planning and optimization strategies (Liagouris et al., 2023; Bater et al., 2018). In what follows, we provide detailed analysis on the asymptotic costs for each SPECIAL operator. Similarly, we assume that all input data sizes mentioned henceforth in this section are of size $n$, and all 1D histograms have $m$ bins. Table 1 summarizes the operator costs.
Table 1. Asymptotic costs for secure operators
Operator Input I/O ($C_{\mathsf{in}}$) Eval. ($C_{\mathsf{eval}}$) Output I/O ($C_{\mathsf{out}}$)
PROJECT $O(n)$ N/A $O(n)$
Agg. $O(n)$ $O(n)$ $O(1)$
Group & Order $O(n)$ $O(n\log^{2}n)$ $O(n)$
SELECT $O(n)$ $O(n)$ $O(n)$
(OP)SELECT $O(n)$ $O(n\log n)$ $\mathsf{hist\_bound}$
(SP)SELECT $O(n)$ $O(n)$ $O(1)$
(DC)SELECT $\mathsf{idx\_bound}$ N/A N/A
JOIN $O(n)$ $O(n^{2})$ $O(n^{2})$
(MX)JOIN $O(n)$ $O(n^{2})^{*}$ $\mathsf{mf\_bound}$
Assuming the maximum size of the indexed buckets is bounded by $O(\frac{n}{\log n})$.
Oblivious sorting and compaction. Oblivious sorting and compaction (OPAC) are crucial for many SPEop operators, so we will begin our analysis with these two primitives. While oblivious sorting algorithms with optimal $O(n\log n)$ complexity exist, they often necessitate either impractically large constants (Ajtai et al., 1983; Goodrich, 2014) or client-side memory (Asharov et al., 2020), both do not fit with SCA scenario. Consequently, we will consider the well-established bitonic sorting based implementation for oblivious sort, which come with $O(n\log^{2}n)$ complexity. Nonetheless, efficient OPAC implementations with $O(n\log n)$ complexity remain achievable (Sasy et al., 2022)
Projection, grouping and aggregation. The PROJECT accesses private relations and discards unnecessary columns independently on each server, without secure computations. Thus, I/O costs dominate this operation, with both input and output costs bounded by $O(n)$. The costs of ORDER-BY, DISTINCT, and GROUP-BY are primarily due to oblivious sorting, resulting in a complexity of $O(n\log^{2}n)$. Additionally, as these operators do not reduce output sizes, both $C_{\mathsf{in}}$ and $C_{\mathsf{out}}$ are bounded by $O(n)$. Finally, the cost of aggregations, i.e. COUNT, SUM, and MIN/MAX subjects to a oblivious linear scan, typically outputting a single secret-shared value. Hence, its $C_{\mathsf{eval}}$ is bounded by $O(n)$, with $C_{\mathsf{in}}$ at $O(n)$ and $C_{\mathsf{out}}$ at $O(1)$.
Selections. The primary cost of SELECT stems from an oblivious linear scan, making $C_{\mathsf{eval}}$ within $O(n)$. Since SELECT does not shrink the output size, both $C_{\mathsf{in}}$ and $C_{\mathsf{out}}$ are within $O(n)$. (OP)SELECT requires an oblivious compaction (OPAC) before writing outputs, where OPAC usually yields an $O(n\log n)$ complexity (Sasy et al., 2022). Consequently, its $C_{\mathsf{eval}}$ is bounded by $O(n\log n)$ with input I/O cost same as SELECT. However, as (OP)SELECT compacts output size, $C_{\mathsf{out}}$ is reduced to $\mathsf{hist\_bound}=O\left(\sum_{B_{i}\cap vals\neq\emptyset}|B_{i}|\right)$, where $B_{i}\cap vals\neq\emptyset$ are bins in the synopsis histogram intersecting with selection conditions. If $\sum_{B_{i}\cap vals\neq\emptyset}|B_{i}|\sim O(1)$, (SP)SELECT becomes preferable, with its running cost dominated by a two-phase linear scan, and thus $C_{\mathsf{eval}}$ is now $O(n)$, and the output cost is $O(1)$. (DC)SELECT is the most efficient selection, though it requires indexable input data. It simply loads indexed data into secure arrays without secure computations or result write-backs. As such, both $C_{\mathsf{evals}}$ and $C_{\mathsf{out}}$ are negligible, with $C_{\mathsf{in}}$ within $\mathsf{idx\_bound}=O\left(\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})-\min_{i}(\mathsf{idx}_{i}.\mathsf{lo})\right)$. Here, $\mathsf{idx}_{i}$ are indexed regions that intersect with selection conditions.
Joins. Both JOIN and (MX)JOIN have $O(n)$ data access costs, but differ in $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$. JOIN, using a Cartesian product approach, doesn’t reduce output size, so both $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$ are within $O(n^{2})$. Compared to JOIN, in the worst-case scenario where the join keys follow a highly biased distribution, i.e. max bucket size reaches $O(n)$, (MX)JOIN’s asymptotic cost is at most $O(n^{2}\log n)$. However, when join keys are distributed more uniformly, the cost can be asymptotically better. For instance, with $m=\log n$ and assuming a max bucket size of $O(\frac{n}{\log n})$, each bucket join costs $O(\frac{n^{2}}{\log n})$, leading to a total cost of $O(n^{2})$, equivalent to JOIN. Recall that bucket joins in (MX)JOIN can be executed concurrently, hence, the processing latency is indeed dominated by the bucket-wise cost, i.e. $O(\frac{n^{2}}{\log n})$. Additionally, the output cost is lowered from $O(n^{2})$ to the sum of per-bucket MF upper bounds (Alg 3:11), which can be substantially less if the join key MFs are relatively low.
6.2. Heuristics
H-1. Filter push down. Filter pushdown is a widely used optimization technique in conventional databases, where it involves moving the selection operation to the earliest stage in the data processing pipeline, such as directly to the base relation. This reduces the amount of data that needs to be loaded, transferred, and processed, by following operators especially complex joins. However, in SCA designs, data obliviousness typically necessitates padding selection sizes to the worst-case scenario, rendering filter pushdown ineffective (Liagouris et al., 2023; Eskandarian and Zaharia, 2017; Zheng et al., 2017; Poddar et al., 2021). With SPEop selections, the selection outputs can be significantly compacted, making filter pushdown effective again. Therefore, we include filter push down as one of the optimization heuristic for SPEplan.
H-2. Predicates fusion. Let $R$ to be any relation, $A_{1},A_{2},...,A_{k}\subseteq attr(D)$, and $\mathbf{v}=\{v_{1},v_{2},...,v_{k}\}$. We say that for multiple selection over $R$ such that $\sigma_{A_{1}\in v_{1}}(...\sigma_{A_{k}\in v_{k}}(R))$, one can always fuse them into one selection $\sigma_{\mathbf{A}\subset\mathbf{v}}(R)$. This can reduce the number of secure computation invocations from $k$ rounds to just one. Additionally, the selection size can be estimated as $\min_{A_{i}}\left(\mathtt{CardEst}(\sigma_{A_{i}\in v_{i}}(R))\right)$.
H-3. Join statistics propagation. A key property of SPECIAL join, (MX)JOIN, is that the output is already indexed and bucketized by the join key. Therefore, for any output $R$ of (MX)JOIN computing $R_{0}\Join_{A_{\mathsf{j}}}R_{1}$, a new index $\textsc{SPEidx}(A_{\mathsf{j}},R)$ across $R$ can be easily derived. Moreover, as per (Hertzschuch et al., 2021), one can also update the MF for $R$ by computing $\widehat{\mathtt{mf}}(A_{\mathsf{j}},R)=\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})\times\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})$. As a result, we say that the output of (MX)JOIN as MF-and-index-ready, enabling direct application of another (MX)JOIN on the same join attribute.
7. Evaluation
In this section, we present evaluation results of our proposed framework. Specifically, we address the following questions:
  • Question-1: Does SPECIAL offer efficiency advantages over current SCA approaches? To what extent can SPECIAL enhance efficiency compared to existing state-of-the-arts (SOTAs)?
  • Question-2: For the DP-based SPECIAL design, is there a trade-off between privacy and efficiency? Can we adjust privacy levels to achieve various efficiency objectives?
  • Question-3: Can SPECIAL scale complex analytical (e.g. multi-way join) queries to large-scale (multi-million rows) datasets?
7.1. Experimental setups
We briefly outline our experimental setup, covering the baseline systems, prototype SPECIAL implementation, workloads, datasets, and default configurations.
Baseline systems and SPECIAL prototype. We compare SPECIAL with two baseline systems: Shrinkwrap (Bater et al., 2018), the SOTA DPSCA, and SMCQL (Bater et al., 2017), a typical SCA system using exhaustive padding for leakage hiding, also used as a baseline for Shrinkwrap. For consistency, we consider the same circuit-model implementations for both baseline systems and the SPECIAL prototype. While some works (Poddar et al., 2021; Liagouris et al., 2023) similar to SMCQL use exhaustive padding but improve efficiency through protocol-level optimizations, we exclude them from our benchmarks for fair comparison concerns. However, it’s important to note that these optimizations are orthogonal to both the baseline systems and our designs. We re-implement important query features for the two baseline systems, and built SPECIAL using the same MPC package, specifically the sh2pc-0.2.2 protocol within EMPtoolkit-0.2.5. Since managing DP synopses and generating indexes do not involve secure computations, thus we implement the related features using Python (3.7) scripts. All implementations are open-sourced (Lovingmage, 2024).
Datasets and workloads. Although SMCQL and Shrinkwrap use the HealthLNK dataset, it is not public data and thus not accessible to the broader research community. Hence, we developed a new benchmark based on a publicly available dataset, and mimic Shrinkwrap’s workload to create a set of testing queries. Specifically, we use the Czech Financial Dataset (Rel, [n.d.]), an anonymized Czech bank transaction dataset for loan applications. This dataset comprises 8 relational tables with a total of 55 columns and 1.1 million rows. For scaling experiments, we use the raw data and schema to generate synthetic data with up to 10 million rows. Shrinkwrap’s benchmark evaluates four queries: a range query, two 2 binary joins, and a 3-way join. Mirroring this approach, we designed eight query workloads based on the Financial data. All workloads are selection-projection-join (SPJ) queries followed by statistical aggregations such as count, count distinct, and group by. The workloads range from simple linear queries to complex multi-way joins. Due to space limitations, a brief summary of the workloads is provided in Table 2, with the full specification in Appendix B.
Table 2. Query workloads
Query No. Type Description
Q1 Linear Point query with single predicate.
Q2 Linear Range query with conjunctive predicates.
Q3 Binary Join Non-expanding binary join.
Q4 Binary Join Expanding binary join.
Q5 Multi Join 3-way mixed non-expanding and expanding joins.
Q6 Multi Join 3-way all expanding joins.
Q7 Multi Join 4-way mixed non-expanding and expanding joins.
Q8 Multi Join 5-way mixed non-expanding and expanding joins.
Default configurations. For SPECIAL, each relational table is assigned a fixed privacy budget with $\epsilon=1.5$ and $\delta=0.00005$ to release synopses. This setup ensures that the total privacy loss per table does not exceed $\epsilon=1.5$. For all equal-width histograms generated in SPECIAL, we configure them to have at most 8 bins. For Shrinkwrap, we use a privacy budget of $\epsilon=1.5$ and $\delta=0.00005$ for each query processing, as specified in (Bater et al., 2018). However, Shrinkwrap does not provide bounded privacy loss guarantees, so the privacy guarantee of Shrinkwrap will be no better than SPECIAL. We conduct all experiments on bare-metal Mac machines with M2 Max CPUs and 96GB unified memory. In addition, we consider Q2, Q4, Q8 are default testing queries in each query type, and are used as the representative workload to guide synopses. Moreover, for baseline systems, as they do not have join ordering optimizations, thus we will assume a random join order for them.
Figure 4. In-depth comparisons in execution plans: (i) The exhaustive padding in SMCQL can lead to significant memory blowup; (ii) Both SMCQL and Shrinkwrap suffer from suboptimal join ordering; (iii) Although Shrinkwrap reduces intermediate sizes, it still requires substantial memory to materialize the Cartesian-style standard oblivious join; (iv) SPECIAL effectively identifies optimal join orders and significantly reduces intermediate materialization sizes; (v) The IdxAcc method in SPECIAL not only bypasses secure selections but also directly reduces input I/O costs.
7.2. End-to-end performance comparisons
To address Question-1, we first conduct an end-to-end performance comparison of SPECIAL, Shrinkwrap, and SMCQL across the full benchmark workloads (Q1-Q8). We use two typical performance metrics, query execution time and memory usage in our evaluations. The results are summarized in Figure 5 and Figure 6. Due to significant memory consumption, SMCQL cannot complete the full benchmark and stops at Q4. Therefore, we only include full statistics up to Q3, and for Q4, we provide a projected performance evaluation using 10% of the input data.
Figure 5. End-to-end comparison: query latency
Figure 6. End-to-end comparison: memory usage
Observation 1. SPECIAL significantly outperforms all baseline systems in query latency, achieving performance gains of up to 3618.3×3618.3$\times$3618.3 × for linear queries, 114×114$\times$114 × for binary joins, and 80.3×80.3$\times$80.3 × for multi-joins. Figure 5 shows significant performance gains of SPECIAL in contrast to Shrinkwrap and SMCQL. For linear queries, SPECIAL exhibits substantial performance gains, achieving up to 3618.3×3618.3$\times$3618.3 × improvement (Q2), primarily due to its index-based fast data access. This approach not only reduces large I/O costs but also eliminates the need for secure computation by enabling direct data retrieval via DP indexes. We observe that in binary joins, SPECIAL has a less pronounced advantage over Shrinkwrap. This is because Shrinkwrap do not have sub-optimal planning issue when dealing with binary join—as there is only one join order. Hence, the performance gains for SPECIAL in Q3 and Q4 stem solely from fast data access and the efficient (MX)Join method. However, for more complex multi-way joins, SPECIAL’s advantage becomes more pronounced again due to its ability to select the optimal join order. For instance, Shrinkwrap shows more than 80×80$\times$80 × slow down in query latency than SPECIAL in group Q7.
Observation 2. SPECIAL shows profound improvement in memory usage (Figure 6) against baseline systems, especially in complex multi-way joins. This is primarily due to two factors: First, the (MX)JOIN used by SPECIAL is more memory-efficient compared to the joins implemented by Shrinkwrap and SMCQL. Second, SPECIAL’s capability to identify optimal execution plans significantly reduces total intermediate sizes, which is particularly beneficial for complex joins that suffer from sub-optimal or exhaustive padding in other systems. To better understand the substantial improvements SPECIAL achieves—for instance, up to 928.2×928.2$\times$928.2 × over Shrinkwrap and more than $10^{5}\times$ over SMCQL—we will zoom into a specific query, Q6, and compare the detailed execution plans of the three systems. The choice of Q6 is strategic because its complexity sufficiently highlights the differences in execution plans, yet it remains simple enough for clear visual representation. Our comparison features four execution plans: a plaintext optimal plan derived directly from the logical plan using actual cardinalities, illustrating the ground truth optimal execution; a hypothetical SMCQL plan (since Q6 cannot be completed by SMCQL so we project the cardinalities); and two actual execution plans from our experiments with Shrinkwrap and SPECIAL. The detailed comparisons and observations are summarized in Figure 4.
7.3. Privacy comparisons
Continue to address Question-1, we now compare the privacy guarantees between Shrinkwrap and SPECIAL, with the results detailed in Figure 7. As privacy loss is independent of query types, our analysis will focus on comparing the privacy composition curves under continual query answering for both Shrinkwrap and SPECIAL. We consider two privacy composition models: advanced composition (Adv.)(Dwork et al., 2014) and concentrated DP composition (CDP)(Bun and Steinke, 2016).
Figure 7. End-to-end comparison privacy loss.
Observation 3. Under continual query answering, SPECIAL demonstrates significantly lower privacy loss compared to Shrinkwrap, achieving up to 89.01×89.01$\times$89.01 × and 38.91×38.91$\times$38.91 × improvements in the Adv. and CDP modes, respectively. The privacy loss of SPECIAL is bounded to the initial synopsis release stage, so continual query answering does not incur additional privacy loss. In contrast, Shrinkwrap’s privacy loss accumulates over time as each new query allocates a fresh privacy budget. Consequently, its privacy loss exhibits a logarithmic growth, as shown in Figure 7. This accumulation can result in significant privacy degradation when processing a large number of queries. For example, answering 100 queries in Shrinkwrap could result in a privacy loss of $\epsilon>100$ in Adv. and $\epsilon\approx 60$ in CDP., respecitvely, even if each query only uses a small privacy budget of $\epsilon=1.5$. As such, SPECIAL demonstrate significant improve in privacy guarantees towards SOTA DPSCA.
7.4. Privacy efficiency tradeoffs
We address Question-2 by evaluating SPECIAL at various privacy levels and observing the performance impacts. Specifically, we maintain $\delta$ constant while varying $\epsilon$ from $0.1$ to $10$ and assess the performance across default testing queries (Q2, Q4, Q8). The results are shown in Figure 8 and Figure 9.
Figure 8. Privacy vs. Performance (query time)
Figure 9. Privacy vs. Performance (memory)
Observation 4. The privacy-efficiency tradeoff generally exists but exhibits varying trends at different privacy levels. For instance, SPECIAL shows a clear tradeoff at higher privacy levels ($\epsilon<1$), while at lower privacy levels ($\epsilon>1$), the tradeoff becomes less pronounced. When $\epsilon$ increases from $0.1$ to $1$, both memory usage and query latency for all test queries significantly decrease. However, increasing $\epsilon$ from $1$ to $10$ shows no significant performance gains. This may indicate that once $\epsilon$ exceeds $1$, the impact of noises on cardinality estimation or index building is alredy minimal, and further reductions in $\epsilon$ do not lead to notable improvements. Therefore, if high privacy protection is required, practitioners should carefully fine-tune privacy parameters to optimize performance. Conversely, if performance is the priority, setting $\epsilon$ near 1 is typically sufficient.
Observation 5. Different queries exhibit varying performance sensitivities to changes in privacy levels. Multi-join queries are significantly affected when privacy levels change. For instance, the performance overhead of Q8 at $\epsilon=0.1$ is more than 3×3$\times$3 × higher than at $\epsilon=1$. In contrast, linear query (Q2) and binary join (Q4) overhead increases to only about 1.5×1.5$\times$1.5 × under the same privacy settings. This difference stems from the error amplification inherent in multi-join operations, which frequently reuse synopses and require multiple cardinality estimations. Therefore, practitioners dealing with complex multi-join workloads may need careful fine-tune privacy parameters, as performance impacts can be substantial.
7.5. Scaling experiments
To address Question-3, we stress SPECIAL with scaled workloads by duplicating the raw financial dataset to sizes of 2×2$\times$2 ×, 4×4$\times$4 ×, and 8×8$\times$8 ×. We then evaluate the default testing queries on these scaled datasets to assess SPECIAL’s efficiency. The results are shown in Figure 10.
Figure 10. End-to-end comparison privacy loss.
Observation 6. SPECIAL demonstrates significant potential to scale up to multi-million records, even when handling complex 5-way joins. Figure 10 shows SPECIAL’s effective scaling: up to 8×8$\times$8 × data for linear queries and binary joins, and up to 4×4$\times$4 × data for complex 5-way joins like Q8. For instance, Q2 can be completed within 290ms under 8×8$\times$8 ×, and in fact, since selection is bypassed (due to index access), thus the cost is mainly on I/O costs. Q4 finishes in 289 minutes at the same scale 8×8$\times$8 ×, while the more complex 5-way join Q8 takes less than 280 minutes for 4×4$\times$4 × data. As a reference, Shrinkwrap would require over 1035 minutes to complete Q8 even with unscaled data (Figure 5). This result can, from another angle, showcase the significant performance improvements of SPECIAL over current SOTA DPSCA designs.
8. Discussion
In this section, we discuss potential extensions to accommodate broader data models and possible adaptations of SPECIAL designs to other secure primitives.
Supporting growing data. In general SPECIAL considers read-intensive analytical workloads and and a static data model, which is the same as current SOTA SCA efforts (Poddar et al., 2021; Bater et al., 2018; Liagouris et al., 2023; Volgushev et al., 2019). In this section, we will briefly discuss how SPECIAL could be extended to support the growing data model where new records are continuously added to the logical database (Wang et al., 2022, 2021; Zhang et al., 2023). One possible strategy involves segmenting the large logical database into discrete, non-overlapping time intervals, each represented by its own database with independent synopses. Specifically, for new incoming data, SPECIAL will initially cache it until a designated time interval concludes. After this period, SPECIAL will batch all the cached data into a new database corresponding to the past time interval and release synopses for this database. As the data within these intervals is disjoint, queries can be easily partitioned and independently optimized and executed across corresponding databases in parallel, before being aggregated. Note that prior to releasing new synopses, all cached data will be processed using standard SCA processing (e.g. SMCQL). Since each synopses release covers only newly inserted data, the total privacy loss adheres to parallel composition (Dwork et al., 2014), and thus the bounded privacy loss still hold.
Adaptability to other secure primitives. Beyond MPC-based SCA, the key design insights from SPECIAL are transferable to other secure outsourced databases such as the TEE-based analytical systems (Qiu et al., [n.d.]; Qin et al., 2022; Eskandarian and Zaharia, 2017; Zheng et al., 2017) or searchable encryption based private key-value store (Amjad et al., 2021; Groce et al., 2019; Chen et al., 2018). This is because SPECIAL’s core components, including synopses management, indexing, query planning, and compaction estimation, can all be executed without relying on secure computation primitives. For example, in TEE databases, one may conduct a similar query planning outside the TEE, with the optimal plan then executed within the TEE for fast oblivious processing. Similarly, SPECIAL’s design principles could be applied to searchable encryption to construct private indexes in advance on sorted data, enable fast and private information retrieval.
9. Related Work
SCA systems. Two main approaches exist for designing MPC-based SCA systems. The first is peer-to-peer paradigm (Aggarwal et al., 2005; Mouris et al., 2024; Poddar et al., 2021; Volgushev et al., 2019; Wang and Yi, 2021), where the goal is to improve efficiency by decomposing analytical queries and pushing them to data owners, so that they can either directly process in clear or running MPCs across a smaller group of parties. Unfortunately, this approach burdens data owners with heavy computational overheads, particularly for complex operations like joins. Moreover, given real-world data owners often lack robust computing resources and reliable service capabilities, these methods are difficult to scale and promise consistent SCA services to external analysts. On the other hand, the server-aided-MPC model (Wang et al., 2021; Bater et al., 2018, 2020; Wang et al., 2022; Zhang et al., 2023; Roy Chowdhury et al., 2020; Kamara et al., 2011; Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023; Mohassel and Zhang, 2017) offers a more practical solution. It allows data owners to outsource both expensive MPC computations and secure data storage to a set of capable cloud service providers (CSPs), who then jointly evaluate MPC to provide reliable SCA services. For the SPECIAL prototype, we focus on the server-aided-MPC model under a strong corruption assumption, allowing up to $n-1$ of $n$ servers to be compromised and collate with others. While, as we discussed before, SPECIAL’s core design principles are protocol-agnostic. This enables interoperability with various MPC models, including the peer-to-peer settings or a weaker corruption setting (for performance purpose) where a supermajority of servers need to be honest (Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023).
DP leakages. Leakage-abuse attacks (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016), exploit data-dependent processing patterns, are persistent threats to SCA systems. To mitigate these risks, oblivious computation techniques (Keller and Scholl, 2014; Pinkas and Reinman, 2010; Stefanov et al., 2011; Chang et al., 2022; Mishra et al., 2018; Crooks et al., 2018; Jönsson et al., 2011; Sasy et al., 2022; Asharov et al., 2020; Wang et al., 2014; Pettai and Laud, 2015; Wang et al., 2023; Tinoco et al., 2023) have become the de facto solution—processing data such that the entire execution transcript (e.g. memory traces and read/write volumes) is padded to a worst-case constant. While these techniques ensure the strongest privacy guarantees by eliminating data-dependent leakages, they also introduce a fundamental contention with modern database optimizations, which often rely heavily on data-dependent operations (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022). To this end, many recent efforts seek a practical balance in the privacy-performance trade-off by allowing controlled leakage under DP (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Groce et al., 2019; He et al., 2017; Zhang et al., 2023; Qin et al., 2022; Chan et al., 2022; Wagh et al., 2016; Chu et al., 2021; Qiu et al., [n.d.]; Patel et al., 2019). However, a common issue of these approaches is unbounded privacy loss. While some works propose to address this (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., [n.d.]), their approaches are restricted to only simple linear queries. Furthermore, existing DP leakage designs also lack lossless guarantees for complex query processing. The inherent randomness of DP noises (Wang et al., 2021; Groce et al., 2019; Wang et al., 2022) and uncertainty in complex join sensitivities (Chu et al., 2021; Bater et al., 2018; Wang et al., 2022) often lead to improper output compaction and lose tuples. SPECIAL addresses all these limitations together, and to our knowledge, is the first system to ensure both bounded privacy and lossless results for securely processing complex queries.
SCA query planning. Query planning (Silberschatz et al., 2011) is crucial in conventional databases. A few studies (Bater et al., 2017; Volgushev et al., 2019; Poddar et al., 2021; Liagouris et al., 2023) explored query planning within SCA frameworks. Unlike conventional planners, which exploit size disparities across different execution plans to choose the most efficient one with minimized plan sizes (Elmasri et al., 2020; Harmouch and Naumann, 2017; Han et al., 2021), these methods typically prohibit the use of such data-dependent information. Consequently, they primarily rely on data-independent metrics for planning. For example, (Bater et al., 2017; Liagouris et al., 2023; Volgushev et al., 2019) optimize based on the MPC execution costs, while (Poddar et al., 2021) focuses on decomposing large MPC tasks into smaller 2/3PC ones. Since these methods do not optimize or compact intermediate sizes, significant query execution overhead from these large sizes remains evident, especially when processing complex multi-joins (Bater et al., 2018; Liagouris et al., 2023). Shrinkwrap (Bater et al., 2018) uses DP leakage models and allows planner to optimize query intermediate sizes in a privacy-preserving manner. However, it cannot estimate cardinalities before runtime, mandating a random selection of plan structures (e.g. join order) before execution. Consequently, planning focuses solely on optimally allocating privacy budgets across the chosen plan, which often result in suboptimal query executions. In contrast, SPECIAL offers an advanced query planner capable of pre-estimating plan sizes and comparing costs among different plan structures before execution.
10. Conclusion
We introduce SPECIAL, the first SCA system that simultaneously supports: (i) handling complex queries with bounded privacy loss; (ii) advanced query planning that effectively exploit plan intermediate sizes before runtime; and (iii) delivering exact query results without missing tuples. This is achieved through a novel synopses-assisted SCA design, where a set of private table statistics are released with one-time privacy cost to guide subsequent secure SCA planning and processing. The benchmark results demonstrate that SPECIAL significantly outperforms SOTA SCA systems and scales complex multi-join queries to multi-million datasets
Appendix A Privacy Proof
Theorem A.1.
The one-sided Laplace mechanism as described in Definition 4.1 satisfies $(\epsilon,\delta)$-DP.
Proof.
WLOG. we illustrate proof details by assuming positive mechanism, $\mathsf{Lap}^{+}$, and by symmetric, the properties proved here also holds for $\mathsf{Lap}^{-}$. In general, we write the mechanism as $\mathcal{M}(D)\leftarrow f(D)+\mathsf{Lap}^{+}(\epsilon,\delta)$, where $f$ is a counting query. Note that the sensitive for counting query is 1, and thus we will implicitly omit the sensitivity notations.
Given any two neighboring databases $D$ and $D^{\prime}$, and let $c\leftarrow f(D)$ and $c^{\prime}\leftarrow f(D^{\prime})$, WLOG. we will assume $c\leq c^{\prime}$. Now we will evaluate three disjoint output domains for $f$, namely $O_{1}=(-\infty,c)$, $O_{2}=[c,c^{\prime}]$, and $O_{3}=(c^{\prime},+\infty)$. Since the noise is non-negative, thus it’s clear that $\textup{Pr}[\mathcal{M}(D)\in O_{1}]=0$ and $\textup{Pr}[\mathcal{M}(D^{\prime})\in O_{1}]=0$. Next, we evaluate the domain $O_{3}$ and consider an auxiliary domain $O^{*}=\{x-c^{\prime}:x\in O_{3}\}$. We compute the following ratio:
(6) $\begin{split}\frac{\textup{Pr}[\mathcal{M}(D)\in O_{3}]}{\textup{Pr}[\mathcal{M}(D^{\prime})\in O_{3}]}&=\frac{\textup{Pr}[c+\mathsf{Lap}^{+}(\epsilon,\delta)\in O_{3}]}{\textup{Pr}[c^{\prime}+\mathsf{Lap}^{+}(\epsilon,\delta)\in O_{3}]}\\ &=\frac{\textup{Pr}[\mathsf{Lap}^{+}(\epsilon,\delta)\in O^{*}]}{\textup{Pr}[\mathsf{Lap}^{+}(\epsilon,\delta)\in O^{*}+(c^{\prime}-c)]}\\ &=\frac{e^{-\epsilon|x-\mu|}}{e^{-\epsilon|x-(c^{\prime}-c)-\mu|}}\\ &=e^{\epsilon(|x-(c^{\prime}-c)-\mu|-|x-\mu|)}\\ &\leq e^{\epsilon(|c^{\prime}-c)|)}~{}~{}(\textup{triangle~{}inequality})\\ &=e^{\epsilon}\end{split}$
Then we evaluate the last output domain $O_{2}$, and we can see that $\textup{Pr}[\mathcal{M}(D^{\prime})\in O_{2}]=0$ and $\textup{Pr}[\mathcal{M}(D)\in O_{2}]\geq 0$, and thus there will be unbounded privacy loss within the domain of $O_{2}=[c^{\prime},c]$. However, we can bound this probability by $\delta$, which is equivalent to prove that with density function
$\textup{Pr}[z=x]=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon|x-\mu|},$
where $\mu=1-\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))$, the probability of sampling a value $z$ within the range $(-\infty,1]$ is less than or equal to $\delta$. Since we consider a discrete distribution, thus we compute the following
(7) $\begin{split}\Pr[z<1]&=\sum_{x=-\infty}^{0}\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon(\mu-x)}\\ &=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon\mu}\sum_{x=-\infty}^{0}e^{\epsilon x}\\ &=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon\mu}\frac{1}{1-e^{-\epsilon}}\quad\text{(geometric series sum of }e^{\epsilon x})\\ &=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon\left(1-\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))\right)}\frac{1}{1-e^{-\epsilon}}\quad\text{(substitute }\mu)\\ &=\frac{e^{\epsilon}}{e^{\epsilon}+1}e^{-\epsilon+\ln(\delta(e^{\epsilon}+1))}\\ &=\frac{e^{\epsilon}}{e^{\epsilon}+1}e^{-\epsilon}\delta(e^{\epsilon}+1)\\ &=\delta.\end{split}$
Combining all these we obtain the following
(8) $\begin{split}\textup{Pr}[\mathcal{M}(D)\in O]&=\textup{Pr}[\mathcal{M}(D)\in(O_{1}\cup O_{2}\cup O_{3})]\\ &=\textup{Pr}[\mathcal{M}(D)\in O_{2}]+\textup{Pr}[\mathcal{M}(D)\in O_{3}]\\ &\leq\delta+e^{\epsilon}\textup{Pr}[\mathcal{M}(D^{\prime})\in(O_{2})]\end{split}$
The claim thus holds. ∎
Next we prove Theorem 4.3.
Proof.
(Theorem 4.3) We prove this theorem by analyzing privacy loss using composition theorems. Generating noisy histograms with one-sided noise is equivalent to releasing multiple counting queries across disjoint datasets. Therefore, by the parallel composition theorem (Dwork et al., 2014), each histogram release is $(\epsilon,\delta)$-DP. The release of the maximum frequency via a noisy max mechanism is $(\epsilon,0)$-DP. Given $c$ pairs, the sequential composition theorem (Dwork et al., 2014) applies, resulting in a total privacy loss under $2c$-fold advanced composition. Consequently, the entire process adheres to $(\hat{\epsilon},\hat{\delta})$-DP, with $\hat{\epsilon}\leq 6\epsilon\sqrt{c\ln(1/\delta)}$.
Theorem A.2.
The privacy of SPECIAL adheres to Definition 3.1
Proof.
Since the entire DP synopses releasing satisfies DP. Thus, we prove this theorem by constructing a simulator, $\mathcal{S}$, which takes only the DP synopses as input and can simulate execution transcripts that are indistinguishable from those of the actual protocol. For ease of notation, we abstract SPECIAL as a composed secure protocol $\pi$.
We first examine a hybrid protocol, $\pi_{1}$, differing from $\pi$ in that the secure outsourcing procedure—where owners upload secret shares of their data to two servers—is replaced by a simulator, $\mathcal{S}_{1}$. This simulator, which takes no input, simulates secret shares with randomly sampled data $(x_{1},x_{2})\xleftarrow[]{\mathsf{rd}}\mathbb{Z}_{2^{32}}$. We say that a $\mathcal{S}_{1}$ must exist that produces results indistinguishable from the secret shares in $\pi$, as the absence of such a simulator would imply the non-existence of secure secret sharing techniques. Next, we consider another hybrid protocol, $\pi_{2}$, which differs from $\pi_{1}$ only in that the query planning function is replaced by a simulator $\mathcal{S}_{2}$. This simulator takes as inputs all synopses and a set of random values (simulated secure outsourced data) generated by $\mathcal{S}_{1}$, and produces a simulated execution plan for every requested query. We say that an $\mathcal{S}_{2}$ should also exist with outputs indistinguishable from $\pi$’s execution plans. Given that Selinger optimizers can pre-generate execution plans using only table statistics, the non-existence of such a simulator, $\mathcal{S}_{2}$, would imply the non-existence of Selinger optimizers, which is contradictory. Finally, we construct $\mathcal{S}=(\mathcal{S}_{1},\mathcal{S}_{2})$, where $\mathcal{S}_{1}$ simulates secure data outsourcing and $\mathcal{S}_{2}$ simulates query planning. $\mathcal{S}$ executes the query plan using randomly generated values from $\mathcal{S}_{1}$ through MPC computation. Since both the execution plans and input data are indistinguishable from those in $\pi$, the MPC execution must also be indistinguishable. Otherwise, it would violate the security guarantees of MPC. ∎
Appendix B Query Workloads
We illustrate the benchmark query workloads as follows:
Listing 1: Benchmark query workload

### v2

@@
 Abstract.
-Secure collaborative analytics (SCA) enable the processing of analytical SQL queries across multiple owners’ data, even when direct data sharing is not feasible.
-Although essential for strong privacy, the large overhead from data-oblivious primitives in traditional SCA has hindered its practical adoption.
-Recent SCA variants that permit controlled leakages under differential privacy (DP) show a better balance between privacy and efficiency.
-However, they still face significant challenges, such as potentially unbounded privacy loss, suboptimal query planning, and lossy processing.
+Secure collaborative analytics (SCA) enables the processing of analytical SQL queries across data from multiple owners, even when direct data sharing is not possible.
+While traditional SCA provides strong privacy through data-oblivious methods, the significant overhead has limited its practical use.
+Recent SCA variants that allow controlled leakages under differential privacy (DP) strike balance between privacy and efficiency but still face challenges like unbounded privacy loss, costly execution plan, and lossy processing.
 To address these challenges, we introduce SPECIAL, the first SCA system that simultaneously ensures bounded privacy loss, advanced query planning, and lossless processing.
-SPECIAL employs a novel synopsis-assisted secure processing model, where a one-time privacy cost is spent to acquire private synopses (table statistics) from owner data.
-These synopses then allow SPECIAL to estimate (compaction) sizes for secure operations (e.g., filter, join) and index encrypted data without extra privacy loss.
-Crucially, these estimates and indexes can be prepared before runtime, thereby facilitating efficient query planning and accurate cost estimations.
-Moreover, by using one-sided noise mechanisms and private upper bound techniques, SPECIAL ensures strict lossless processing for complex queries (e.g., multi-join).
-Through a comprehensive benchmark, we show that SPECIAL significantly outperforms cutting-edge SCAs, with up to 80×80$\times$80 × faster query times and over 900×900$\times$900 × smaller memory for complex queries.
-Moreover, it also achieves up to an 89×89$\times$89 × reduction in privacy loss under continual processing.
+SPECIAL employs a novel synopsis-assisted secure processing model, where a one-time privacy cost is used to generate private synopses from owner data.
+These synopses enable SPECIAL to estimate compaction sizes for secure operations (e.g., filter, join) and index encrypted data without additional privacy loss.
+These estimates and indexes can be prepared before runtime, enabling efficient query planning and accurate cost estimations.
+By leveraging one-sided noise mechanisms and private upper bound techniques, SPECIAL guarantees lossless processing for complex queries (e.g., multi-join).
+Our comprehensive benchmarks demonstrate that SPECIAL outperforms state-of-the-art SCAs, with up to $80\times$ faster query times, $900\times$ smaller memory usage for complex queries, and up to $89\times$ reduced privacy loss in continual processing.
+PVLDB Reference Format:
+PVLDB, 18(4): 1035 - 1048, 2024.
+doi:10.14778/3717755.3717764 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
+For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
+Copyright is held by the owner/author(s).
+Publication rights licensed to the VLDB Endowment.
+Proceedings of the VLDB Endowment, Vol. 18, No. 4 ISSN 2150-8097.
+doi:10.14778/3717755.3717764
+PVLDB Artifact Availability:
+The source code, data, and/or other artifacts have been made available at https://github.com/lovingmage/caplan.
 1. Introduction
@@
 Consequently, data-dependent processing patterns such as memory traces and read/write volumes can still reveal critical information, risking privacy breaches (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016) even when the core data remains encrypted.
-To ensure strong privacy, modern SCA systems also utilize data-oblivious primitives that exhaustively pad query processing complexities to a worst-case, data-independent upper bound (Bater et al., 2017; Liagouris et al., 2023; Poddar et al., 2021).
+To ensure strong privacy, modern SCA systems utilize data-oblivious primitives that exhaustively pad query processing complexities to a worst-case and data-independent upper bound (Bater et al., 2017; Liagouris et al., 2023; Poddar et al., 2021).
 However, such stringent protections can largely reduce system efficiency and hinder the generalization of conventional optimization techniques to SCA, which are typically data-dependent (Liagouris et al., 2023).
 To address this, recent efforts (Qin et al., 2022; Wang et al., 2021, 2022; Bater et al., 2018) have introduced Differentially Private SCA (DPSCA).
-This approach allows controlled information leakage under differential privacy (Dwork et al., 2014) to mitigate constant worst-case overhead.
-For instance, systems under this model can dynamically compact intermediate query sizes to a noisy estimate close to actual sizes, avoiding exhaustive padding.
+This approach allows controlled information leakage under DP (Dwork et al., 2014) to mitigate constant worst-case overhead.
+For instance, systems under this model can dynamically compact an intermediate query size to a noisy estimate close to the actual size, avoiding exhaustive padding.
 As such, queries under DPSCA experience largely boosted efficiencies (e.g., up to $10^{5}\times$ faster (Wang et al., 2022)) compared to their “no leakage” counterparts.
@@
 This approach can lead to either unbounded privacy loss or the forced cessation of query responses upon budget exhaustion.
-To mitigate this, some studies (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., [n.d.]) propose private, locality-sensitive grouping, incurring a one-time privacy cost to pre-group data based on specific attributes.
+To mitigate this, some studies (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., 2023) propose private, locality-sensitive grouping, incurring a one-time privacy cost to pre-group data based on specific attributes.
 Subsequent queries on those attributes can be directly applied to a smaller subset and need no additional privacy budget.
@@
 $\bullet$ L-2.
-Suboptimal execution plan.
+Unoptimized execution plan.
 Conventional query planners can pre-estimate sizes for equivalent plans of a given query and select the most efficient plan with minimized intermediate sizes before execution (Silberschatz et al., 2011; Blasgen et al., 1981).
 In contrast, SCA systems lack this capability, and even DPSCA designs (Qin et al., 2022; Wang et al., 2022; Zhang et al., 2023; Wagh et al., 2021; Bater et al., 2018, 2020) can only reactively determine plan sizes during runtime.
-This inherent limitation often forces existing systems to settle for less efficient query plans, such as suboptimal join orders, which lead to significantly inflated intermediate sizes (§ 7.2) and substantially hinder performance.
+This inherent limitation often forces existing systems to settle for less efficient query plans, such as costly join orders, which lead to significantly inflated intermediate sizes (§ 7.2) and substantially hinder performance.
 $\bullet$ L-3.
 Lossy processing.
-Noise from randomized mechanisms in DPSCA also introduces a unique accuracy issue (e.g., losing tuples after compaction), and unfortunately, no existing design can provide deterministic accuracy guarantees for complex queries (Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Zhang et al., 2023; Groce et al., 2019).
-Furthermore, stronger privacy settings can further increase noise variance, which amplifies errors, significantly impacting the utility of SCA systems.
-1.1. Overview of SPECIAL.
-In this work, we introduce SPECIAL, an innovative SCA system designed to address all these challenges simultaneously through a new paradigm called synopsis-assisted secure processing.
+Noise from randomized mechanisms in DPSCA also introduces a unique accuracy issue (e.g., conventional DP mechanisms may generate negative noise, applying which to obfuscate the sizes of intermediate query results can cause losing qualified real tuples), and unfortunately, no existing DPSCA can mitigate such loss for complex queries (Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Zhang et al., 2023; Groce et al., 2019).
+Furthermore, stronger DP settings can further increase noise variance, which amplifies errors, significantly impacting the utility of SCA systems.
+1.1. Overview of SPECIAL
+In this work, we introduce SPECIAL, an innovative SCA system that resolves the aforementioned limitations all at once through a new paradigm called synopsis-assisted secure processing.
 At its core, SPECIAL incurs a one-time privacy cost to gather DP synopses (statistics of base tables) from owners’ data.
-These synopses are then used to compact query intermediate results, index encrypted data, and enhance SCA query planning.
-Notably, SPECIAL is the first system to provide all of the following benefits: (1) Bounded privacy—it manages complex queries, such as multi joins, within strict privacy limits; (2) Advanced query planning—it builds an advanced SCA planner that can exploit plan sizes before runtime; and (3) Lossless processing—it ensures exact results with no data omissions.
-An overview of SPECIAL’s architecture is shown in Figure 1.
+These synopses are then used to accelerate complex query processing, and enhance SCA query planning.
+Notably, SPECIAL is the first system to provide all of the following benefits: (1) Bounded privacy—the privacy loss in SPECIAL is strictly limited to the one-time synopses release stage, with absotely no additional privacy cost during complex query processing and planning; (2) Advanced query planning—it builds an advanced SCA planner that can exploit plan sizes before runtime; and (3) Lossless processing—it ensures exact results with no data omissions.
+An overview of SPECIAL is shown in Figure 1.
 Figure 1. Overview of SPECIAL workflow.
 SPECIAL operates under a standard server-aided MPC model (Kamara et al., 2011) with three key participants: data owners, at least two SPECIAL servers, and a vetted analyst.
-The process starts with data owners securely outsourcing their data (e.g.
-using secret sharing (Kamara et al., 2011; Wang et al., 2021; Zhang et al., 2023; Wang et al., 2022)), and privately release a set of DP synopses (table statistics) to the servers.
-Once the data and synopses are ready, analysts can then submit queries on the outsourced data.
-To process queries, SPECIAL uses synopses to privately estimate tight result compaction bounds or build private indexes for faster access, thus avoiding exhaustive padding.
-Crucially, this information is exploit before runtime, so that SPECIAL can pre-identify an optimal secure execution plan that minimizes intermediate sizes and overall costs.
-Finally, the optimized plan is executed via MPC across the secret-shared data, and the results are securely returned to the analyst.
-In the SPECIAL prototype, we adopt the same privacy model used in existing DPSCA designs, assuming a computationally bounded, “honest-but-curious” adversary (Goldreich, 2004) who can compromise up to $n-1$ out of $n$ owners and $m-1$ out of $m$ servers.
-SPECIAL ensures that such an adversary can only learn a limited leakage profile $\mathsf{Lkg}$ about honest parties’ data, which strictly adheres to DP constraints.
-Note that, SPECIAL’s core components, including synopses management and query planning do not leverage raw data at all.
-Thus, any of the servers, alone or jointly, can manage these tasks.
-Also, once analysts are authorized by data owners, they are considered trusted, and additional randomization of query results is not required.
-It is worth mentioning that SPECIAL’s corruption model (e.g.
-all except one) is entirely determined by the underlying MPC, and SPECIAL’s design does not modify them at all.
-Hence, one may also opt to a weaker corruption model, such as a supermajority of owners and servers must remain uncorrupted (Liagouris et al., 2023; Tan et al., 2021), to enhance efficiency.
+The process begins with data owners securely outsourcing their data, typically through secret sharing (Kamara et al., 2011; Wang et al., 2021; Zhang et al., 2023; Wang et al., 2022), and privately releasing corresponding DP synopses (§ 4) to the servers.
+SPECIAL introduces a set of novel primitives (§ 5) that can leverage these synopses to accelerate secure query operations.
+Once the data and synopses are in place, analysts can submit Select-Project-Join-Aggregation (SPJA) queries (Silberschatz et al., 2011) for analytics.
+To process queries, a private planner (§ 6), running on SPECIAL servers, strategically orchestrates SPECIAL primitives to process the query and optimize performance.
+Finally, the results are securely returned to the analyst.
 1.2. Unique challenges and key contributions
@@
 How to select proper synopses?
-Even a single relation can have numerous options for generating synopses, such as using different attributes or their combinations.
-However, improper selection can lead to large errors (e.g.
-using too many synopses or leverage high-dimensional attributes (Zhang et al., 2021)), or reduced functionalities (e.g., using only simple attributes (Zhang et al., 2023)).
+Even for a single relation, one can find numerous attribute combinations for generating synopses.
+Improper selection can lead to large errors (e.g.
+using too many synopses or high-dimensional attributes (Zhang et al., 2021)), or reduced functionalities (e.g., using only simple attributes (Zhang et al., 2023)).
 Hence, a key challenge is selecting a limited set of DP synopses to optimize the privacy budget for complex query processing.
@@
 Thus, a second challenge is designing practical approaches to achieve lossless results without violating privacy goals.
-To address this, we employ one-sided DP noise (either strictly positive or negative, § 4) for generating synopses and design novel primitives (§ 5) based on them to pessimistically estimate filter cardinalities and intervals of index structures.
-This enable us to efficiently discard unnecessary records during processing without breaking privacy and lossless guarantees.
-To ensure lossless processing of complex joins, we extend upon cutting-edge join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join compaction sizes using available synopses.
+To address this, we employ one-sided DP noise (either strictly positive or negative, § 4) in generating synopses, and design novel primitives (§ 5) based on them to pessimistically estimate filter cardinalities and intervals of index structures.
+To ensure lossless processing of complex joins, we extend upon cutting-edge join upper bound techniques (Hertzschuch et al., 2021) to privately estimate lossless join sizes using DP synopses.
 To our knowledge, this is the first study to support private join upper bound estimation.
@@
 To navigate this potential, we explore various use of synopses in accelerating secure processing including private indexes SPEidx (§ 5.2), and compacted oblivious operations SPEop (§ 5.3).
-We then develop a Selinger-style query planner (§ 6) that supports both standard secure operators and SPEop.
-Employing a customized cost model and tailored heuristics, this planner efficiently generates optimal execution plans for queries using a mix of standard and SPEop operators.
-This in essence, also address the limitations of sub-optimal execution plans inherent in standard DPSCA.
+We also design a novel private query planner (§ 6) that efficiently orchestrate the execution of SPECIAL primitives (e.g., SPEidx, SPEop) to process SPJA queries.
+The planner uses available synopses to privately estimate intermediate result sizes and operation costs for a set of equivalent execution plans of a given query.
+It then executes the one with the lowest estimated cost.
 $\bullet$ C-4.
 How to systematically evaluate SPECIAL?
-A major challenge in SCA design evaluation is the absence of open-source benchmarks.
-We address this by initiating an open-source evaluation platform accessible to the public, utilizing public Financial data (Rel, [n.d.]) (1.1M rows and 59 columns), and designing eight test queries ranging from simple linear to complex 5-way joins.
-We evaluated our prototype SPECIAL alongside the state-of-the-art DPSCA system, Shrinkwrap (Bater et al., 2018), and conventional SCA, SMCQL (Bater et al., 2017).
-Results indicate that SPECIAL outperforms Shrinkwrap, reducing query latency by up to 80.3×80.3$\times$80.3 ×, and SMCQL, with at least a 114×114$\times$114 × reduction in query latency.
-Additionally, SPECIAL improves memory efficiency in complex join processing by more than 900×900$\times$900 × compared to both systems.
-Moreover, scaling experiments show that SPECIAL can effectively scale linear and binary joins up to 8×8$\times$8 × (8.8M rows) and 5-way joins up to 4×4$\times$4 × (4.4M rows) workloads.
+A major is the absence of open benchmarks.
+We address this by initiating an open-source evaluation set, accessible to the public.
+Specifically, we use public financial data (Rel, [n.d.]) and design eight test queries, ranging from simple linear queries to complex 5-way joins.
+We also re-produce an open version of the HealthLNK benchmark.
+We evaluate our prototype, SPECIAL, against the state-of-the-art (SOTA) DPSCA system, Shrinkwrap (Bater et al., 2018), and the conventional SCA system, SMCQL (Bater et al., 2017).
+Results indicate that SPECIAL outperforms Shrinkwrap, reducing query latency by up to $80.3\times$, and SMCQL, with at least a $114\times$ reduction in query latency.
+Additionally, SPECIAL improves memory efficiency in complex join processing by more than $900\times$ compared to both systems.
+Moreover, scaling experiments show that SPECIAL can effectively scale up to 8.8M rows dataset and up to 9-way complex joins.
 All benchmarks, including our prototype implementation, are open-sourced and available at (Lovingmage, 2024).
-The remainder of the paper is structured as follows: We begin with essential background in § 2 and a formal definition of our privacy model in § 3 before delving into design details.
-§ 4 discusses the SPECIAL catalog design, including synopses selection and generation.
-§ 5 details novel synopses-assisted secure database operators, and § 6 describes our planner design.
-Benchmark evaluations are presented in § 7, followed related works and conclusions.
 2. Background
 General notations.
-We consider the logical database $\mathcal{D}$ to contain multiple private relations $\{D_{1},D_{2},...\}$, where each relation $D_{i}$ is belongs owned by a specific party $P_{i}$, where $D\in\mathcal{D}$ is owned by a specific owner and has a set of attributes as $attr(D)$.
-The domain for attribute $A\in attr(D)$ is given by $dom(A)$, and the combined domain for a collection of attributes $\mathbf{A}=\{A_{1},A_{2},...\}\subseteq attr(D)$ is denoted as $dom(\mathbf{A})=\prod_{A\in\mathbf{A}}dom(A)$.
-For a tuple $t\in D$, and $\mathbf{A}\subseteq attr(D)$, we denote $t.\mathbf{A}$ as the attribute value of $\mathbf{A}$ in $t$
+We consider the logical database $\mathcal{D}$ to contain multiple private (base) relations $\{D_{1},D_{2},$...\}, where each relation $D_{i}$ is owned by a specific party $P_{i}$.
+A base relation $D$ (we omit subscript for simplicity) has a set of attributes $attr(D)$.
+The domain of an attribute $A\in attr(D)$ is denoted by $dom(A)$, and the combined domain of a collection of attributes $\mathbf{A}=\{A_{1},A_{2},...\}\subseteq attr(D)$ is denoted by $dom(\mathbf{A})=\prod_{A\in\mathbf{A}}dom(A)$.
+For a tuple $t\in D$, and $\mathbf{A}\subseteq attr(D)$, we use $t.\mathbf{A}$ to denote the attribute value of $\mathbf{A}$ in $t$.
+A logical query, represented by $q(\mathcal{D})$, applies transformations and computations on $\mathcal{D}$.
+In this work, we focus on SPJA (Silberschatz et al., 2011) queries.
 Frequency (count).
-Given $D$, and $\mathbf{A}\subseteq attr(D)$, and $\mathbf{v}\in dom(\mathbf{A})$, the frequency (count) of $\mathbf{v}$ in $D$ is the total number of tuples $t\in D$ with $t.\mathbf{A}=\mathbf{v}$.
+Given $D$, $\mathbf{A}\subseteq attr(D)$, and a set of values $\mathbf{v}\in dom(\mathbf{A})$, the frequency (count) of $\mathbf{v}$ in $D$ is the total number of tuples $t\in D$ with $t.\mathbf{A}=\mathbf{v}$.
 In addition, the max frequency moments (MF) of $\mathbf{A}$ is defined as ${\mathtt{mf}}(\mathbf{A},D)=\max_{\mathbf{v}\in dom(\mathbf{A})}|\{t\in D\mid t.\mathbf{A}=\mathbf{v}\}|$.
@@
 Given $D$, and $\mathbf{A}\subseteq attr(D)$, the (equal-width) histogram $\mathbf{h}(\mathbf{A},D)=(c_{1},c_{2},...,c_{m})$ is a list of counts for the attribute values in $\mathbf{A}$.
-Specifically, $\mathbf{h}$ partitions [^1] We use a condensed notation that omits arguments when $\mathbf{A}$ and $D$ are obvious.
-$dom(\mathbf{A})$ into $m$ “equal-sized” domain intervals $(B_{1},...,B_{m})$, and each count $c_{i}\in\mathbf{h}$ corresponds to the total number of $t\in D$ such that $t.\mathbf{A}$ fall within $B_{i}$.
-Multi-party secure computation (MPC).
-MPC (Yao, 1986; Goldreich, 2009; Micali et al., 1987; Ben-Or et al., 2019) is a cryptographic techniques to allow a group of participants $P_{1},P_{2},...$ to jointly compute a function $f(x_{1},x_{2},...)$ over their respective private input $x_{i}$.
-MPC ensures that no unauthorized information is revealed to any involved party, excepted the desired output of function $f$, emulating a computation as if performed by a trusted third party.
-Earlier MPC typically require all parties to participate in intensive computations.
-Recent schemes (Kamara et al., 2011; Scholl et al., 2017; Mohassel and Zhang, 2017) allow these computations to be offloaded to multiple powerful servers without sacrificing original security guarantees, known as the server-aided MPC.
-In this approach, each party, say $P_{i}$, secretly shares their inputs, $x_{i}$ to servers.
-These servers then jointly evaluate an MPC that functionally replicates the logic of first reconstructing the secrets $x_{1},x_{2},\ldots$ and then evaluating $f(x_{1},x_{2},\ldots)$.
-Depending on designs, the result may be directly returned to the parties or re-shared among the servers for later retrieval.
-By default, SPECIAL considers server-aided MPC with two non-colluding servers.
-Secret sharing and secure array.
-SPECIAL uses the $2$-out-of-$2$ boolean secret share (Araki et al., 2016) over ring $\mathbb{Z}_{2^{32}}$ for securely outsourcing owners’ data and storing query execution results.
-Specifically, each data, $x$, is divided into two shares: $x1,x2$ that are uniformly distributed over the ring $\mathbb{Z}_{2^{32}}$ such that $x=x1\oplus x2$.
-Each server $S_{i}$ receives one secret shares, $s_{i}$, where $i\in\{0,1\}$.
-By retrieving shares from any two servers, an authorized party can successfully reconstruct the value of $x$.
-However, a single server alone learns nothing about $x$.
-For clarity and to abstract out the lower-level details, we leverage a logically unified data structure, namely the secure array (Wang et al., 2022; Bater et al., 2018), denoted as $\langle\mathbf{x}\rangle=\left(\langle x_{1}\rangle,\langle x_{2}\rangle,...\right)$, which is a collection of secret-shared relational tuples.
+Specifically, $\mathbf{h}$ partitions $dom(\mathbf{A})$ into $m$ “equal-sized” domain intervals $(B_{1},...,B_{m})$, and a count $c_{i}\in\mathbf{h}$ is the number of tuples $t\in D$ with $t.\mathbf{A}$ in the interval of $B_{i}$.
+Query planning.
+Modern databases parse queries into physical plans (Silberschatz et al., 2011) that can be executed by the underlying query engine.
+These plans specify the operations like scans, joins, and sorts, and the order in which they’re performed.
+The same query can have multiple equivalent execution plans, but their performance can vary greatly depending on resource usage and data access patterns.
+Query planning (Blasgen et al., 1981), done before runtime, involves selecting cost-efficient plans from these options.
+A key part of this process is accurately estimating intermediate result sizes, known as cardinality estimation (CE) (Harmouch and Naumann, 2017), which relies heavily on table statistics.
+Two crucial statistics in modern cardinality estimation methods are: (i) histograms, which are vital for estimating selectivities in filters, and (ii) max frequency, which is crucial for estimating join sizes.
+[…]
+Differential privacy (Dwork et al., 2014).
+DP ensures that modifying a single input tuple to a mechanism produces only a negligible change in its output.
+To elaborate, consider $D$ and $D^{\prime}$ as two relations differing by just one tuple, then DP defines the following.
+Definition 2.1 ($(\epsilon,\delta)$-DP).
+Given $\epsilon>0$, and $\delta\in(0,1)$.
+A randomized mechanism $\mathcal{M}$ is said to be $(\epsilon,\delta)$-DP if for all ${D}\sim{D^{\prime}}$ pairs, and any possible output $o\subset Range(\mathcal{M})$, the following holds: Pr[M(D) o] e^ϵ Pr[M(D’) o] + δ
+[…]
 Oblivious (relational) operators.
-Oblivious operators are specific MPC protocols, applied over secure arrays, that implement database operator functionalities such as filter, join, and aggregation.
-Specifically, an oblivious operator $\langle\mathbf{o}\rangle\leftarrow\mathsf{\textbf{op}}(\langle\mathbf{a}\rangle,\langle\mathbf{b}\rangle,...)$ accepts one or more secure arrays as inputs, performs database operations over the inputs, and then writes the result to an output secure array.
-Moreover, the oblivious property stipulates that absolutely no memory access pattern information is disclosed (Chang et al., 2022; Ramachandran and Shi, 2021; Liagouris et al., 2023) (or in other words, the memory access is fully independent from input data) during the protocol’s execution.
-Since revealing output size can also enable numerous reconstruction attacks, recent designs of oblivious operators also consider hiding output volumes, i.e., by padding the outputs to their maximum sizes.
-Thus, unless otherwise stated, we assume that all the subsequent oblivious operators mentioned have considered volume-hiding.
-Differential privacy (Dwork et al., 2014).
-DP ensures that modifying a single input data tuple to a mechanism produces only a negligible change in its output.
-To elaborate, consider $D$ and $D^{\prime}$ as two databases differing by just one tuple—termed as neighboring databases.
-Then:
-Definition 2.1 ($(\epsilon,\delta)$-differential privacy).
-Given $\epsilon>0$, and $\delta\in(0,1)$.
-A randomized mechanism $\mathcal{M}$ is said to be $(\epsilon,\delta)$-DP if for all pairs of neighboring databases $D,D^{\prime}$, and any possible output $o\subset Range(\mathcal{M})$, the following holds:
-$\textup{Pr}\left[\mathcal{M}(D)\in o\right]\leq e^{\epsilon}\textup{Pr}\left[\mathcal{M}(D^{\prime})\in o\right]+\delta$
-3. Privacy Model
-In this section, we formally define the privacy definitions of SPECIAL and explain the corresponding semantics.
-Similar to previous work on DPSCA (Wang et al., 2022; Zhang et al., 2023), we follow computational DP (SIM-CDP) (Mironov et al., 2009) framework for privacy model formulation.
-Definition 3.1 (Secure protocol with DP leakage).
-Given a party (owner) $P$ with private data $D$ and a randomized mechanism $\mathsf{Lkg}(D)$, referred to as the leakage profile of $P$’s data.
-Consider $\Pi$ a secure query processing protocol applied to $D$.
-The protocol $\Pi$ is said to be secure with DP leakage if, for any honest party $P$ with data $D$ and any probabilistic polynomial time (p.p.t.) adversary $\mathcal{A}$:
-  • $\mathsf{Lkg}(D)$ satisfies $(\epsilon,\delta)$-DP (definition 2.1).
-  • There exists a p.p.t. simulator $\mathcal{S}$ with only access to public parameters $\mathsf{pp}$ and the output of $\mathsf{Lkg}(D)$ that satisfies:
-    (1) $\begin{split}&\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\Pi}(\mathcal{D},\mathsf{pp})=1\right)\right]\\ &\leq\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\mathcal{S}}(F(\mathcal{D}),\mathsf{pp})\right)=1\right]+\mathsf{negl}(\kappa)\end{split}$
-where $\mathsf{VIEW}^{\Pi}$, and $\mathsf{VIEW}^{\mathcal{S}}$ denotes the adversary’s view against the protocol execution and the simulator’s outputs, respectively.
-And $\mathsf{negl}(\kappa)$ is a negligible function related to a security parameter $\kappa$.
-Simply put, Definition 3.1 specifies that as long as there is at least one honest owner, the knowledge any p.p.t.
-adversary can obtain about this owner’s data from observing $\Pi$ execution transcripts is bounded to what can be inferred from the outputs of an $(\epsilon,\delta)$-DP mechanism $\mathsf{Lkg}$.
-By default, we define $\mathsf{Lkg}$ within $(\epsilon,\delta)$-tuple level DP (Dwork et al., 2010), that’s said primary privacy objective is to protect individual tuples.
-However, due to group-privacy properties (Kifer and Machanavajjhala, 2011; Xiao and Xiong, 2015) of DP, privacy can extend across multiple tuples and to user-level, as long as a user owns a limited number of tuples.
-Throughout this paper, our focus remains on developing algorithms that guarantee tuple-level privacy, while upholding all stated privacy guarantees, albeit with potentially different privacy parameters.
-Theorem 3.2.
-The privacy of SPECIAL adheres to Definition 3.1
-Proof.
-Due to space concerns, we defer the complete privacy proof of SPECIAL to Appendix A. ∎
+Oblivious operators are data-independent MPC protocols that implement the same functionalities as their plaintext database counterparts (e.g., filter and join).
+Data-independent execution requires that the control flow and memory access patterns of a function are indistinguishable given different inputs of the same size, and typically requires costly computation.
+For example, a linear scan is required to fulfill oblivious filtering (Zheng et al., 2017), and join requires nested-loop over the two inputs (Eskandarian and Zaharia, 2017).
+The output sizes of such operators are usually padded with dummy tuples to the worst case: $N$ rows for filters and $N^{2}$ for joins, given size $N$ inputs.
+The dummy tuples will not affect the query result but can significantly impact the performance (Zheng et al., 2017; Eskandarian and Zaharia, 2017).
+To enhance efficiency while maintaining strong privacy, DPSCAs introduce a new type of oblivious operators (Bater et al., 2018; Qin et al., 2022; Qiu et al., 2023; Wang et al., 2022).
+These operators typically involve two steps: Compute and Compact.
+The Compute step is fully oblivious, while the Compact resizes the output, often by obliviously sorting valid tuples to the front and trimming the output to a noisy DP size (the true size plus DP noise).
+While this approach can significantly reduce the computation cost and query sizes, it may lead to lossy query processing if the DP size is smaller than the true size (e.g., negative DP noise), as valid tuples could be excluded during the compaction (Wang et al., 2021, 2022; Zhang et al., 2023).
+We emphasize that when DP sizes exceed true sizes, there is no accuracy loss as it only includes extra dummy data that do not impact accuracy (Zheng et al., 2017; Eskandarian and Zaharia, 2017).
+Private indexes.
+In conventional databases, indexes are powerful data structures that map attribute values to positions in a sorted array, allowing a predicate selection to quickly access the desired data via index lookup without the need for full table scan.
+However, traditional indexes are unsuitable in SCAs due to their data-dependent nature, which can easily lead to privacy breaches.
+To address this, recent research has proposed DP indexes (Roy Chowdhury et al., 2020), where the mapping of attribute values to their positions is intentionally distorted with DP noise.
+To process queries, the system first pre-fetch a small set of data using DP indexes, followed by oblivious selection.
+This effectively avoids full table oblivious scan and sorting-based result compaction.
+However, the uncertainty inherent in DP indexes can lead to the loss of valid tuples.
+For example, in the DP index of (Roy Chowdhury et al., 2020), a true index range of positions $[10,20]$ might be distorted into positions $[12,17]$, causing data at position $10,11,18,19$ and $20$ to be missed.
+Nevertheless, if DP indexes overestimate the range, subsequent oblivious selection can losslessly identify all valid tuples.
+3. System and Privacy Model
+In general, SPECIAL follows a standard server-aided MPC (Kamara et al., 2011) model, involving (i) a set of mutually distrustful data owners $P_{1},...,P_{n}$}, (ii) two non-colluding servers $S_{0}$ and $S_{1}$, and (iii) a trusted analyst.
+We assume an admissible adversary (Mohassel and Zhang, 2017) $\mathcal{A}$, capable of corrupting n1n-1 out of $n$ clients and at most one of the two servers.
+An instance of such adversary can be a malicious server that creates Sybil owners to form a malicious collation, attempting to steal sensitive information from an honest owner.
+Additionally, $\mathcal{A}$ is considered honest-but-curious, meaning it follows the protocol without deviation but may try to infer information from observed protocol transcripts, such as randomness, memory access patterns, and communication messages.
+The combination of these information is referred to as the view of $\mathcal{A}$.
+We also assume $\mathcal{A}$ is computationally bounded as a probabilistic polynomial time (p.p.t.) adversary, which is a standard requirement in MPC protocols to ensure that adversaries cannot break cryptographic primitives.
+This threat model is consistent with prior SCA designs (Bater et al., 2017, 2018; Wang et al., 2022; Mohassel and Zhang, 2017; Wang et al., 2021).
+Given this setup, we design SPECIAL to satisfy the following:
+Definition 3.1 (MPC protocol with DP leakage).
+Given a set of parties (owners) $P_{i}$ with private data $D_{i}$ and a secure query protocol $\Pi$ that applies over $\mathcal{D}=\{D_{1},D_{2},\cdots\}$.
+We define a randomized mechanism $\mathsf{Lkg}(\mathcal{D})=\{\mathsf{Lkg}(D_{1}),\mathsf{Lkg}(D_{2}),\cdots\}$ as the leakage profile, consisting of the control flow and access patterns of running $\Pi$ over $\mathcal{D}$.
+The protocol $\Pi$ is said to be secure with DP leakage if, for the subset of uncorrupted parties with data $\mathbf{D}\subseteq\mathcal{D}$, leakage profile $\mathsf{Lkg}(\mathbf{D})\subseteq\mathsf{Lkg}(\mathcal{D})$, and any p.p.t.
+adversary $\mathcal{A}$:
+  • $\mathsf{Lkg}(\mathbf{D})$ satisfies $(\epsilon,\delta)$-DP (definition 2.1).
+  • There exists a p.p.t. simulator $\mathcal{S}$ with only access to public parameters $\mathsf{pp}$ and $\mathsf{Lkg}(\mathbf{D})$ that satisfies:
+    (1) $\begin{split}&\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\Pi}(\mathcal{D},\mathsf{pp})=1\right)\right]\\ &\leq\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\mathcal{S}}(\mathsf{Lkg}(\mathbf{D}),\mathsf{pp})\right)=1\right]+\mathsf{negl}(\kappa)\end{split}$
+where $\mathsf{VIEW}^{\Pi}$ is $\mathcal{A}$’s view in $\Pi$’s execution and $\mathsf{VIEW}^{\mathcal{S}}$ is a simulated view produced by $\mathcal{S}$ using $\mathsf{Lkg}$; $\mathsf{pp}$ denotes all public parameters, and $\mathsf{negl}(\kappa)$ is a negligible function related to a security parameter $\kappa$.
+Simply put, Definition 3.1 requires that the knowledge any p.p.t.
+adversary adversary can gain about each individual tuple of an honest owner, by observing the protocol execution, is bounded to what can be inferred from the outputs of the $(\epsilon,\delta)$-DP mechanism $\mathsf{Lkg}$.
+We stress that this notion focuses on DP at the event (tuple) level without loss of generality.
+Due to the group-privacy properties of DP (Dwork et al., 2010; Kifer and Machanavajjhala, 2011; Xiao and Xiong, 2015; Vadhan, 2017), event-level DP can be extended to user-level DP.
+For instance, in a logical database $\mathcal{D}$ where any single user owns at most $l$ tuples, if a protocol satisfies $(\epsilon,\delta)$ event-level DP, it also satisfies $(l\epsilon,le^{(l-1)\epsilon}\delta)$ user-level DP.
+Moreover, we say that SPECIAL can be relaxed to employ a weaker corruption model, such as requiring a supermajority of owners and servers to remain uncorrupted, to enhance efficiency (Liagouris et al., 2023; Tan et al., 2021).
+This adjustment does not change the privacy guarantee outlined in Definition 3.1, but it does affect the security assumptions.
+Under the relaxed corruption model, Definition 3.1 is only satisfied when at least two-thirds of the parties remain uncorrupted.
+Due to space concerns, we defer the complete privacy proof of SPECIAL to our full version (Chenghong Wang, 2024).
 4. SPECIAL Synopses
-SPECIAL maintains independent DP synopses for all outsourced data.
-Each synopsis holds noisy statistics that reflect the corresponding data’s distributional traits.
-As per our core design idea, these statistics will be used to dynamically optimize query processing—for example, by estimating tight compaction bounds or enabling fast access to eliminate unnecessary tuples, which significantly compacts intermediate query sizes, all without any extra privacy cost.
+We now discuss the details of private synopses used in SPECIAL, while in later sections we will show how they accelerate query processing (§ 5) and aid in query planning (§ 6).
 Challenges.
-A major design challenge is selecting proper synopses to achieve our design goals.
-This involves decisions such as attribute selection, combinations thereof, and the types of statistics used (e.g., histograms, min/max/avg, frequency counts, etc.).
-Using too many attributes or leveraging complex (high-dimensional) combinations can introduce significant noise (Zhang et al., 2021) and increase storage costs.
-Another challenge involves privately releasing synopses to aid in lossless processing, such as result compaction and private indexing.
-Ensuring lossless compaction requires consistently overestimating the output sizes of subplans.
-Supporting lossless indexes is even more complex: for a given indexed range $D[\mathsf{lo},\mathsf{hi}]$, we need to overestimate the upper index $\mathsf{hi}$ and underestimate the lower index $\mathsf{lo}$.
-Unfortunately, traditional DP methods, which typically use symmetric additive noises, struggle to ensure these conditions.
+We reiterate the main challenges in generating private synopses for a relation include: (C-1) selecting a set of attribute combinations that enable functional and efficient query processing; (C-2) ensuring that the subsequent query processing based on the private synopses is lossless.
 Key ideas.
-We observe that joins are more resource-intensive than any other operations.
-For instance a $k$-way oblivious join can incur a $O(n^{k})$ complexity if without optimisations (Bater et al., 2017, 2018; Wang et al., 2022; Zheng et al., 2017; Eskandarian and Zaharia, 2017).
-Hence, joins require prioritized acceleration.
-Moreover, synopses available for frequently queried filter attributes on base relations are vital (Wang et al., 2022) as well.
-These statistics not only help to compact input data at an early stage before heavy joins, but also enable fast indexing (§ 5.2) for quick data access.
-Hence, to address the first challenge, our key idea is to focus on commonly queried join and base relation filter attributes, as well as their low-dimensional combinations.
-To tackle the second challenge, our key idea is twofold.
-First, to ensure lossless result compaction of filters and indexing, we employ one-sided DP noise to generate private histograms that either overestimate or underestimate attribute distributions consistently.
-Second, to achieve lossless compaction of join outputs, we integrate noisy max frequency moments (MF) into the synopses.
-This integration enables us to extend upon advanced join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join compaction sizes without losing data (§ 5.2).
+Given a SPJA query, join operations typically need prioritized acceleration, as they are more resource-intensive than other operations.
+A $k$-way join can have $O(n^{k})$ complexity without optimizations (Bater et al., 2017, 2018; Wang et al., 2022; Zheng et al., 2017; Eskandarian and Zaharia, 2017).
+Additionally, synopses for frequently queried filter attributes play an important role in efficient query processing, as they enable fast indexing (§ 5.2) and effective filtering of unnecessary data before heavy joins.
+As such, our first key idea is to focus on histograms-based synopses that cover frequently queried join and filter attributes.
+To minimize noise, we focus on low-dimensional synopses: only 1D or 2D histograms.
+To address the second challenge, our approach combines two strategies.
+First, to support accurate indexing and filtering, we use one-sided DP noise to generate DP histograms that consistently overestimate or underestimate attribute distributions.
+We will show later that such special histograms allow lossless filtering and indexing that reliably overestimate true filter sizes and indexing ranges (§ 5.2).
+Second, for lossless join output compaction, we incorporate noisy max frequency moments (MF) into the synopses.
+MF allows us to build on advanced join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join sizes without data loss (§ 5.2).
 4.1. Synopses generation
-Steered by our key ideas, we now elaborate on our synopsis generation details.
-The general process is outlined as follows: Initially, the SPECIAL servers selects appropriate attributes to generate private synopses, which are then distributed to owners.
-The owners create the synopses using a DP mechanism and upload them to SPECIAL servers.
-Finally, the server applies post-processing to these synopses to aid in query processing.
+We now elaborate on the details of synopsis generation, which mainly contains two phases: (i) Attributes selection, where the SPECIAL servers select appropriate attributes for the generation, which are then distributed to owners; (ii) Local synopses release, where the owners create corresponding synopses using a DP mechanism and upload them to SPECIAL servers.
 Attributes selection (servers).
-The first step is to identify a smaller set of attributes for deriving synopses.
-In general, we consider the existence of a representative workload, $Q_{\mathsf{R}}$ (Kotsogiannis et al., 2019), which can be sourced from a warm-up run of SPECIAL or annotated by the administrator.
-Note that this process do not incur any private data thus is leakage-free.
+The first step is to identify a set of attributes for deriving synopses.
+In general, we consider the existence of a representative workload, $Q_{\mathsf{R}}$ (Kotsogiannis et al., 2019), which can be sourced from a warm-up run or annotated by the administrator.
+Note that the representative workload does not involve any private data and thus is leakage-free.
 The servers first identify representative attribute pairs, $\mathsf{pair}=\{\mathsf{pair}_{k}\}_{k\geq 1}$, for each private relation $D\in\mathcal{D}$ via $Q_{\mathsf{R}}$.
@@
 Synopses release (owners).
-Next, servers pushes the identified pairs to each owner, and subsequently, the owner independently dispatches a private synopsis and returns it to servers.
-In what follows, we focus on the DP synopses generation mechanism evaluated by each owner, Algorithm 1 illustrates the overall flow.
-Algorithm 1 DP synopsis generation $\mathcal{M}_{\mathtt{synop}}$ (in the view of $P$)
+Next, servers pushes the identified pairs to owners, and subsequently, the owners independently dispatche private synopses and return them to servers.
+We now focus on the DP synopses generation mechanism run by each owner, Algorithm 1 illustrates the workflow.
+Algorithm 1 DP synopsis gen $\mathcal{M}_{\mathtt{synop}}$ (in the view of $P$)
 1:Input: $\mathsf{pair}=\{\mathsf{pair}_{k}\}_{k\geq 1}$ from servers; private data $D$.
@@
 7:  $\mathbf{h}^{-}(\mathsf{pair}_{k},D)\leftarrow\mathbf{h}+\mathsf{Lap}^{-}(\epsilon,\delta,\mathbf{h}.\mathsf{shape})$
-8$\triangleright$ adding independently sampled noise to every bin of $\mathbf{h}^{+}$, $\mathbf{h}^{-}$
+[…]
 9:DP max frequencies:
@@
 11:  else if $A_{\mathsf{ft}}\in\mathsf{pair}_{k}\neq\emptyset$ then
-12$\triangleright$ assuming $\mathbf{h}$ partitions $dom(A_{\mathsf{ft}})$ into $\{B_{1},...,B_{m}\}$
+[…]
 13:   $D^{\ell}\leftarrow\sigma_{A_{\mathsf{ft}}\in B_{\ell}}(D)$ for $\ell=1,2,...,m$
@@
 17:release $\mathtt{synop}$, $\epsilon$, $\delta$ to servers
-Generally, we expect owners to set a desired privacy budget for their data (using parameters $\epsilon$ and $\delta$) before initiating the synopsis generation.
-For each $\mathsf{pair}_{k}$, the owner first creates a histogram $\mathbf{h}(\mathsf{pair}_{k},D)$.
-By default, we assume there exist global parameters, i.e.
-bin sizes of each attribute, which guide all owners to partition attributes consistently.
-Next, the owner derives two noisy histograms, $\mathbf{h}^{+}(\mathsf{pair}_{k},D)$, and $\mathbf{h}^{-}(\mathsf{pair}_{k},D)$ by adding independently sampled one-sided Laplace noises to every bin of $\mathbf{h}(\mathsf{pair}_{k},D)$.
+In general, we expect owners to set a desired privacy budget for their data (using parameters $\epsilon$ and $\delta$). Algorithm 1 produces synopses formalized as :
+Definition 4.1 (SPECIAL synopses).
+Given $Q_{\mathsf{R}}$, we consider for each relation $D$, its corresponding synopsis $\mathtt{synop}$ is the collection of $\{(\mathsf{pair}_{k},\mathbf{H}(\mathsf{pair}_{k},D),\mathtt{MF}_{k})\}_{k\geq 1}$, such that
+  • $\mathsf{pair}_{k}=(A_{\mathsf{ft}},A_{\mathsf{j}})\in Q_{\mathsf{R}}$ is a frequently queried attribute pair.
+  • $\mathbf{H}(\mathsf{pair}_{k},D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$ is the DP bounding histogram for $\mathsf{pair}_{k}$, where $\mathbf{h}^{+}$ (resp.
+$\mathbf{h}^{-}$) is a DP histogram that overestimate (resp.
+underestimate) the true histogram of $\mathsf{pair}_{k}$.
+  • $\mathtt{MF}_{k}$ represents a collection of privately overestimated join key MFs categorized by $\mathsf{pair}_{k}.A_{\mathsf{ft}}$.
+We now present a detailed explanation of generating these synopsis structures, starting with the private bounding histograms (Alg 1, lines 2:5).
+Specifically, for each $\mathsf{pair}_{k}$, each owner first constructs a histogram $\mathbf{h}(\mathsf{pair}_{k},D)$.
+By default, we assume there exists global parameters (e.g., bin sizes) for each attribute, that ensure consistent data partitioning among all owners.
+Next, the owner derives two noisy histograms, $\mathbf{h}^{+}(\mathsf{pair}_{k},D)$ and $\mathbf{h}^{-}(\mathsf{pair}_{k},D)$, by adding independently sampled one-sided Laplace noise (Definition 4.2) to every bin of $\mathbf{h}(\mathsf{pair}_{k},D)$.
+This guarantees that $\mathbf{h}^{+}$ always overestimates the true histogram, while $\mathbf{h}^{-}$ consistently underestimates it.
 Definition 4.2 (One-sided Laplace variable).
 $\mathsf{Lap}^{+}(\epsilon,\delta)=\max(0,z)$ (resp.
-$\mathsf{Lap}^{-}(\epsilon,\delta)=\min(0,z)$) is a one-sided Laplace random variable in the range of $[0,\infty)$ (resp.
+$\mathsf{Lap}^{-}(\epsilon,\delta)=\min(0,z)$) is a one-sided Laplace random variable in the range of $0,\infty)$ (resp.
 $(-\infty,0]$) if $z$ is drawn from a distribution with the following density function
@@
 where $\mu=1-\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))$ (resp. $\mu=\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))-1$).
-The one-sided Laplace noise can be strictly positive ($\mathsf{Lap}^{+}$) or negative ($\mathsf{Lap}^{-}$).
-This ensures that $\mathbf{h}^{+}$ always overestimates the actual histogram, while $\mathbf{h}^{-}$ consistently underestimates it.
-For the ease of description, we refer $\mathbf{H}(\mathsf{pair}_{k},D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$ as the bounding histogram of attribute $\mathsf{pair}_{k}$, with $\mathbf{h}^{+}$, and $\mathbf{h}^{-}$ to be the upper and lower histograms, respectively.
-We continue with the details of the (noisy) MFs generation.
-Assume both $A_{\mathsf{ft}},A_{\mathsf{j}}\neq\emptyset$, and the $\mathbf{h}(\mathsf{pair}_{k},D)$ partitions $dom(A_{\mathsf{ft}})$ into $\{B_{1},...,B_{m}\}$.
-The mechanism $\mathcal{M}_{\mathtt{synop}}$ generates a table of noisy MFs, $\{\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\}_{\ell\geq 1}$, where each entry $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})$ is an independent MF statistic for $A_{\mathsf{j}}$ over specific filtered data on $A_{\mathsf{ft}}$, such that
+Next, we detail the generation of (noisy) join key MFs (Alg 1 lines 6:10).
+We assume that both $A_{\mathsf{ft}}$ and $A_{\mathsf{j}}$ are non-empty and that $\mathbf{h}(\mathsf{pair}_{k},D)$ partitions $dom(A_{\mathsf{ft}})$ into bins $\{B_{1},\dots,B_{m}\}$.
+Each owner then generates a table of noisy MFs, $\{\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\}_{1\leq\ell\leq m}$, where each entry $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})$ represents an independently generated MF statistic for the join key attribute $A_{\mathsf{j}}$, calculated over a specific subset of data filtered by the attribute $A_{\mathsf{ft}}$, such that
 (3) $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\leftarrow\widehat{\max}_{\epsilon}\left(\mathcal{G}_{\mathbf{count}(A_{\mathsf{j}})}\left(\sigma_{A_{\mathsf{ft}}\in B_{\ell}}(D)\right)\right)$
@@
 Nevertheless, as exponential noises are non-negative, thus $\widehat{\mathtt{mf}}\geq\mathtt{mf}$ holds for all cases.
-Given the discussion so far, we summarize SPECIAL synopses as:
-Definition 4.2 (SPECIAL synopses).
-Given representative workload $Q_{\mathsf{R}}$, we consider for each relation $D$, its corresponding synopsis $\mathtt{synop}$ to be the collection of $\{(\mathsf{pair}_{k},\mathbf{H}(\mathsf{pair}_{k},D),\mathtt{MF}_{k})\}_{k\geq 1}$, such that
-  • $\mathsf{pair}_{k}$ is a frequently queried attribute (pair) over $D$ found in $Q_{\mathsf{R}}$.
-  • $\mathbf{H}$ is the private bounding histogram for $\mathsf{pair}_{k}$ over $D$.
-  • $\mathtt{MF}_{k}$ represents a collection of privately overestimated join key MFs that corresponds to a group of tuples in $\mathbf{h}$, categorized by $\mathsf{pair}_{k}.A_{\mathsf{ft}}$.
 Theorem 4.3.
-Given $|\mathsf{pair}|=c$, $\epsilon,\delta>0$, the synopsis generation (Algorithm 1) is $(\hat{\epsilon},\hat{\delta})$-DP where $\hat{\epsilon}\leq 6\epsilon\sqrt{c\ln(1/\delta)}$, and $\hat{\delta}=(c+1)\delta$
-For space concern, we move the complete proof to Appendix A.
+[…]
+For space concern, we move complete proofs to the full version (Chenghong Wang, 2024).
 In a sketch, adding $\mathsf{Lap}^{+}$ (or $\mathsf{Lap}^{-}$) to a single bin is $(\epsilon,\delta)$-DP.
 By parallel and sequential composition, generating $\mathbf{{H}}$ is $(2\epsilon,2\delta)$-DP.
-Moreover, each report noisy max is $(\epsilon,0)$-DP, and by parallel composition, the generation of the entire MF table is also $(\epsilon,0)$-DP.
-In this way, we know that the generation of each $(\mathsf{pair}_{k},\mathbf{H},\mathtt{MF}_{k})$ is at most $(3\epsilon,\delta)$-DP.
+Moreover, each noisy max is $(\epsilon,0)$-DP, and by parallel composition, the generation of the entire MF table is also $(\epsilon,0)$-DP.
+In this way, we know that the generation of each $(\mathsf{pair}_{k},\mathbf{H},\mathtt{MF}_{k})$ is at most $(3\epsilon,2\delta)$-DP.
 Given there are in total $c$ such pairs, and thus the total privacy loss is subject to $c$-fold advanced composition (Dwork et al., 2014).
@@
 We say that one can perform transformations on released synopses without incurring extra privacy loss, per the post-processing theorem of DP (Dwork et al., 2014).
-Now, we briefly outline key synopsis transformations relevant to the SPECIAL design.
+Now, we outline the key synopsis transformation relevant to SPECIAL’s design.
 First, given any (2d) bounding histogram $\mathbf{H}(\mathsf{pair},D)$ with both $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}$ are non-empty, one can derive the (1d) bounding histograms, i.e.
@@
 This enables the creation of statistics on individual attributes, even when $A_{\mathsf{ft}}$ and $A_{\mathsf{j}}$ are not included as a standalone synopsis attribute.
-Moreover, it’s possible to derive relevant join key statistics even following a selection on the base relation.
-For example, given $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}\neq\emptyset$, and let $D^{\prime}\leftarrow\sigma_{A_{\mathsf{ft}}\in vals}(D)$, one can obtain the (1d) bounding histogram $\mathbf{H}(A_{\mathsf{j}},D^{\prime})$ by conducting a selective marginal sum of $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over bins of $A_{\mathsf{ft}}$ that intersect with $vals$.
+Moreover, it’s possible to derive relevant join key statistics following a selection on the base relation.
+For example, given $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}\neq\emptyset$, and let $D^{\prime}\leftarrow\sigma_{A_{\mathsf{ft}}\in vals}(D)$, one can obtain the (1d) bounding histogram $\mathbf{H}(A_{\mathsf{j}},D^{\prime})$ by conducting a selective marginal sum of $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over bins of $A_{\mathsf{ft}}$ that intersect with valsvals.
 Beyond bounding histograms, join key MFs over pre-filtered data can also be computed by
@@
 5. SPECIAL Primitives
-We now outline key secure database operations offered by SPECIAL.
-A major challenge in designing synopsis-assisted secure processing is the evident knowledge gap on how private synopses could potentially accelerate oblivious query processing.
-To bridge this gap, we explore various uses of synopses, such as creating private indexes (SPEidx 5.2) and designing compacted oblivious operations (SPEop 5.3).
-Moreover, as discussed earlier, joins are the most intensive operations.
-Hence, we extensively optimize join algorithms, synergizing both private indexing and compaction techniques to develop a novel, parallel-friendly oblivious join (§ 5.3).
-Another challenge in primitive design arises from ensuring lossless guarantees.
-To address this, we integrate mechanisms into our designs that pessimistically estimate selection cardinalities, indexing ranges, and complex join sizes, using a combination of available synopses and advanced upper bound techniques (Hertzschuch et al., 2021).
-For clarity, we’ll assume, WLOG., that all input relations below are of size $n$ and all 1d histograms contain $m$ bins.
+We next introduce the secure primitives in SPECIAL.
+One major challenge in designing these primitives is the knowledge gap on how private synopses can accelerate oblivious query processing, i.e., C-3.
+To address this, we explore various usage of synopses, including creating private indexes (SPEidx § 5.2) and designing compacted oblivious operations (SPEop § 5.3).
+Given that joins are the most resource-intensive operations, we optimize join algorithms by combining private indexing and compaction techniques to develop a novel, parallel-friendly oblivious join (§ 5.3).
+Another challenge is ensuring lossless processing, which we tackle by integrating mechanisms that pessimistically estimate selection cardinalities, indexing ranges, and join sizes using synopses (§ 4) and advanced upper bound techniques (Hertzschuch et al., 2021).
+For simplicity, we assume all input relations are of size $n$ and all 1D histograms contain $m$ bins.
 5.1. Basic Operations
-SPECIAL supports conventional fully-oblivious operators (Bater et al., 2017).
-In general, these operators function logically the same as non-private ones, but their execution transcripts are completely data independent, with results consistently padded to the worst-case maximum.
-Default data access SeqACC.
-By default, all query executions proceed with a standard data access, loading all data into a secure array using a sequential scan [^2] In the context of a secret-shared array $\langle D\rangle$, this step can be each server sequentially load the respective secret shares of $D$ into memory, readying them for MPC..
-Additionally, each loaded tuple gets a secret bit (initially ‘0’), $\mathsf{ret}$, used to mark tuple validity.
-SELECT.
-is the secure version of filter $\sigma_{p}(R)$.
-It conducts a linear scan over secure array $\langle R\rangle$ for relation $R$ and updates the $\mathsf{ret}$ bit: ‘1’ for tuples satisfying predicate $p$ and ‘0’ for others.
-Every tuple’s $\mathsf{ret}$ bit is updated, regardless of whether it matches the predicate.
-PROJECT. discards irrelevant attributes for relation stored in $\langle R\rangle$. Note that the secret bit $\mathsf{ret}$ cannot be discarded by a PROJECT.
-JOIN.
-implements the secure version of $\theta$-join, $R_{0}\Join_{\theta}R_{1}$.
-It processes two secure arrays $\langle R_{0}\rangle$ and $\langle R_{1}\rangle$, computing their cartesian product $\langle R_{0}\times R_{1}\rangle$ and then using SELECT to mark the joined tuples.
-It’s important to note that JOIN expands output, always padding it to the worst-case max (product of input sizes).
-COUNT, SUM, MIN/MAX are the aggregation primitives. The operators linearly scan a secure array and continuously update a secret-shared aggregation value after each tuple access.
-ORDER-BY, DISTINCT, GROUP-BY(AGG).
-are built on the oblivious sort primitive (Batcher, 1968).
-Specifically, ORDER-BY obliviously sorts a secure array according to a given attribute.
-DISTINCT sorts a secure array first, then performs a linear scan to identify unique tuples.
-For instance, among consecutively sorted identical tuples, it sets the $\mathsf{ret}$ bit of the last tuple to ‘1’ and the rest to ‘0’.
-GROUP-BY(AGG) starts by using DISTINCT to identify unique tuples.
+SPECIAL supports conventional fully-oblivious operators (Bater et al., 2017), which logically the same as non-private ones but with data-independent execution and worst-case padding for results.
+We briefly introduce these operations: (i) Default data access (SeqACC).
+By default, query execution begins with loading all data into a secure array via sequential scan.
+Each loaded tuple gets a secret bit $\mathsf{ret}$ (initially ‘0’), marking its validity; (ii) SELECT.
+This secure filter $\sigma_{p}(R)$ performs a linear scan over the secure array $\langle R\rangle$, updating the $\mathsf{ret}$ bit to ‘1’ for tuples satisfying predicate $p$ and ‘0’ for others; (iii) PROJECT.
+Removes irrelevant attributes from relation $\langle R\rangle$, but retains the $\mathsf{ret}$ bit; (iv) JOIN.
+Implements a secure $\theta$-join $R_{0}\Join_{\theta}R_{1}$ by computing the cartesian product $\langle R_{0}\times R_{1}\rangle$ and marking joined tuples with SELECT.
+The output is padded to the worst-case maximum size; (v) COUNT, SUM, MIN/MAX.
+These aggregation operators scan the secure array and update a secret-shared aggregation value for each tuple; (vi) ORDER-BY, DISTINCT, GROUP-BY(AGG).
+Built on the oblivious sort primitive (Batcher, 1968).
+ORDER-BY sorts the array by a given attribute.
+DISTINCT sorts and then identifies unique tuples, marking only the last in a sequence of identical tuples with $\mathsf{ret}=\text{`1'}$.
+GROUP-BY(AGG) first uses DISTINCT to find unique tuples.
 For each distinct tuple ($\mathsf{ret}$ set to ‘1’), it appends an aggregation value derived from the tuple and a dummy attribute (e.g., ‘-1’) for non-distinct tuples.
 5.2. SPECIAL Index SPEidx
-In conventional databases, indexes are powerful data structures that map attribute values to positions within a sorted array, enabling fast data access.
-However, generalizing such feature to SCA remains largely underexplored with existing solutions facing significant limitations such as high data loss (Roy Chowdhury et al., 2020), dependence on complex data structures filled with dummy tuples (Zhang et al., 2023; Bogatov et al., 2021), and limited query capabilities (i.e., linear queries across fixed attributes).
-Moreover, these methods only index base relations.
-SPEidx innovatively overcomes these issues by enabling the creation of lossless indexes directly on outsourced data or query intermediate results, without the need for additional structures or dummy injections.
+Existing index techniques in SCA have several drawbacks: loss of qualified data (Roy Chowdhury et al., 2020), reliance on intricate data structures with large overhead (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., 2023), and restricted query support (Zhang et al., 2023; Bogatov et al., 2021).
+Furthermore, all these techniques only support indexing the base relations.
+SPEidx offers a breakthrough by enabling the creation of lossless indexes directly on outsourced data and the query of intermediate results, eliminating the need for extra structures or storing dummy data.
 In general, SPEidx builds upon the typical indexing model that utilizes cumulative frequencies (CF) (Kraska et al., 2018).
-Specifically, given $D$ sorted by $A\in attr(D)$, all records $t\in D$ where $t.A=x$ can be indexed by the interval $[g(x-1),g(x)]$, where $g(x)=|\{t\mid t.A\leq x\}|$ is the CF function.
+Specifically, given $D$ sorted by $A\in attr(D)$, all records $t\in D$ where $t.A=x$ can be indexed by the interval [g(x1),g(x)][g(x-1),g(x)], where $g(x)=|\{t\mid t.A\leq x\}|$ is the CF function.
 For better illustration, we show an example index lookup in Figure 2: to get all records with an attribute value of 24, one may compute $[g(23),g(24)]=[217,248]$ and access the relevant data from the subset $D[217:248]$.
-To make this indexing method private and lossless, the key idea of SPEidx is to derive two noisy CF curves from synopses (e.g.
-bounding histograms).
+To make this indexing method private and lossless, the key idea of SPEidx is to derive two noisy CF curves from synopses (i.e., bounding histograms).
 One curve, $g^{+}(x)$, consistently overestimates $g(x)$, while the other, $g^{-}(x)$, consistently underestimates it.
-Then for any attribute value $x$, we can now derive a private interval $[g^{-}(x-1),g^{+}(x)]$ that losslessly indexes all desired records.
-For instance, as illustrated in Figure 2, the SPECIAL index might estimate the index range for attribute value 24 as $[g^{-}(23),g^{+}(24)]=[198,267]$.
-Figure 2. True index (left) vs. SPECIAL index (right) for $x=24$
+Then for any attribute value $x$, we can now derive a private interval $g^{-}(x-1),g^{+}(x)]$ that losslessly indexes all the desired records.
+As illustrated in Figure 2, the SPECIAL index might estimate the index range for attribute value 24 as $g^{-}(23),g^{+}(24)]=[198,267]$.
+Figure 2. True (left) vs. SPECIAL (right) index for $x=24$
 In what follows, we provide the formal explanations on how SPEidx derives indexes from DP synopses.
@@
 Definition 5.1 (SPECIAL index).
-Given $D$ sorted by $A$, the bounding histogram $\mathbf{H}(A,D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$, and assume $\mathbf{h}^{+}=(c^{+}_{1},...,c^{+}_{m})$, $\mathbf{h}^{-}=(c^{-}_{1},...,c^{-}_{m})$ partitions $dom(A)$ into $\{B_{1},...,B_{m}\}$.
-We say $\textsc{SPEidx}(A,D)=\{\mathsf{idx}_{i}=[\mathsf{lo}_{i},\mathsf{hi}_{i}]\}_{1\leq i\leq m}$ is the SPECIAL index of $D$ over $A$ with:
-  • $\forall~{}i\geq 1$, $\mathsf{hi}_{i}=\min(|D|,\sum^{i}_{k=1}c^{+}_{k})$.
+[…]
+[…]
   • $\mathsf{lo}_{1}=0$, and $\forall~i\geq 2$, $\mathsf{lo}_{i}=\sum^{i-1}_{k=1}\max(0,c^{-}_{k})$.
@@
 Here, SPEidx can derive $\mathbf{H}(A,D^{\prime})$ from $\mathbf{H}(A^{*},D)$ and subsequently build indexes on $D^{\prime}$.
-Importantly, since index creation is a post-processing activity using available DP synopses, it incurs no additional privacy loss.
+Importantly, since index creation is a post-processing procedure using available DP synopses, it incurs no additional privacy loss.
 Indexed store and fast data access IdxAcc.
-SPEidx enables the exploration of a new storage layout for organizing outsourced data, namely indexed datastore, that facilities private data access.
+SPEidx enables a new storage layout for outsourced data, namely indexed datastore.
 Specifically, by analyzing a representative workload $Q_{\mathsf{R}}$, one may identify the “hottest” attribute per base relation, sort them according to the “hottest” attribute, and then build indexes over the sorted data.
-This storage layout enables fast indexed access (IdxAcc) to retrieve a compact subset of data from the outsourced relations, thereby eliminating the need for a full table sequential scan (SeqAcc) potentially bypassing selection operations (§ 5.3).
+This storage layout enables fast indexed access (IdxAcc) to retrieve a compact subset of data from the outsourced relations, thereby eliminating the need for a full table sequential scan (SeqAcc) and can directly produce a compact input.
 We emphasize that the SPECIAL design does not require replicating the outsourced datastore to accommodate multiple query types (Bogatov et al., 2021; Zhang et al., 2023).
 However, creating compact replicas (e.g., column replicas (Huang et al., 2020) over frequently queried attributes) can be optionally employed to enhance query processing speed.
-Moreover, the generation of all aforementioned objects (indexed store and column replicas) requires only three primitives: projection, oblivious sorting, and SPEidx, all achieved without extra privacy loss.
-In other words, this implies that one can selectively adjust these objects to align with dynamic query workloads, without incurring any privacy costs.
+Moreover, the generation of all aforementioned objects (indexed store and column replicas) requires only three primitives: projection, oblivious sorting, and SPEidx.
+In other words, this implies that one can selectively adjust these objects to align with dynamic query workloads, without incurring extra privacy loss.
 5.3. SPECIAL Operators SPEop
-We introduce SPEop, a set of novel synopsis-assisted operators that maintain full obliviousness (e.g., operators’ execution cause no privacy loss) while enabling lossless compaction.
-To our knowledge, SPEop is the first primitive of its kind in any SCA system.
+We introduce SPEop, a set of novel synopsis-assisted operators that maintain full obliviousness, while enabling lossless compaction.
+To our knowledge, SPEop is the first primitive of its kind in any SCA.
 Oblivious compaction: OPAC.
-We first introduce the fundamental operation of oblivious compaction, a key component for many SPEop primitives.
+is a fundamental operation critical to other SPEop primitives.
 Given input $\langle R\rangle$, OPAC sorts it based on the secret bit $\mathsf{ret}$, moving tuples with $\mathsf{ret}=$ ‘1’ to the front.
@@
 SPECIAL selections: (OP)SELECT, (SP)SELECT, (DC)SELECT. Let $R$ to be a relation and $A\in attr(R)$, we now introduce three advanced selections that implements $\sigma_{A\in vals}(R)$.
-Algorithm 2 $\mathtt{CardEst}(\sigma_{A\in vals}(R),\mathtt{synop})$
-1:$\mathbf{h}=\emptyset$, $c=0$
-2:if $\mathbf{H}(A,R)\in\mathtt{synop}$ then $\mathbf{h}\leftarrow\mathbf{h}^{+}\in\mathbf{H}(A,R)$
-3:else if $\exists~{}\mathsf{pair}\in\mathtt{synop}$, s.t. $A\in\mathsf{pair}$ then
-4:     $\mathbf{h}\leftarrow$ marginal sum $\mathbf{h}^{+}\in\mathbf{H}(A,R)$ over $\left(\mathsf{pair}\setminus A\right)$.
-5:else return $cs=|R|$
-6:return $cs=\min(|R|,\sum_{i=1}^{m}c_{i}\in\mathbf{h}:\left(B_{m}\cap vals\neq\emptyset\right))$
 (OP)SELECT.
@@
 Moreover, as OPAC is fully oblivious and $cs$ is determined completely from post-processing over DP synopsis, thus, (OP)SELECT causes no privacy loss.
+Algorithm 2 $\mathtt{CardEst}(\sigma_{A\in vals}(R),\mathtt{synop})$
+[…]
+2:if $\mathbf{H}(A,R)\in\mathtt{synop}$ then $\mathbf{h}\leftarrow\mathbf{h}^{+}\in\mathbf{H}(A,R)$
+3:else if $\exists~\mathsf{pair}\in\mathtt{synop}$, s.t. $A\in\mathsf{pair}$ then
+4:  $\mathbf{h}\leftarrow$ marginal sum $\mathbf{h}^{+}\in\mathbf{H}(A,R)$ over $\left(\mathsf{pair}\setminus A\right)$.
+5:else return $cs=|R|$
+6:return $cs=\min(|R|,\sum_{i=1}^{m}c_{i}\in\mathbf{h}:\left(B_{m}\cap vals\neq\emptyset\right))$
 (SP)SELECT.
@@
 Specifically, in the second scan, (SP)SELECT internally maintains the last actual write position $\mathsf{idx}$ in $\langle R_{o}\rangle$.
-Then for every newly accessed tuple $\langle t\rangle$ in $\langle D\rangle$, a write action occurs on all tuples in $\langle R_{o}\rangle$.
-If $\langle t\rangle$ is selected, then an actual write is made that writes $\langle t\rangle$ to $\langle R_{o}[\mathsf{idx}+1]\rangle$ and a dummy write [^3] In the context of the secret-shared secure array $\langle\mathbf{a}\rangle$, a dummy write to $\langle\mathbf{a}[i]\rangle$ is simply a re-sharing of $\mathbf{a}[i]$ through secure protocols without changing its value.
-is made to elsewhere.
+Then for every newly accessed tuple $\langle t\rangle$ in $\langle R\rangle$, a write action occurs on all tuples in $\langle R_{o}\rangle$.
+If $\langle t\rangle$ is selected, then an actual write is made that writes $\langle t\rangle$ to $\langle R_{o}[\mathsf{idx}+1]\rangle$ and a dummy write is made to elsewhere.
 If not, dummy writes are made throughout $\langle R_{o}\rangle$.
-Similarly, (SP)SELECT does not yield additional privacy loss, and the complexity is linear in its input.
+We say that, in the context of the secret-shared secure array $\langle\mathbf{a}\rangle$, a dummy write to $\langle\mathbf{a}[i]\rangle$ is simply a re-sharing of $\mathbf{a}[i]$ through secure protocols without changing its value.
 (DC)SELECT.
-Finally, if the underlying data is already indexable on $A$, a direct selection over the data can be applied that bypasses secure computations.
-The operator simply looks up $\textsc{SPEidx}(A,R)$, and accesses $R[a,b]$, where $a=\min_{i}(\mathsf{idx}_{i}.\mathsf{lo}),b=\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})$, where $\mathsf{idx}_{i}$ dentoes the index in $\textsc{SPEidx}(A,R)$ where $B_{i}\cap vals\neq\emptyset$.
+Finally, if the underlying data is already indexable on $A$, a direct pre-fetch can be applied to avoid full table scan and compaction.
+The operator simply looks up $\textsc{SPEidx}(A,R)$, and accesses $R[a,b]$, where $a=\min_{i}(\mathsf{idx}_{i}.\mathsf{lo}),b=\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})$, and $\mathsf{idx}_{i}$ dentoes the index in $\textsc{SPEidx}(A,R)$ with bin $B_{i}\cap[a,b]\neq\emptyset$.
+A standard SELECT is then applied to $R[a,b]$.
 SPECIAL join: (MX)JOIN.
 We now introduce a novel MF-Index based oblivious join operation.
-The advancements of (MX)JOIN stand out in two key aspects compared to existing oblivious joins.
+The advancements of (MX)JOIN stand out in three aspects.
 First, compared to the standard JOIN, (MX)JOIN stands out for its ability to significantly compact the output size, coupled with a highly parallelizable fast processing mode.
-Second, existing oblivious joins with DP leakages typically require additional privacy budget to learn join sensitivity (Dong et al., 2022) or necessitate truncations on joined tuples (Wang et al., 2022; Bater et al., 2018).
+Second, existing DP oblivious joins typically require spending privacy budget (Dong et al., 2022) to learn join sensitivity (Dong et al., 2022) or necessitate truncation on joined tuples (Wang et al., 2022; Bater et al., 2018).
+(MX)JOIN eliminates this need.
 Moreover, (MX)JOIN is unique as the first oblivious join that enables lossless output compaction without extra privacy loss.
@@
 Algorithm 3 (MX)JOIN (base and pre-filtered relations)
-1:Input: relations $R_{0}$, $R_{1}$; join attribute $A_{\mathsf{j}}$.
+1:Input: relations $R_{0}$, $R_{1}$; join attribute $A_{\mathsf{j}}$; we consider synopses (histograms) of $A_{\mathsf{j}}$ are partitioned into bins $B_{1},...B_{m}$.
 2:if $\mathsf{MXReady}(R_{0},R_{1})==\mathsf{True}$ then  $\mathsf{BucketJoin}(R_{0},R_{1},A_{\mathsf{j}})$
@@
 9:  else assert “not applicable for (MX)JOIN”
-10:$\mathsf{BucketJoin}(R_{0},R_{1},A_{\mathsf{j}})$:
-11:for $i=1,2,...,m$ do
-12:     compute $O_{i}\leftarrow(R_{0}[\mathsf{idx}_{i}]\Join_{A_{\mathsf{j}}}R_{1}[\mathsf{idx}_{i}])$ via basic JOIN
-13:     $cs_{i}\leftarrow\min\left(\frac{|R_{0}[\mathsf{idx}_{i}]|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})},\frac{|R_{1}[\mathsf{idx}_{i}]|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}\right)\times{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})}\cdot{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}$
-14:     $R_{\mathsf{out}}\leftarrow R_{\mathsf{out}}\cup{\texttt{OPAC}}(O_{i},cs_{i})$
-15:return $R_{\mathsf{out}}$
-In general, (MX)JOIN can be applied to two types of data: the base and pre-filtered relations where the join key attribute is contained by the synopses.
+[…]
+11:for i=1,2,,mi=1,2,...,m do
+12:  $R^{(i)}_{0,1}\leftarrow\sigma_{A_{j}\in B_{i}}(R_{0,1})$ using (DC)SELECT/ $\textsc{SPEidx}(A_{\mathsf{j}},R_{0,1})$
+13:  compute $O_{i}\leftarrow(R^{(i)}_{0}\Join_{A_{\mathsf{j}}}R^{(i)}_{1})$ via standard JOIN
+14:  $cs_{i}\leftarrow\min\left(\frac{|R^{(i)}_{0}|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})},\frac{|R^{(i)}_{1}|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}\right)\times{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})}\cdot{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}$
+15:  $R_{\mathsf{out}}\leftarrow R_{\mathsf{out}}\cup{\texttt{OPAC}}(O_{i},cs_{i})$
+16:return $R_{\mathsf{out}}$
+In general, (MX)JOIN can be applied to two types of data: the base and pre-filtered relations where the join key attribute is included in $\mathtt{synop}$.
 Specifically, (MX)JOIN starts with computing the join key MFs (Alg 3:4) and constructing private indexes (Alg 3:5) for both inputs.
@@
 As bucket-wise operations are independent, the aforementioned steps lend themselves well to parallelized processing.
-For better illustration, we visualize (MX)JOIN’s processing flow in Figure 3.
-Figure 3. Example processing flow of (MX)JOIN.
-Since (MX)JOIN derives join compaction sizes completely from post-processing of DP synopses, it thus incurs no extra privacy loss.
-Moreover this design choice also eliminates the need for computing join sensitivities — a major hurdle in existing — as accurate join sensitivity estimation is hard and typically requires a significant privacy budget (Dong et al., 2022).
+As (MX)JOIN derives join compaction sizes completely from post-processing of DP synopses, it thus incurs no extra privacy loss.
 Additionally, the noisy MF bounds guarantee that compaction sizes are consistently overestimated, ensuring lossless compaction of join results.
 6. SPECIAL Planner
-A major limitation of existing DPSCA designs is sub-optimal execution planning.
-Without the ability to pre-estimate intermediate result sizes, traditional methods cannot identify the most efficient execution strategies.
-To address this, SPECIAL employs an innovative query planner that leverages available synopses for plan size estimation, enabling optimal execution planning.
-At a high level, our planner is modeled on the Selinger-style (cost-based) optimizer (Blasgen et al., 1981).
-It uses a bottom-up, dynamic programming approach to enumerate all equivalent secure execution plans for a given query and estimate their costs to select the optimal one.
-However, challenges in the design space still remain.
-First, the introduction of SPECIAL primitives can significantly impact the cost modeling for oblivious operations, rendering existing models (Bater et al., 2018; Liagouris et al., 2023) ineffective.
-Additionally, the equivalent plan search space is extensive, requiring further strategies to simplify the search space.
-To address these challenges, we first conduct a comprehensive analysis of the complexities of SPECIAL primitives and develop a new cost model (§ 6.1) for SPECIAL planner.
-Next, we design protocol-specific heuristics (§ 6.2) tailored for the SPECIAL planner to simplify the plan search space.
-Note that existing features of the Selinger-style optimizer, such as plan enumerations, are not within our contributions.
-For detailed understanding of these features, we encourage readers to explore the original IBM System R paper (Blasgen et al., 1981).
+Current DPSCA designs struggle with costly execution plans because they cannot pre-estimate query intermediate sizes, and thus unable to identify effective execution plans with minimized cost.
+SPECIAL overcomes this challenge by introducing a novel query planner that uses synopses for size estimation, and thus enabling both private and efficient SCA query planning.
+At a high level, our planner is modeled after the Selinger-style optimizer (Blasgen et al., 1981).
+It uses a bottom-up, dynamic programming approach to enumerate all equivalent plans for a given query, estimates their costs (heavily influenced by intermediate sizes) using available synopses, and selects the plan with minimal cost.
+The introduction of SPECIAL primitives significantly impacts cost modeling for oblivious operations, rendering existing models (Bater et al., 2018; Liagouris et al., 2023) inadequate.
+Furthermore, the design-space challenge of query planning persists, and the extensive equivalent plan search space necessitates strategies to simplify the search process.
+To address these, we first systematically analyze the complexities of SPECIAL primitives and develop a new cost model (§ 6.1).
+We then design protocol-specific heuristics (§ 6.2) tailored to our planner to narrow the search space.
 6.1. Cost Model
-Similar to the cost model proposed by Shrinkwrap (Bater et al., 2018), we view the cost of a secure execution plan cost as a combination of each operator’s I/O and secure evaluation costs.
-In addition, we formulate these costs as functions over their processed data sizes.
+We adopt the standard SCA cost framework (Bater et al., 2018) to develop SPECIAL’s cost model, viewing the cost of a secure execution plan as the sum of each operator’s I/O and secure evaluation costs.
 Specifically, given a plan with $\ell$ operators, $\mathsf{op}_{1},...,\mathsf{op}_{\ell}$, and let $\mathbf{I}=\{{I}_{1},...,{I}_{\ell}\}$, $\mathbf{O}=\{{O}_{1},...,{O}_{\ell}\}$, to be the input and output sizes of each operators.
-The plan cost is formulated as:
+The plan cost is:
 (5) $\mathsf{Cost}=\textstyle\sum_{i=1}^{\ell}C^{\mathsf{op}_{i}}_{\mathsf{in}}(I_{i})+C_{\mathsf{eval}}^{\mathsf{op}_{i}}(I_{i})+C^{\mathsf{op}_{i}}_{\mathsf{out}}(O_{i})$
 Here, $C_{\mathsf{in}}$ represents the data access cost (input I/O), primarily capturing the expenses when moving data from persistent storage to an in-memory secure array.
-If the accessed data has already loaded into memory, $C_{\mathsf{in}}$ becomes negligible.
 $C_{\mathsf{out}}$ denotes the output I/O cost, modeling the expenses when writing operator results into output arrays.
-$C_{\mathsf{eval}}$ accounts for the secure computing expenses for evaluating each operators, typically constituting the dominant cost.
+$C_{\mathsf{eval}}$ accounts for the secure computing cost for evaluating an operator, typically constituting the dominant cost.
 Note that, in practice, the exact formulas for $C_{\mathsf{in}},C_{\mathsf{out}}$, and $C_{\mathsf{eval}}$ can vary depending on the specific secure protocol employed (garbled circuits (Yao, 1986), secret sharing (Micali et al., 1987), etc.) as well as the particular hardware configurations in use.
-Nonetheless, the understanding of the asymptotic costs is adequate for comprehending the principles of our query planning and optimization strategies (Liagouris et al., 2023; Bater et al., 2018).
+Nonetheless, the understanding of the asymptotic costs is adequate for comprehending the principles of SCA query planning and optimization strategies (Liagouris et al., 2023; Bater et al., 2018).
 In what follows, we provide detailed analysis on the asymptotic costs for each SPECIAL operator.
@@
 (MX)JOIN $O(n)$ $O(n^{2})^{*}$ $\mathsf{mf\_bound}$
-Assuming the maximum size of the indexed buckets is bounded by $O(\frac{n}{\log n})$.
+Assuming the max size of the indexed buckets is bounded by $O(\frac{n}{\log n})$.
 Oblivious sorting and compaction.
-Oblivious sorting and compaction (OPAC) are crucial for many SPEop operators, so we will begin our analysis with these two primitives.
 While oblivious sorting algorithms with optimal $O(n\log n)$ complexity exist, they often necessitate either impractically large constants (Ajtai et al., 1983; Goodrich, 2014) or client-side memory (Asharov et al., 2020), both do not fit with SCA scenario.
 Consequently, we will consider the well-established bitonic sorting based implementation for oblivious sort, which come with $O(n\log^{2}n)$ complexity.
-Nonetheless, efficient OPAC implementations with $O(n\log n)$ complexity remain achievable (Sasy et al., 2022)
+Nonetheless, efficient OPAC implementations with $O(n\log n)$ complexity remain achievable (Sasy et al., 2022).
 Projection, grouping and aggregation.
-The PROJECT accesses private relations and discards unnecessary columns independently on each server, without secure computations.
+The PROJECT accesses private relations and discards unnecessary columns independently on each server, which is naturally oblivious.
 Thus, I/O costs dominate this operation, with both input and output costs bounded by $O(n)$.
-The costs of ORDER-BY, DISTINCT, and GROUP-BY are primarily due to oblivious sorting, resulting in a complexity of $O(n\log^{2}n)$.
+The costs of ORDER-BY, DISTINCT, and GROUP-BY are dominated by oblivious sorting, resulting in a complexity of $O(n\log^{2}n)$.
 Additionally, as these operators do not reduce output sizes, both $C_{\mathsf{in}}$ and $C_{\mathsf{out}}$ are bounded by $O(n)$.
@@
 (DC)SELECT is the most efficient selection, though it requires indexable input data.
-It simply loads indexed data into secure arrays without secure computations or result write-backs.
-As such, both $C_{\mathsf{evals}}$ and $C_{\mathsf{out}}$ are negligible, with $C_{\mathsf{in}}$ within $\mathsf{idx\_bound}=O\left(\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})-\min_{i}(\mathsf{idx}_{i}.\mathsf{lo})\right)$.
+All costs are directly related to the size of the indexed data, so $C_{\mathsf{in}}$, $C_{\mathsf{evals}}$ and $C_{\mathsf{out}}$ are all bounded by $\mathsf{idx\_bound}=O\left(\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})-\min_{i}(\mathsf{idx}_{i}.\mathsf{lo})\right)$.
 Here, $\mathsf{idx}_{i}$ are indexed regions that intersect with selection conditions.
@@
 Both JOIN and (MX)JOIN have $O(n)$ data access costs, but differ in $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$.
-JOIN, using a Cartesian product approach, doesn’t reduce output size, so both $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$ are within $O(n^{2})$.
+JOIN, conducting a Cartesian product for two input tables, has $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$ both bounded by $O(n^{2})$.
 Compared to JOIN, in the worst-case scenario where the join keys follow a highly biased distribution, i.e.
@@
 $O(\frac{n^{2}}{\log n})$.
-Additionally, the output cost is lowered from $O(n^{2})$ to the sum of per-bucket MF upper bounds (Alg 3:11), which can be substantially less if the join key MFs are relatively low.
+Additionally, the output cost is lowered from $O(n^{2})$ to the sum of per-bucket MF upper bounds (Alg 3:11), which can be substantially less if the join key MFs are low.
 6.2. Heuristics
@@
 Filter push down.
-Filter pushdown is a widely used optimization technique in conventional databases, where it involves moving the selection operation to the earliest stage in the data processing pipeline, such as directly to the base relation.
-This reduces the amount of data that needs to be loaded, transferred, and processed, by following operators especially complex joins.
-However, in SCA designs, data obliviousness typically necessitates padding selection sizes to the worst-case scenario, rendering filter pushdown ineffective (Liagouris et al., 2023; Eskandarian and Zaharia, 2017; Zheng et al., 2017; Poddar et al., 2021).
-With SPEop selections, the selection outputs can be significantly compacted, making filter pushdown effective again.
-Therefore, we include filter push down as one of the optimization heuristic for SPEplan.
+is a common query planning optimization, moves selection operations to the earliest possible stage to reduce data processed by subsequent operations.
+In conventional SCAs, data obliviousness often requires padding selection sizes, making filter pushdown ineffective (Liagouris et al., 2023; Eskandarian and Zaharia, 2017; Zheng et al., 2017; Poddar et al., 2021).
+However, SPECIAL’s innovative selection methods enable compacting selections to approximate true cardinalities without compromising privacy, restoring the effectiveness of filter pushdown.
+Hence, we include filter push down as one of the optimization heuristic for SPEplan.
 H-2.
 Predicates fusion.
-Let $R$ to be any relation, $A_{1},A_{2},...,A_{k}\subseteq attr(D)$, and $\mathbf{v}=\{v_{1},v_{2},...,v_{k}\}$.
+Let $R$ to be any relation, $A_{1},A_{2},...,A_{k}\subseteq attr(R)$, and $\mathbf{v}=\{v_{1},v_{2},...,v_{k}\}$.
 We say that for multiple selection over $R$ such that $\sigma_{A_{1}\in v_{1}}(...\sigma_{A_{k}\in v_{k}}(R))$, one can always fuse them into one selection $\sigma_{\mathbf{A}\subset\mathbf{v}}(R)$.
@@
 7. Evaluation
-In this section, we present evaluation results of our proposed framework. Specifically, we address the following questions:
-  • Question-1: Does SPECIAL offer efficiency advantages over current SCA approaches? To what extent can SPECIAL enhance efficiency compared to existing state-of-the-arts (SOTAs)?
-  • Question-2: For the DP-based SPECIAL design, is there a trade-off between privacy and efficiency? Can we adjust privacy levels to achieve various efficiency objectives?
-  • Question-3: Can SPECIAL scale complex analytical (e.g. multi-way join) queries to large-scale (multi-million rows) datasets?
+In this section, we present evaluation results of our proposed framework.
+Specifically, we address the following questions: Question-1: Does SPECIAL offer efficiency and privacy advantages over existing SCAs?
+Question-2: For SPECIAL design, is there a trade-off between privacy guarantees and system efficiency?
+How different synopses scenarios affect system performance?
+Question-3: Can SPECIAL scale complex analytical (e.g.
+multi-way join) queries to large-scale (multi-million rows) datasets?
 7.1. Experimental setups
-We briefly outline our experimental setup, covering the baseline systems, prototype SPECIAL implementation, workloads, datasets, and default configurations.
 Baseline systems and SPECIAL prototype.
-We compare SPECIAL with two baseline systems: Shrinkwrap (Bater et al., 2018), the SOTA DPSCA, and SMCQL (Bater et al., 2017), a typical SCA system using exhaustive padding for leakage hiding, also used as a baseline for Shrinkwrap.
+We compare SPECIAL with two baseline systems: Shrinkwrap (Bater et al., 2018), the SOTA DPSCA, and SMCQL (Bater et al., 2017) (also used as a baseline for Shrinkwrap).
 For consistency, we consider the same circuit-model implementations for both baseline systems and the SPECIAL prototype.
 While some works (Poddar et al., 2021; Liagouris et al., 2023) similar to SMCQL use exhaustive padding but improve efficiency through protocol-level optimizations, we exclude them from our benchmarks for fair comparison concerns.
-However, it’s important to note that these optimizations are orthogonal to both the baseline systems and our designs.
-We re-implement important query features for the two baseline systems, and built SPECIAL using the same MPC package, specifically the sh2pc-0.2.2 protocol within EMPtoolkit-0.2.5.
-Since managing DP synopses and generating indexes do not involve secure computations, thus we implement the related features using Python (3.7) scripts.
+We re-implemented key query features for the two baseline systems and built SPECIAL using the same MPC package ( EMPtoolkit-0.2.5).
 All implementations are open-sourced (Lovingmage, 2024).
 Datasets and workloads.
-Although SMCQL and Shrinkwrap use the HealthLNK dataset, it is not public data and thus not accessible to the broader research community.
-Hence, we developed a new benchmark based on a publicly available dataset, and mimic Shrinkwrap’s workload to create a set of testing queries.
-Specifically, we use the Czech Financial Dataset (Rel, [n.d.]), an anonymized Czech bank transaction dataset for loan applications.
+We developed two open benchmarks.
+The first reproduces the HealthLNK benchmark used by Shrinkwrap and SMCQL which simulates a real-world scenario where medical researchers want to perform secure analytics across multiple cohorts’ sensitive data.
+We use an open schema (SMCQL, [n.d.]) to generate a scalable synthetic dataset with 4 tables,  [^2] K rows and 24 columns.
+The benchmark involves four multi-cohort medical study queries, identical to those used in Shrinkwrap and SMCQL.
+Our second benchmark simulates secure collaborative analytics within the financial sector.
+Imagine multiple banks needing to analyze their combined, private data to study loan and financial statistics—all without compromising sensitive customer information.
+We use the anonymized Czech Financial Dataset (Rel, [n.d.]) for this, assuming each entry represents data owned by a different bank or financial organization unable to directly share information.
 This dataset comprises 8 relational tables with a total of 55 columns and 1.1 million rows.
-For scaling experiments, we use the raw data and schema to generate synthetic data with up to 10 million rows.
-Shrinkwrap’s benchmark evaluates four queries: a range query, two 2 binary joins, and a 3-way join.
-Mirroring this approach, we designed eight query workloads based on the Financial data.
-All workloads are selection-projection-join (SPJ) queries followed by statistical aggregations such as count, count distinct, and group by.
-The workloads range from simple linear queries to complex multi-way joins.
-Due to space limitations, a brief summary of the workloads is provided in Table 2, with the full specification in Appendix B.
+For testing workloads, we design eight query workloads, ranging from simple linear queries to complex multi-way join-aggregation queries.
+A brief summary of the workloads is provided in Table 2.
+Default configurations.
+For SPECIAL, we employ a per-table privacy budget allocation strategy.
+Each table is allocated a one-time privacy budget of $\epsilon=1.5$ and $\delta=0.00005$ for synopses generation, and is evenly distributed across all DP synopses.
+For Shrinkwrap, we adopt their default per-query privacy allocation, assigning a privacy budget of $\epsilon=1.5$ and $\delta=0.00005$ to each query, as outlined in (Bater et al., 2018).
+It is important to note that this configuration means Shrinkwrap will not offer guarantees on the bounded privacy loss across multiple queries.
+For the HealthLNK benchmark, we use Dosage and Aspirin as representative workloads, and for the Financial benchmark, we use FQ2, FQ4, and FQ8.
+Unless further elaborated, these workloads will also serve as default testing queries for our evaluation.
+For all equal-width histograms generated in SPECIAL, we configure them to have at most 8 bins.
+Moreover, for baseline systems, as they do not have join ordering optimizations, thus we will assume a random join order for them.
+We conduct all experiments on bare-metal Mac machines with M2 Max CPUs and 96GB unified memory.
 Table 2. Query workloads
-Query No. Type Description
-Q1 Linear Point query with single predicate.
-Q2 Linear Range query with conjunctive predicates.
-Q3 Binary Join Non-expanding binary join.
-Q4 Binary Join Expanding binary join.
-Q5 Multi Join 3-way mixed non-expanding and expanding joins.
-Q6 Multi Join 3-way all expanding joins.
-Q7 Multi Join 4-way mixed non-expanding and expanding joins.
-Q8 Multi Join 5-way mixed non-expanding and expanding joins.
-Default configurations.
-For SPECIAL, each relational table is assigned a fixed privacy budget with $\epsilon=1.5$ and $\delta=0.00005$ to release synopses.
-This setup ensures that the total privacy loss per table does not exceed $\epsilon=1.5$.
-For all equal-width histograms generated in SPECIAL, we configure them to have at most 8 bins.
-For Shrinkwrap, we use a privacy budget of $\epsilon=1.5$ and $\delta=0.00005$ for each query processing, as specified in (Bater et al., 2018).
-However, Shrinkwrap does not provide bounded privacy loss guarantees, so the privacy guarantee of Shrinkwrap will be no better than SPECIAL.
-We conduct all experiments on bare-metal Mac machines with M2 Max CPUs and 96GB unified memory.
-In addition, we consider Q2, Q4, Q8 are default testing queries in each query type, and are used as the representative workload to guide synopses.
-Moreover, for baseline systems, as they do not have join ordering optimizations, thus we will assume a random join order for them.
+Bench. Query Type Description
+HealthLNK Dosage Study Binary Join Expanding binary join.
+Comorbidity Binary Join Non-expanding binary join.
+Aspirin count Multi Join 3 way mixed join
+3 Join Aspirin Multi Join 4 way mixed join
+Financial FQ1 Linear Point query.
+FQ2 Linear Range query.
+FQ3 Binary Join Non-expanding binary join.
+FQ4 Binary Join Expanding binary join.
+FQ5 Multi Join 3-way mixed joins.
+FQ6 Multi Join 3-way all expanding joins.
+FQ7 Multi Join 4-way mixed joins.
+FQ8 Multi Join 5-way mixed joins.
 Figure 3.
-In-depth comparisons in execution plans: (i) The exhaustive padding in SMCQL can lead to significant memory blowup; (ii) Both SMCQL and Shrinkwrap suffer from suboptimal join ordering; (iii) Although Shrinkwrap reduces intermediate sizes, it still requires substantial memory to materialize the Cartesian-style standard oblivious join; (iv) SPECIAL effectively identifies optimal join orders and significantly reduces intermediate materialization sizes; (v) The IdxAcc method in SPECIAL not only bypasses secure selections but also directly reduces input I/O costs.
-7.2. End-to-end performance comparisons
-To address Question-1, we first conduct an end-to-end performance comparison of SPECIAL, Shrinkwrap, and SMCQL across the full benchmark workloads (Q1-Q8).
-We use two typical performance metrics, query execution time and memory usage in our evaluations.
-The results are summarized in Figure 5 and Figure 6.
-Due to significant memory consumption, SMCQL cannot complete the full benchmark and stops at Q4.
-Therefore, we only include full statistics up to Q3, and for Q4, we provide a projected performance evaluation using 10% of the input data.
+In-depth comparisons in execution plans: (i) The exhaustive padding in SMCQL can lead to significant memory blowup; (ii) Both SMCQL and Shrinkwrap suffer from unoptimized join ordering; (iii) Although Shrinkwrap reduces intermediate sizes, it still requires substantial memory to materialize join outputs; (iv) SPECIAL can identify efficient join execution orders to reduce intermediate sizes; (v) IdxAcc can significantly reduce input I/O costs.
+7.2. End-to-end comparisons
+To address Question-1, we first conduct an end-to-end performance comparison of SPECIAL, Shrinkwrap, and SMCQL across all benchmark workloads.
+The results are summarized in Figure 45.
+We cannot complete full benchmark for SMCQL due to high memory cost, so we project evaluations for FQ4 (using 10% data) and omit results for other complex workloads.
 Figure 4. End-to-end comparison: query latency
@@
 Observation 1.
-SPECIAL significantly outperforms all baseline systems in query latency, achieving performance gains of up to 3618.3×3618.3$\times$3618.3 × for linear queries, 114×114$\times$114 × for binary joins, and 80.3×80.3$\times$80.3 × for multi-joins.
-Figure 5 shows significant performance gains of SPECIAL in contrast to Shrinkwrap and SMCQL.
-For linear queries, SPECIAL exhibits substantial performance gains, achieving up to 3618.3×3618.3$\times$3618.3 × improvement (Q2), primarily due to its index-based fast data access.
-This approach not only reduces large I/O costs but also eliminates the need for secure computation by enabling direct data retrieval via DP indexes.
-We observe that in binary joins, SPECIAL has a less pronounced advantage over Shrinkwrap.
-This is because Shrinkwrap do not have sub-optimal planning issue when dealing with binary join—as there is only one join order.
-Hence, the performance gains for SPECIAL in Q3 and Q4 stem solely from fast data access and the efficient (MX)Join method.
-However, for more complex multi-way joins, SPECIAL’s advantage becomes more pronounced again due to its ability to select the optimal join order.
-For instance, Shrinkwrap shows more than 80×80$\times$80 × slow down in query latency than SPECIAL in group Q7.
+SPECIAL outperforms Shrinkwrap and SMCQL in query latency across all benchmarks, reaching up to $3618.3\times$ for linear queries, $114\times$ for binary joins, and $80.3\times$ for multi-joins.
+Figure 4 shows the comparison results in query time.
+First, SPECIAL shows significant speedups for linear queries, reaching up to $3618.3\times$ (FQ2).
+This large performance gain is mainly attributed to its index-based fast data access.
+By directly fetching private data through DP indexes, SPECIAL eliminate substantial I/O costs (e.g., sequential reads) and bypass the need for secure computations (e.g., oblivious filter).
+Second, we observe that in binary joins, SPECIAL has a less pronounced advantage over Shrinkwrap.
+This is because binary joins have a single join order, eliminating the potential for join ordering problems (where different join orders lead to significantly different performance).
+Consequently, even though Shrinkwrap does not optimize join orders, it doesn’t experience efficiency losses in this scenario.
+However, for more complex multi-way joins, SPECIAL’s advantage becomes more pronounced again.
+For instance, more than $80\times$ speedup in FQ7.
+This is because SPECIAL can pre-select efficient join orders before runtime.
 Observation 2.
@@
 Second, SPECIAL’s capability to identify optimal execution plans significantly reduces total intermediate sizes, which is particularly beneficial for complex joins that suffer from sub-optimal or exhaustive padding in other systems.
-To better understand the substantial improvements SPECIAL achieves—for instance, up to 928.2×928.2$\times$928.2 × over Shrinkwrap and more than $10^{5}\times$ over SMCQL—we will zoom into a specific query, Q6, and compare the detailed execution plans of the three systems.
-The choice of Q6 is strategic because its complexity sufficiently highlights the differences in execution plans, yet it remains simple enough for clear visual representation.
-Our comparison features four execution plans: a plaintext optimal plan derived directly from the logical plan using actual cardinalities, illustrating the ground truth optimal execution; a hypothetical SMCQL plan (since Q6 cannot be completed by SMCQL so we project the cardinalities); and two actual execution plans from our experiments with Shrinkwrap and SPECIAL.
+To better understand the substantial improvements SPECIAL achieves—for instance, up to $928.2\times$ over Shrinkwrap and more than $10^{5}\times$ over SMCQL—we will zoom into a specific query, FQ6, and compare the detailed execution plans of the three systems.
+The choice of FQ6 is strategic because its complexity sufficiently highlights the differences in execution plans, yet it remains simple enough for clear visual representation.
+Our comparison features four execution plans: a plaintext optimal plan, illustrating the ground truth optimal execution; a hypothetical SMCQL plan (with projected cardinalities); and two actual execution plans from our experiments with Shrinkwrap and SPECIAL.
 The detailed comparisons and observations are summarized in Figure 3.
-7.3. Privacy comparisons
-Continue to address Question-1, we now compare the privacy guarantees between Shrinkwrap and SPECIAL, with the results detailed in Figure 7.
-As privacy loss is independent of query types, our analysis will focus on comparing the privacy composition curves under continual query answering for both Shrinkwrap and SPECIAL.
-We consider two privacy composition models: advanced composition (Adv.)(Dwork et al., 2014) and concentrated DP composition (CDP)(Bun and Steinke, 2016).
+To continue address Question-1, we now compare the privacy guarantees of SPECIAL with the baseline systems (Figure 6).
+Specifically, we focus on comparing cumulative privacy loss in multiple query answering, w.r.t.
+two composition models: advanced composition (Adv.)(Dwork et al., 2014) and concentrated composition (CDP.)(Bun and Steinke, 2016).
 Figure 6. End-to-end comparison privacy loss.
 Observation 3.
-Under continual query answering, SPECIAL demonstrates significantly lower privacy loss compared to Shrinkwrap, achieving up to 89.01×89.01$\times$89.01 × and 38.91×38.91$\times$38.91 × improvements in the Adv.
+Under continual query answering, SPECIAL demonstrates significantly lower privacy loss compared to Shrinkwrap, achieving up to $89.01\times$ and $38.91\times$ improvements in the Adv.
 and CDP modes, respectively.
@@
 As such, SPECIAL demonstrate significant improve in privacy guarantees towards SOTA DPSCA.
+Even when compared to standard SCA (e.g., SMCQL) with no privacy loss due to exhaustive padding, our system incurs only a small and fixed privacy cost (e.g.
+$\epsilon=1.5$ per table) while delivering substantial performance gains.
 7.3. Privacy efficiency tradeoffs
-We address Question-2 by evaluating SPECIAL at various privacy levels and observing the performance impacts.
-Specifically, we maintain $\delta$ constant while varying $\epsilon$ from $0.1$ to $10$ and assess the performance across default testing queries (Q2, Q4, Q8).
-The results are shown in Figure 8 and Figure 9.
-Figure 8. Privacy vs. Performance (query time)
-Figure 9. Privacy vs. Performance (memory)
+We address Question-2 by evaluating SPECIAL at various privacy levels.
+Specifically, we maintain $\delta$ constant while varying $\epsilon$ from $0.1$ to $10$ and assess the performance across default testing queries.
+The results are shown in Figure 7.
+Figure 7. Privacy vs. Performance trade-offs
 Observation 4.
@@
 Conversely, if performance is the priority, setting $\epsilon$ near 1 is typically sufficient.
-Observation 5.
-Different queries exhibit varying performance sensitivities to changes in privacy levels.
-Multi-join queries are significantly affected when privacy levels change.
-For instance, the performance overhead of Q8 at $\epsilon=0.1$ is more than 3×3$\times$3 × higher than at $\epsilon=1$.
-In contrast, linear query (Q2) and binary join (Q4) overhead increases to only about 1.5×1.5$\times$1.5 × under the same privacy settings.
-This difference stems from the error amplification inherent in multi-join operations, which frequently reuse synopses and require multiple cardinality estimations.
-Therefore, practitioners dealing with complex multi-join workloads may need careful fine-tune privacy parameters, as performance impacts can be substantial.
+7.4. Synopses impacts micro benchmarks
+We continue to address Question-2 to explore how synopses scenarios may affect SPECIAL’s performance.
+Specifically, we study two key settings: (i) How different bin numbers (BinNum) in synopses can impact the efficiency of (IDX)JOIN, and (ii) how synopsis coverage levels for a single query can affect its overall execution.
+We will conduct micro-benchmarks for a thorough investigation.
+Note that simulate different synopses scenario on both HealthLNK and Financial benchmarks can be challenging (e.g., joins typically occur on the same key, so it is hard to simulate partial coverage), hence, to better control experimental variables and accurately assess impacts, we will now use synthetic data and workloads.
+Figure 8. BinNum experiments.
+We first study how the BinNums impact the performance of (IDX)JOIN.
+To study this, we synthesize two join queries on fixed input data, generate join key synopses with varying bin numbers (2 to 64), and measure the performance of (IDX)JOIN in processing the queries.
+The results are shown in Figure 8.
+Observation 6.
+The running time of (IDX)JOIN initially decreases but then increases as the BinNum grows.
+The memory usage consistently increases.
+(IDX)JOIN partitions larger joins into smaller sub-joins, but since we rely on DP indexes for partitioning, each sub-join inevitably includes additional dummy data.
+This increased plan size directly translates to higher memory usage and will grow when BinNum increases (more noises in DP indexes).
+On the other hand, partitioning large joins into smaller ones can enhance join efficiency, which is why we initially observe a decrease in execution time as BinNums increase (e.g., 2 to 8).
+However, the trade-off arises when the number of bins becomes excessive (e.g., >8>8).
+The overhead of handling the increased dummy data starts to outweigh the benefits gained from partitioning.
+At this point, the performance improvement plateaus and starts to degrade as the system struggles with the inflated plan sizes.
+Table 3. Synopses coverage experiments.
+Synopses Coverage Time (ms) Improv. Mem. (bytes) Improv.
+No coveragee 1831703 baseline 129048576 baseline
+Join Key
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+* We synthesize a random selectivity between (0, 0.33) for each filter operation.
+We now study how synopses coverage impacts query performance.
+We synthesize and test a 4-way join query under two controlled scenarios: (i) JK coverage: We focus on varying the level of JK coverage, starting from 1 out of 3 JKs to full coverage, while ensuring no filter synopses are present.
+We then measure how this impacts query performance; and (ii) Filter coverage: We maintain full JK coverage and change the coverage of filter synopses on the query’s input tables, ranging from 1 out of 4 inputs to full coverage.
+This allows us to examine the isolated effect of filter synopsis coverage on performance.
+Results are in Table 3.
+[…]
 7.5. Scaling experiments
-To address Question-3, we stress SPECIAL with scaled workloads by duplicating the raw financial dataset to sizes of 2×2$\times$2 ×, 4×4$\times$4 ×, and 8×8$\times$8 ×.
-We then evaluate the default testing queries on these scaled datasets to assess SPECIAL’s efficiency.
-The results are shown in Figure 10.
-Figure 10. End-to-end comparison privacy loss.
-Observation 6.
-SPECIAL demonstrates significant potential to scale up to multi-million records, even when handling complex 5-way joins.
-Figure 10 shows SPECIAL’s effective scaling: up to 8×8$\times$8 × data for linear queries and binary joins, and up to 4×4$\times$4 × data for complex 5-way joins like Q8.
-For instance, Q2 can be completed within 290ms under 8×8$\times$8 ×, and in fact, since selection is bypassed (due to index access), thus the cost is mainly on I/O costs.
-Q4 finishes in 289 minutes at the same scale 8×8$\times$8 ×, while the more complex 5-way join Q8 takes less than 280 minutes for 4×4$\times$4 × data.
-As a reference, Shrinkwrap would require over 1035 minutes to complete Q8 even with unscaled data (Figure 5).
-This result can, from another angle, showcase the significant performance improvements of SPECIAL over current SOTA DPSCA designs.
-8. Discussion
-In this section, we discuss potential extensions to accommodate broader data models and possible adaptations of SPECIAL designs to other secure primitives.
-Supporting growing data.
-In general SPECIAL considers read-intensive analytical workloads and and a static data model, which is the same as current SOTA SCA efforts (Poddar et al., 2021; Bater et al., 2018; Liagouris et al., 2023; Volgushev et al., 2019).
-In this section, we will briefly discuss how SPECIAL could be extended to support the growing data model where new records are continuously added to the logical database (Wang et al., 2022, 2021; Zhang et al., 2023).
-One possible strategy involves segmenting the large logical database into discrete, non-overlapping time intervals, each represented by its own database with independent synopses.
-Specifically, for new incoming data, SPECIAL will initially cache it until a designated time interval concludes.
-After this period, SPECIAL will batch all the cached data into a new database corresponding to the past time interval and release synopses for this database.
-As the data within these intervals is disjoint, queries can be easily partitioned and independently optimized and executed across corresponding databases in parallel, before being aggregated.
-Note that prior to releasing new synopses, all cached data will be processed using standard SCA processing (e.g.
-SMCQL).
-Since each synopses release covers only newly inserted data, the total privacy loss adheres to parallel composition (Dwork et al., 2014), and thus the bounded privacy loss still hold.
-Adaptability to other secure primitives.
-Beyond MPC-based SCA, the key design insights from SPECIAL are transferable to other secure outsourced databases such as the TEE-based analytical systems (Qiu et al., [n.d.]; Qin et al., 2022; Eskandarian and Zaharia, 2017; Zheng et al., 2017) or searchable encryption based private key-value store (Amjad et al., 2021; Groce et al., 2019; Chen et al., 2018).
-This is because SPECIAL’s core components, including synopses management, indexing, query planning, and compaction estimation, can all be executed without relying on secure computation primitives.
-For example, in TEE databases, one may conduct a similar query planning outside the TEE, with the optimal plan then executed within the TEE for fast oblivious processing.
-Similarly, SPECIAL’s design principles could be applied to searchable encryption to construct private indexes in advance on sorted data, enable fast and private information retrieval.
+To address Question-3, we stress SPECIAL with two types of scaled workloads: (i) scaled data: we duplicate the raw dataset to sizes of $2\times$, $4\times$, and $8\times$ and evaluate default testing queries; (ii) scaled query complexity: We use standard inputs, but simulate complex multi joins (up to 9-way) by chaining together multiple join workloads.
+The results are shown in Figure 9 and Table 4.
+Figure 9. Scaling data experiments.
+Observation 8.
+SPECIAL shows large potential to scale up to multi-million data, even for complex 5-way joins.
+Figure 9 shows SPECIAL’s effective scaling: up to $8\times$ data for linear queries and binary joins, and up to $4\times$ data for complex 5-way joins like Q8.
+For instance, Q2 can be completed within 290ms under $8\times$, and in fact, since selection is bypassed (due to index access), thus the cost is mainly on I/O costs.
+Q4 finishes in 289 minutes at the same scale $8\times$, while the more complex 5-way join Q8 takes less than 280 minutes for $4\times$ data.
+As a reference, Shrinkwrap would require over 1035 minutes to complete Q8 even with unscaled data.
+[…]
+Table 4. Scaling query complexity experiments
+Join scale Query time (ms) Improv. Memory (bytes) Improv.
+Shrinkwrap Q6 58342782 baseline 3354165360 baseline
+[…]
+[…]
+[…]
 8. Related Work
@@
 Two main approaches exist for designing MPC-based SCA systems.
-The first is peer-to-peer paradigm (Aggarwal et al., 2005; Mouris et al., 2024; Poddar et al., 2021; Volgushev et al., 2019; Wang and Yi, 2021), where the goal is to improve efficiency by decomposing analytical queries and pushing them to data owners, so that they can either directly process in clear or running MPCs across a smaller group of parties.
-Unfortunately, this approach burdens data owners with heavy computational overheads, particularly for complex operations like joins.
-Moreover, given real-world data owners often lack robust computing resources and reliable service capabilities, these methods are difficult to scale and promise consistent SCA services to external analysts.
-On the other hand, the server-aided-MPC model (Wang et al., 2021; Bater et al., 2018, 2020; Wang et al., 2022; Zhang et al., 2023; Roy Chowdhury et al., 2020; Kamara et al., 2011; Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023; Mohassel and Zhang, 2017) offers a more practical solution.
-It allows data owners to outsource both expensive MPC computations and secure data storage to a set of capable cloud service providers (CSPs), who then jointly evaluate MPC to provide reliable SCA services.
-For the SPECIAL prototype, we focus on the server-aided-MPC model under a strong corruption assumption, allowing up to $n-1$ of $n$ servers to be compromised and collate with others.
-While, as we discussed before, SPECIAL’s core design principles are protocol-agnostic.
-This enables interoperability with various MPC models, including the peer-to-peer settings or a weaker corruption setting (for performance purpose) where a supermajority of servers need to be honest (Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023).
+The first is peer-to-peer (P2P) paradigm (Aggarwal et al., 2005; Mouris et al., 2024; Poddar et al., 2021; Volgushev et al., 2019; Wang and Yi, 2021), where the goal is to improve efficiency by decomposing analytical queries and pushing them to data owners, so that they can either directly process in clear or running MPCs across a small group of parties.
+Unfortunately, this approach imposes large overhead on data owners, especially for complex operations like joins.
+Given that real-world data owners often lack robust computing resources and service capabilities, the P2P paradigm is hard to scale and support reliable SCA services to external analysts.
+The other paradigm is the server-aided-MPC model (Wang et al., 2021; Bater et al., 2018, 2020; Wang et al., 2022; Zhang et al., 2023; Roy Chowdhury et al., 2020; Kamara et al., 2011; Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023; Mohassel and Zhang, 2017).
+This model allows data owners to outsource both expensive MPC computations and secure data storage to a set of capable servers, which can then jointly evaluate MPC to provide reliable SCA services.
+SPECIAL is built upon the server-aided-MPC model and under a strong “all but one”corruption assumption.
+Moreover, SPECIAL’s core design is protocol-agnostic, which allows interoperability with various MPC models, including the P2P or a weaker corruption where a supermajority of servers need to be honest (Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023).
 DP leakages.
 Leakage-abuse attacks (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016), exploit data-dependent processing patterns, are persistent threats to SCA systems.
-To mitigate these risks, oblivious computation techniques (Keller and Scholl, 2014; Pinkas and Reinman, 2010; Stefanov et al., 2011; Chang et al., 2022; Mishra et al., 2018; Crooks et al., 2018; Jönsson et al., 2011; Sasy et al., 2022; Asharov et al., 2020; Wang et al., 2014; Pettai and Laud, 2015; Wang et al., 2023; Tinoco et al., 2023) have become the de facto solution—processing data such that the entire execution transcript (e.g.
-memory traces and read/write volumes) is padded to a worst-case constant.
-While these techniques ensure the strongest privacy guarantees by eliminating data-dependent leakages, they also introduce a fundamental contention with modern database optimizations, which often rely heavily on data-dependent operations (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022).
-To this end, many recent efforts seek a practical balance in the privacy-performance trade-off by allowing controlled leakage under DP (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Groce et al., 2019; He et al., 2017; Zhang et al., 2023; Qin et al., 2022; Chan et al., 2022; Wagh et al., 2016; Chu et al., 2021; Qiu et al., [n.d.]; Patel et al., 2019).
+To mitigate these risks, oblivious computation (Keller and Scholl, 2014; Pinkas and Reinman, 2010; Stefanov et al., 2011; Chang et al., 2022; Mishra et al., 2018; Crooks et al., 2018; Jönsson et al., 2011; Sasy et al., 2022; Asharov et al., 2020; Wang et al., 2014; Pettai and Laud, 2015; Wang et al., 2023; Tinoco et al., 2023) have become the de facto solution.
+While this technique ensures the strongest privacy guarantees by eliminating any data-dependent leakages, it also introduces a fundamental contention with modern database optimizations, which often rely heavily on data-dependent operations (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022).
+To this end, many recent efforts seek a practical balance in the privacy-performance trade-off by allowing controlled leakage under DP (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Groce et al., 2019; He et al., 2017; Zhang et al., 2023; Qin et al., 2022; Chan et al., 2022; Wagh et al., 2016; Chu et al., 2021; Qiu et al., 2023; Patel et al., 2019).
 However, a common issue of these approaches is unbounded privacy loss.
-While some works propose to address this (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., [n.d.]), their approaches are restricted to only simple linear queries.
-Furthermore, existing DP leakage designs also lack lossless guarantees for complex query processing.
-The inherent randomness of DP noises (Wang et al., 2021; Groce et al., 2019; Wang et al., 2022) and uncertainty in complex join sensitivities (Chu et al., 2021; Bater et al., 2018; Wang et al., 2022) often lead to improper output compaction and lose tuples.
-SPECIAL addresses all these limitations together, and to our knowledge, is the first system to ensure both bounded privacy and lossless results for securely processing complex queries.
+While some works propose to address this (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., 2023), their approaches are restricted to only simple linear queries.
+SPECIAL addresses all these limitations together, and to our knowledge, is the first SCA system that can simultaneously ensure both bounded privacy and lossless results for complex SPJA queries.
 SCA query planning.
 Query planning (Silberschatz et al., 2011) is crucial in conventional databases.
-A few studies (Bater et al., 2017; Volgushev et al., 2019; Poddar et al., 2021; Liagouris et al., 2023) explored query planning within SCA frameworks.
-Unlike conventional planners, which exploit size disparities across different execution plans to choose the most efficient one with minimized plan sizes (Elmasri et al., 2020; Harmouch and Naumann, 2017; Han et al., 2021), these methods typically prohibit the use of such data-dependent information.
-Consequently, they primarily rely on data-independent metrics for planning.
-For example, (Bater et al., 2017; Liagouris et al., 2023; Volgushev et al., 2019) optimize based on the MPC execution costs, while (Poddar et al., 2021) focuses on decomposing large MPC tasks into smaller 2/3PC ones.
-Since these methods do not optimize or compact intermediate sizes, significant query execution overhead from these large sizes remains evident, especially when processing complex multi-joins (Bater et al., 2018; Liagouris et al., 2023).
-Shrinkwrap (Bater et al., 2018) uses DP leakage models and allows planner to optimize query intermediate sizes in a privacy-preserving manner.
-However, it cannot estimate cardinalities before runtime, mandating a random selection of plan structures (e.g.
-join order) before execution.
-Consequently, planning focuses solely on optimally allocating privacy budgets across the chosen plan, which often result in suboptimal query executions.
-In contrast, SPECIAL offers an advanced query planner capable of pre-estimating plan sizes and comparing costs among different plan structures before execution.
+Conventional planners can exploit size disparities across different query plans to choose efficient ones with smaller sizes (Elmasri et al., 2020; Harmouch and Naumann, 2017; Han et al., 2021).
+However, such techniques use data-dependent information and are typically prohibited in SCA.
+A handful of studies (Bater et al., 2017; Volgushev et al., 2019; Poddar et al., 2021; Liagouris et al., 2023) that explore query planning within SCA frameworks primarily rely on data-independent metrics for planning, which usually lead to only moderate optimizations.
+Shrinkwrap (Bater et al., 2018) introduced a private planning method that optimally compacts intermediate sizes by efficiently allocating privacy budgets to minimize dummy data.
+However, it cannot pre-determine an optimal join order.
+SPECIAL offers an advanced query planner capable of pre-estimating intermediate sizes and comparing execution costs among different plan structures before runtime.
 9. Conclusion
@@
 This is achieved through a novel synopses-assisted SCA design, where a set of private table statistics are released with one-time privacy cost to guide subsequent secure SCA planning and processing.
-The benchmark results demonstrate that SPECIAL significantly outperforms SOTA SCA systems and scales complex multi-join queries to multi-million datasets
-Appendix A Privacy Proof
-Theorem A.1.
-The one-sided Laplace mechanism as described in Definition 4.1 satisfies $(\epsilon,\delta)$-DP.
-Proof.
-WLOG.
-we illustrate proof details by assuming positive mechanism, $\mathsf{Lap}^{+}$, and by symmetric, the properties proved here also holds for $\mathsf{Lap}^{-}$.
-In general, we write the mechanism as $\mathcal{M}(D)\leftarrow f(D)+\mathsf{Lap}^{+}(\epsilon,\delta)$, where $f$ is a counting query.
-Note that the sensitive for counting query is 1, and thus we will implicitly omit the sensitivity notations.
-Given any two neighboring databases $D$ and $D^{\prime}$, and let $c\leftarrow f(D)$ and $c^{\prime}\leftarrow f(D^{\prime})$, WLOG.
-we will assume $c\leq c^{\prime}$.
-Now we will evaluate three disjoint output domains for $f$, namely $O_{1}=(-\infty,c)$, $O_{2}=[c,c^{\prime}]$, and $O_{3}=(c^{\prime},+\infty)$.
-Since the noise is non-negative, thus it’s clear that $\textup{Pr}[\mathcal{M}(D)\in O_{1}]=0$ and $\textup{Pr}[\mathcal{M}(D^{\prime})\in O_{1}]=0$.
-Next, we evaluate the domain $O_{3}$ and consider an auxiliary domain $O^{*}=\{x-c^{\prime}:x\in O_{3}\}$.
-We compute the following ratio:
-(6) $\begin{split}\frac{\textup{Pr}[\mathcal{M}(D)\in O_{3}]}{\textup{Pr}[\mathcal{M}(D^{\prime})\in O_{3}]}&=\frac{\textup{Pr}[c+\mathsf{Lap}^{+}(\epsilon,\delta)\in O_{3}]}{\textup{Pr}[c^{\prime}+\mathsf{Lap}^{+}(\epsilon,\delta)\in O_{3}]}\\ &=\frac{\textup{Pr}[\mathsf{Lap}^{+}(\epsilon,\delta)\in O^{*}]}{\textup{Pr}[\mathsf{Lap}^{+}(\epsilon,\delta)\in O^{*}+(c^{\prime}-c)]}\\ &=\frac{e^{-\epsilon|x-\mu|}}{e^{-\epsilon|x-(c^{\prime}-c)-\mu|}}\\ &=e^{\epsilon(|x-(c^{\prime}-c)-\mu|-|x-\mu|)}\\ &\leq e^{\epsilon(|c^{\prime}-c)|)}~{}~{}(\textup{triangle~{}inequality})\\ &=e^{\epsilon}\end{split}$
-Then we evaluate the last output domain $O_{2}$, and we can see that $\textup{Pr}[\mathcal{M}(D^{\prime})\in O_{2}]=0$ and $\textup{Pr}[\mathcal{M}(D)\in O_{2}]\geq 0$, and thus there will be unbounded privacy loss within the domain of $O_{2}=[c^{\prime},c]$.
-However, we can bound this probability by $\delta$, which is equivalent to prove that with density function
-$\textup{Pr}[z=x]=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon|x-\mu|},$
-where $\mu=1-\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))$, the probability of sampling a value $z$ within the range $(-\infty,1]$ is less than or equal to $\delta$.
-Since we consider a discrete distribution, thus we compute the following
-(7) $\begin{split}\Pr[z<1]&=\sum_{x=-\infty}^{0}\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon(\mu-x)}\\ &=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon\mu}\sum_{x=-\infty}^{0}e^{\epsilon x}\\ &=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon\mu}\frac{1}{1-e^{-\epsilon}}\quad\text{(geometric series sum of }e^{\epsilon x})\\ &=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon\left(1-\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))\right)}\frac{1}{1-e^{-\epsilon}}\quad\text{(substitute }\mu)\\ &=\frac{e^{\epsilon}}{e^{\epsilon}+1}e^{-\epsilon+\ln(\delta(e^{\epsilon}+1))}\\ &=\frac{e^{\epsilon}}{e^{\epsilon}+1}e^{-\epsilon}\delta(e^{\epsilon}+1)\\ &=\delta.\end{split}$
-Combining all these we obtain the following
-(8) $\begin{split}\textup{Pr}[\mathcal{M}(D)\in O]&=\textup{Pr}[\mathcal{M}(D)\in(O_{1}\cup O_{2}\cup O_{3})]\\ &=\textup{Pr}[\mathcal{M}(D)\in O_{2}]+\textup{Pr}[\mathcal{M}(D)\in O_{3}]\\ &\leq\delta+e^{\epsilon}\textup{Pr}[\mathcal{M}(D^{\prime})\in(O_{2})]\end{split}$
-The claim thus holds. ∎
-Next we prove Theorem 4.3.
-Proof.
-(Theorem 4.3) We prove this theorem by analyzing privacy loss using composition theorems.
-Generating noisy histograms with one-sided noise is equivalent to releasing multiple counting queries across disjoint datasets.
-Therefore, by the parallel composition theorem (Dwork et al., 2014), each histogram release is $(\epsilon,\delta)$-DP.
-The release of the maximum frequency via a noisy max mechanism is $(\epsilon,0)$-DP.
-Given $c$ pairs, the sequential composition theorem (Dwork et al., 2014) applies, resulting in a total privacy loss under $2c$-fold advanced composition.
-Consequently, the entire process adheres to $(\hat{\epsilon},\hat{\delta})$-DP, with $\hat{\epsilon}\leq 6\epsilon\sqrt{c\ln(1/\delta)}$.
-Theorem A.2.
-The privacy of SPECIAL adheres to Definition 3.1
-Proof.
-Since the entire DP synopses releasing satisfies DP.
-Thus, we prove this theorem by constructing a simulator, $\mathcal{S}$, which takes only the DP synopses as input and can simulate execution transcripts that are indistinguishable from those of the actual protocol.
-For ease of notation, we abstract SPECIAL as a composed secure protocol $\pi$.
-We first examine a hybrid protocol, $\pi_{1}$, differing from $\pi$ in that the secure outsourcing procedure—where owners upload secret shares of their data to two servers—is replaced by a simulator, $\mathcal{S}_{1}$.
-This simulator, which takes no input, simulates secret shares with randomly sampled data $(x_{1},x_{2})\xleftarrow[]{\mathsf{rd}}\mathbb{Z}_{2^{32}}$.
-We say that a $\mathcal{S}_{1}$ must exist that produces results indistinguishable from the secret shares in $\pi$, as the absence of such a simulator would imply the non-existence of secure secret sharing techniques.
-Next, we consider another hybrid protocol, $\pi_{2}$, which differs from $\pi_{1}$ only in that the query planning function is replaced by a simulator $\mathcal{S}_{2}$.
-This simulator takes as inputs all synopses and a set of random values (simulated secure outsourced data) generated by $\mathcal{S}_{1}$, and produces a simulated execution plan for every requested query.
-We say that an $\mathcal{S}_{2}$ should also exist with outputs indistinguishable from $\pi$’s execution plans.
-Given that Selinger optimizers can pre-generate execution plans using only table statistics, the non-existence of such a simulator, $\mathcal{S}_{2}$, would imply the non-existence of Selinger optimizers, which is contradictory.
-Finally, we construct $\mathcal{S}=(\mathcal{S}_{1},\mathcal{S}_{2})$, where $\mathcal{S}_{1}$ simulates secure data outsourcing and $\mathcal{S}_{2}$ simulates query planning.
-$\mathcal{S}$ executes the query plan using randomly generated values from $\mathcal{S}_{1}$ through MPC computation.
-Since both the execution plans and input data are indistinguishable from those in $\pi$, the MPC execution must also be indistinguishable.
-Otherwise, it would violate the security guarantees of MPC.
-∎
-Appendix B Query Workloads
-We illustrate the benchmark query workloads as follows:
-Listing 1: Benchmark query workload
+Acknowledgements.
+This work was supported by the NSF grant OAC-2419821, IUIAS Collaborative Award, and IU Faculty Startup Grant.